A Complete Bibliography of Publications in The Journal of Supercomputing: 2025–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 May 2025 Version 1.04

Title word cross-reference

(K_{11} − C_{11})^{n} [218]. (n, k) [562]. ^{2} [519, 661]. C [448]. γ [250]. H [216]. K [23, 65, 79, 274, 466, 532, 647, 712]. p [53, 197]. Q [103, 625]. r, s, t [173]. Z [230].

-Block [230]. -dominant [216]. -factor-critical [466]. -graphyne [250]. -means [65, 274, 448, 532, 712]. -nearest [23, 79, 647]. -network-based [625]. -rung [103]. -sequence [197]. -spherical [173]. -star [562]. -type [53].

1 [201]. 1-Extra [164]. 16 [714]. 19 [113, 698].

2-connected [511]. 2-disjoint [273].

2-transformer [695]. 2-tuple [103]. 2.0 [405]. 23 [405]. 2D [203, 639].

3-component [164]. 3-path-connectivity [245]. 3000 [38]. 3D [19, 39, 276, 459, 465, 476, 497, 689].

4 [464]. 4-D [464]. 4D [482].

512 [362]. 5G [72, 335, 606, 690]. 5G/ [606].

6G [567, 589, 606].

AAMB [700]. ABIDS [713]. ABIDS-VEM [713]. accelerated [463]. Accelerating [322, 356, 690]. Acceleration [264, 371, 602]. accelerator [190, 599]. accelerators [23, 548, 586]. access [44, 91, 213, 233, 403, 445, 549]. access-control [91]. accident [558, 703]. AccPRET [548]. accuracy [408, 720]. accurate [354, 380, 558]. achieve [553]. ACO [257]. acoustic [573, 608]. acquiring [631]. across [249]. action [10, 12, 37, 318, 417]. activity [349]. actor [357]. actor-critic [357]. adaptation [383, 510, 588]. adapting [137]. Adaptive [66, 80, 92, 98, 112, 139, 212, 302, 343, 399, 468, 491, 511, 526, 530, 585, 601, 614, 640, 653, 663, 683, 688, 695]. adder [492]. additive [563]. Addressing [629]. ADH [515]. ADH-YOLO [515]. adjusted [513]. ADOX [212]. ADOX-SSA [212]. ADV [34]. ADV-YOLO [34]. advanced [559, 617]. Advancements [369]. advancing [19, 418]. adversarial [168, 239, 289, 349, 623]. AE [338]. AE-UNet [338]. Aerial [95, 105, 302, 423, 502, 618]. affine [510]. against [530]. agent [150, 279, 388]. aggregate [621]. aggregation [43, 63, 269, 318, 703]. aggregations [294]. agreement [9, 149]. agriculture [280]. AI [272, 280, 347, 356, 390, 450, 551, 565, 615, 616]. AI-based [280]. AI-driven [272, 390]. AI-powered [565, 616]. AIGC [426]. AIGC-generated [426]. AIoT [472]. air [73, 401]. airborne [491]. aircraft [344, 563, 590]. airport [515]. AISBench [450]. Aitken [264]. ALCARO [337]. algebra [334]. algebra-enhanced [334]. Algebraic [53]. algorithm [7, 8, 13, 18, 21, 22, 32, 47, 52, 54, 55, 66, 67, 79, 91, 94, 98, 99, 110, 113, 116, 124, 135, 137, 144, 150, 170, 182, 188, 202, 209, 227, 236, 244, 251, 252, 261, 264, 266, 285, 290, 292, 295, 302, 320, 325, 329, 330, 355, 360, 372, 381, 384, 387, 391, 399, 412, 441–443, 449, 460, 481, 493, 504, 529, 530, 533, 544, 570, 578, 584, 592, 595, 609, 610, 612, 620, 626, 633, 636– 638, 657, 685, 697, 699, 708, 712, 714, 716]. algorithm-based [137]. algorithms [20, 88, 132, 172, 247, 257, 478, 485]. alignment [365, 383, 612]. allergic [53]. allocation [11, 118, 192, 237, 340, 367, 469, 514, 570, 574]. ALMASH [301]. along [386]. alternative [708]. altitude [423]. Alzheimer [432]. AlzONet [432]. analysis [1, 9, 46, 51, 53, 55, 58, 64, 86, 93, 131, 136, 179, 195, 336, 348, 349, 358, 456, 487, 494, 507, 508, 543, 562, 613, 624, 630, 649, 658, 673, 719]. analytical [268]. analytics [559]. Analyzing [611, 639, 665]. Android [501]. ANFIS [201]. angle [98]. angle-based [98]. annealing [147]. annotation [600]. Anomalies [243]. anomaly [195, 282, 297, 327, 382, 413, 457, 505, 628, 713]. anomaly-based [713]. anonymity [301]. anonymity-based [301]. anonymization [134]. ant [47]. antenna [325]. antenna-array [325]. Antivirus [151]. any [475]. AODV [386]. AP [72, 340]. Apache [214, 518]. API [594]. APPBoost [530]. Application [7, 53, 118, 197, 205, 280, 294, 630, 660, 686]. applications [94, 108, 109, 136, 167, 173, 253, 367, 437, 454, 601, 671, 691]. applied [159, 431]. approach [1, 6, 24, 29, 58, 88, 92, 123, 128, 134, 137, 142, 165, 166, 184, 198, 206, 207, 254, 260, 263, 274, 283, 285, 321, 323, 346, 369, 412, 428, 433, 467, 488, 489, 511, 518, 538, 559, 573, 577, 651, 661, 664, 698, 717]. approaches [20, 76, 354]. approximate [109, 121, 408]. approximation [468]. Arabic [187]. arbitrary [305]. arbitration [483]. architecture [172, 252, 271, 337, 351, 366, 548, 567, 672]. architectures [322]. area [6, 291, 527]. ARGCN [86]. ARIMA [373]. arithmetic [23, 302]. ARM [214]. Arnold [39]. arrangement [218]. arrangements [39]. array [325, 586]. artery [267, 692]. articles [388]. Artificial [62, 110, 150, 270]. ASOD [618]. ASOD-YOLOX [618]. Aspect [46, 64, 125, 131, 189, 200, 348, 543, 634, 658, 673]. Aspect-aware [64]. Aspect-based [64, 131, 348, 543, 658]. aspect-focused [348]. aspect-level [46, 125, 189, 200, 673]. Assessing [180, 662]. assessment [73, 205, 285, 445, 471]. Assignment [95]. assist [419]. assisted [8, 149, 165, 231, 315, 589, 660]. associated [243]. association [335, 713]. association-based [335]. associations [36, 72]. assurance [247]. AST [600]. asynchronous [7, 496, 635]. atomic [494]. atomistic [554]. Attack [62, 453, 553, 659]. attacks [69, 252, 320, 484, 518]. attention [18, 26, 28, 31, 40, 70, 74, 82, 115, 125, 126, 157, 191, 199, 240, 242, 267, 293, 324, 338, 352, 414, 438, 480, 484, 539, 541, 563, 573, 604, 608, 632, 634, 652, 693, 711]. attention-based [324]. attention-enhanced [125]. attentional [86]. attribute [44, 167, 181, 430, 540]. attribute-based [44, 430]. attributes [712]. attribution [25]. audio [683]. audio-visual [683]. Audiovisual [31]. auditing [551]. augmentation [114, 208, 382, 542]. augmented [185]. authenticated [9, 149]. authentication [5, 119, 233, 301, 700]. authentication-based [119]. authorial [151]. authorization [655]. auto [228, 353, 480, 596]. auto-encoder [353, 480]. auto-encoder-based [596]. auto-negotiation [228]. autoencoder [5, 16, 62, 118, 231, 234]. autoencoder-based [234]. automata [691]. Automated [1, 613]. automatic [610]. automatically [150]. automation [690]. automaton [99]. autonomic [367]. autonomous [386, 387]. autotuning [283]. Availability [445]. average [99]. average-roulette [99]. avoidance [617]. AVX [362]. AVX-512 [362]. aware [2, 3, 15, 27, 45, 52, 64, 121, 122, 226, 299, 306, 383, 384, 435, 495, 687, 692, 693, 705]. awareness [84, 382, 651].

B5G [550]. backoff [213]. backpropagation [405]. backscatter [369]. backup [444]. backward [65]. bag [285]. balance [468, 621]. Balanced [60, 273, 683]. Balancing [52, 109, 335, 452, 479, 657]. bald [251]. Bandwidth [175]. bank [278]. barbed [222]. base [27]. Based [4, 6, 9, 10, 16, 18, 21, 22, 24, 31, 33, 34, 37, 39, 42, 57, 62, 65, 66, 68, 69, 85, 89, 98, 100, 113, 121, 124, 129, 144, 153, 160, 162, 171, 173, 181, 210, 215, 240, 241, 250, 255, 270, 280, 302, 304, 329, 330, 340, 344, 350, 355, 358, 360, 370, 376, 379, 385, 386, 388, 389, 392, 411, 415, 422, 424, 439, 440, 449, 456, 457, 468, 473, 478, 484, 493, 497, 504, 509, 531, 540, 552, 555, 556, 574, 577, 592, 595, 598, 601, 603, 609, 625, 628, 635, 641, 651, 653, 657, 658, 666, 672, 676– 678, 683, 690, 706, 712, 715, 716]. based [5, 8, 12–14, 25, 28, 35, 40, 44, 47, 54, 58, 64, 76, 79, 96, 97, 110, 119, 123, 126, 128, 130, 132, 134, 137, 140, 142, 143, 156, 172, 184, 188, 189, 196, 203, 209, 212, 219, 227–229, 232, 260, 262, 266, 267, 276, 277, 290, 292, 300, 301, 324, 325, 331, 335, 337, 342, 348, 356, 359, 372, 380, 396, 414, 430, 433, 442, 444, 447, 460, 461, 463, 472, 485, 486, 495, 500, 501, 505, 527, 532, 535, 541, 543, 544, 561, 563, 570, 571, 596, 600, 613, 618, 627, 638, 640, 647, 654, 659, 664, 700, 707, 711, 713]. based [49, 67, 111, 131, 148, 169, 213, 221, 234, 248, 264, 269, 351, 357, 464, 471, 481, 498, 502, 515, 520, 537, 546, 566, 575]. bases [243]. basis [152]. bat [137]. Bayesian [55]. BDLT [271]. BDLT-IoMT [271]. beam [275, 656]. bearing [262]. Beetle [90, 412]. behavior [467, 552, 630, 649, 707]. behaviour [284]. benchmark [450, 485, 649]. benefits [662]. Bernoulli [201]. BERT [163, 187, 324, 388]. Beyond [72, 354, 589, 606, 690]. BFL [385]. BFT [496]. bi [31, 177, 349, 363, 541, 544]. bi-clique [363]. bi-directional [349]. bi-layer [31]. bi-level [541]. bias [552]. bidirectional [117]. big [165, 181, 629]. Bilateral [136]. Bilinear [446]. BiLSTM [162, 512]. binary [190, 209, 225, 363]. binary/ternary [190]. biomedical [171]. biometric [540, 583]. bionic [236]. bipartite [48, 363]. Bipolar [294]. bird [524]. bit [183]. bit-level [183]. Bitcoin [77]. bits [210]. BKA [471]. BL [512]. black [626]. blindness [526]. block [77, 160, 224, 230, 431]. blockage [523]. blockchain [61, 97, 111, 128, 166, 219, 232, 271, 324, 328, 339, 385, 420, 473, 479, 508, 534, 598]. blockchain-based [473, 598]. blockchain-enabled [479]. blockchain-federated [385]. blockchains [700]. blood [419, 674]. Bluelight [575]. BN [557]. body [106, 349]. body-worn [349]. boosted [375]. boosting [229, 530]. bot [611]. botnet [518]. bound [102]. bounds [547]. BR [55]. brain [19, 235, 352, 432, 512, 689]. branch [127, 519, 604]. breakpoints [222]. breast [679]. broadcast [247]. broadcasting [635]. brooms [53]. bubble [164]. bubble-sort [164]. budget [286]. budget-feasible [286]. building [711]. built [109]. built-in [109]. bullying [284]. burden [621]. bushing [372]. BXIv3 [537]. BXIv3-based [537]. Byzantine [716].

C&M [512]. C2 [337]. C3 [571]. CA [442]. cache [45, 409]. caching [651]. calculation [218]. calculations [475]. calculus [90]. camera [25]. camouflage [157]. camouflaged [572]. cancellation [608]. cancer [679]. Canny [230]. capabilities [196]. Capsule [86, 433, 499]. caption [658]. capuchin [236]. Caputo [720]. carbon [498]. cardiovascular [411]. Carlo [381, 575]. carotid [267]. carrier [589]. Cartesian [529]. case [315, 358, 454, 456, 554]. cases [686]. cat [113, 325]. categories [213]. categorization [342]. categorization-based [342]. category [382]. category-guided [382]. causal [247, 365, 393, 588]. causal-contrastive [588]. cause [195]. cavity [392]. CBM [155]. CBR [77]. CD [190]. CD-MAC [190]. CDMANet [242]. CDNRocks [57]. CE [592]. cell [419, 469, 663, 691]. Cellular [99, 329, 340, 691]. center [4, 211]. centers [184]. central [242]. centric [166, 549]. CEPDNet [100]. Cerebral [106]. certificateless [667]. CGWRIME [375]. chain [89, 379, 403, 574]. chain-based [89]. chains [331]. challenges [291, 662]. Change [54, 711, 715]. change-point [715]. changes [71]. Channel [126, 169, 252, 270, 414, 465, 471, 479, 509, 604]. chaotic [108, 152, 533]. character [631]. chargers [298]. charging [298, 485]. ChatGPT [643]. chatHPC [194]. Chebyshev [179, 252, 531]. ChebyshevNet [179]. check [296, 708]. chemical [359]. China [358, 456]. Chinese [50, 200]. Chinese-oriented [200]. chlorophyll [265]. chlorophyll-a [265]. churn [178]. CILOS [103]. CIMS [229]. CIMS-based [229]. cipher [224]. circuit [117, 361, 398, 460]. cities [97, 412, 518]. city [651]. class [16, 74, 167, 248, 644, 706]. class-consistent [167]. class-imbalanced [248]. class-wise [16]. classical [354, 465]. Classification [23, 26, 54, 66, 101, 114, 125, 126, 189, 200, 217, 248, 285, 351, 352, 377, 383, 394, 402, 404, 414, 418, 510, 512, 519, 553, 622, 644, 650, 668, 672, 675, 684, 688]. classifier [154]. classify [538]. CLDE [569]. client [69, 196, 564]. client-side [69]. CLIP [174]. CLIP-driven [174]. clique [363]. clockwork [421]. closure [67]. Cloud [4, 12, 65, 89, 127, 128, 135, 141, 165, 176, 181, 237, 295, 368, 394, 430, 433, 472, 479, 484, 551, 556, 577, 601, 603, 625, 629, 633, 657, 671]. cloud-assisted [165]. Cloud-based [12, 472]. cloud-edge [551]. cloud-edge-terminal [625]. Cloud-WAVECAP [433]. Cluster [2, 480, 574, 630]. cluster-based [574]. Cluster-guided [480]. Clustering [2, 65, 98, 135, 143, 150, 171, 193, 209, 274, 386, 397, 448, 607, 654, 663, 712]. Clustering-based [143, 171]. clusters [104, 436, 470, 706]. CNN [5, 24, 100, 144, 162, 412, 463, 472, 512, 642, 689]. CNN-based [5, 24, 100]. CNN-swin [689]. CNNs [60, 408]. co [462, 584]. co-execution [462]. co-location [584]. coalition [132]. coalitions [570]. Coarse [490]. Coarse-to-fine [490]. code [496, 600]. coding [85, 210]. coefficient [401]. coefficients [681]. coevolutionary [59, 287]. cognitive [133, 559]. cold [234]. collaboration [84, 158, 375, 507]. Collaborative [54, 59, 79, 114, 124, 287, 376, 403, 497, 570]. collection [123]. color [108]. colored [247]. coloring [316]. combat [73, 279]. Combating [69]. combined [161, 333]. combining [94, 258]. Comment [149]. commerce [229]. commercialization [163]. common [584, 715]. communication [11, 283, 340, 369, 541, 550, 589, 656]. communication-constrained [11]. communications [277]. community [182, 216, 460]. compact [77]. compactions [156]. compensation [109, 673]. competition [375]. competition-boosted [375]. Competitive [569, 590]. compiler [451]. complementarity [572]. complete [180, 341]. completion [585]. Complex [36, 159, 182, 411, 513, 582, 626, 718]. component [120, 164, 562]. composite [338]. composites [13]. compound [359]. comprehension [50]. Comprehensive [291, 366, 485, 680, 716]. compressed [144]. compression [96, 399]. computable [57]. computation [175, 469, 576]. computational [152, 487]. computer [172, 581]. computing [15, 24, 29, 35, 41, 52, 76, 84, 100, 116, 121, 127, 186, 190, 192, 196, 254, 257, 288, 289, 295, 317, 367, 384, 386, 408, 427, 428, 430, 435, 445, 479, 518, 551, 577, 591, 615, 626, 633, 653, 657, 662, 705, 710]. concave [327]. concentrations [265]. conception [170]. concise [197, 584]. condition [576]. Conditional [14, 238, 521, 717]. conditioned [30]. conditions [718]. confluence [339]. conjugate [55]. connected [511]. connection [250]. connections [207]. connectivity [72, 164, 180, 245, 536]. Consensus [61, 91, 97, 330, 716]. consistency [299]. consistent [167]. consolidation [426]. consortium [420]. constrained [11, 190, 286, 571, 650]. constraint [116, 132]. constraints [4, 473]. Constructing [259, 607]. construction [341, 663, 668]. constructions [564]. consumption [321, 336, 412, 470, 664]. container [15, 260, 461]. container-based [461]. content [168, 651]. content-type [651]. contention [213, 483]. context [71, 318, 422, 607, 673, 680]. contextual [176]. Continuous [5, 396]. Continuous-time [396]. contour [230]. contract [379, 520, 600]. contraction [361]. contracts [44, 379]. contrast [521]. contrastive [37, 101, 139, 153, 239, 256, 588, 596, 627]. contribution [255]. contributions [513]. control [44, 91, 185, 204, 403, 431, 640, 701]. Controllable [108, 540]. controller [444, 527]. Conv2D [557]. Conv2D-ReLU-BN [557]. ConvBiFuseNet [352]. convective [525]. convergence [148]. Convergent [475]. conversational [153]. convex [605]. ConvFormer [539]. convolution [26, 235, 566, 627]. convolutional [18, 25, 27, 193, 231, 277, 326, 348, 397, 402, 425, 446, 465, 476, 543, 693]. cooperation [439]. Cooperative [59, 287, 704]. coordinate [191]. coordination [597]. core [59, 287, 506]. cores [248]. coronary [692]. corpora [187]. Correction [287, 390, 391, 413, 414, 428, 456, 494, 603, 616, 702]. correction-optimized [702]. correlated [505]. correlation [355]. correlation-guided [355]. cost [116, 132, 288, 601, 629]. cost-effective [288]. coverage [83]. covers [48, 273]. COVID [113, 698]. COVID-19 [113, 698]. CovMedCare [339]. CPCS [439]. CPU [59, 287, 437, 462, 483]. CPU-GPU [462]. criteria [89]. criterion [1]. critic [357]. critical [466]. criticality [394]. CRKG [258]. Crns [174]. cross [127, 139, 299, 323, 440, 471, 519, 604, 688, 695, 700]. cross-aware [299]. cross-domain [139, 688, 700]. cross-enhancement [519]. cross-fusion [604]. cross-modal [127]. cross-project [323]. cross-stage [440]. crossroad [67]. crowd [707]. crowdfunding [263]. crowding [636]. crowdsensing [286, 467]. CRS [153]. cryptographic [598]. Cs [716]. Cs-pbft [716]. CT [338]. cube [180, 400, 536]. cubes [96, 341]. cubic [103]. curiosity [227]. currencies [442]. custom [648]. customer [178]. customized [405]. CV [201]. cyber [51, 142, 284, 559]. cyber-bullying [284]. cyber-physical [51, 559]. cyberattacks [714]. cyberphysical [455]. cybersecurity [271, 277]. cybertwin [439]. cylinder [185].

D [242, 464, 572]. damage [387, 563]. damped [709]. dandelion [468]. DAP [77]. DAP-CBR [77]. Dara [383]. Data [4, 41, 57, 63, 91, 123, 134, 143, 154, 165– 167, 181, 184, 210, 211, 225, 238, 271, 272, 300, 323, 349, 366, 382, 385, 390, 399, 403, 437, 444, 452, 475, 482, 505, 507, 519, 521, 542, 547, 559, 578, 584, 613, 629, 631, 643, 647, 651, 661, 678, 701, 712, 713, 717]. Data-dependent [547]. Data-driven [475, 701]. data-free [521]. database [81]. dataset [31, 649]. datasets [118, 315, 426, 452, 642]. DBSCAN [150]. DCLMD [607]. Dcsca [604]. Dcsca-Net [604]. DDoS [484, 611]. DDQN [276]. deadline [4]. debiasing [542]. deblur [440]. Decentralized [121, 479]. decimal [492]. decision [29, 229, 360, 361, 527, 556, 603]. decision-making [29]. decoder [157]. decomposition [160]. Deep [6, 35, 72, 92, 130, 161, 213, 223, 241, 280, 284, 288, 295, 322, 343, 344, 368, 380, 397, 402, 419, 432, 447, 457, 472, 557, 567, 573, 577, 602, 609, 611, 625, 628, 645, 648, 690]. deep-reinforcement-learning-based [577]. Deepat [344]. defect [32, 40, 155, 239, 261, 424, 499, 504]. defined [291, 435, 495]. Defocus [440]. deformable [490]. degree [145, 331]. degree-based [331]. degree-wise [145]. dehazing [169]. delay [52, 577]. delay-sensitive [577]. deletion [675]. deliver [263]. delivery [651]. demand [55, 485]. demonstrations [607]. denial [453]. denoising [3, 100, 159]. Density [143]. dependencies [189]. Dependency [45, 82, 125, 693]. Dependency-aware [45]. dependency-oriented [82]. dependent [203, 547, 662]. deploy [454]. deployment [84, 535, 550]. depression [93]. DepressionFeature [93]. depth [440]. depth-of-field [440]. derivation [225]. derivatives [720]. descent [328, 526]. Design [9, 148, 149, 251, 437, 492, 545, 590]. designing [254, 428]. Detecting [284, 714]. Detection [16, 22, 28, 32, 34, 40, 42, 62, 67, 75, 105, 106, 110, 142, 151, 155, 177, 191, 202, 222, 230, 231, 239, 249, 254, 261, 275, 277, 282, 288, 290, 297, 299, 303, 307, 315, 319, 320, 327, 334, 338, 372, 382, 387, 388, 391, 413, 419, 422, 424, 426, 428, 433, 447, 453, 457, 459, 460, 465, 476, 477, 484, 486, 501, 502, 504, 505, 515, 520, 526, 527, 544, 553, 560, 563, 565, 566, 571, 572, 582, 585, 592, 593, 596, 616, 618, 623, 628, 632, 642, 643, 645, 646, 650, 652, 659, 675, 680, 697, 704, 707, 708, 711, 713, 715, 718]. detection-based [460]. detector [423]. deterministic [343]. DETR [290]. DETRmg [307]. development [586]. device [9, 119, 606]. devices [232, 399, 593]. dew [149]. dew-assisted [149]. DGN [130]. Dhcache [409]. Dhcm [597]. diabetic [166]. DiabeticChain [166]. Diagnosis [111, 262, 419, 432, 472, 539, 648]. diagnosis-based [472]. diagnostic [334]. diagonal [589]. diagram [262]. diagrams [361]. difference [242, 681]. differences [662]. differencing [709]. different [715]. differential [49, 116, 333, 569, 590, 602, 609]. differentiating [213]. DiffREE [30]. Diffusion [30, 203, 208, 224, 345, 446, 720]. DiFuseR [21]. digit [23]. digital [5, 229, 478, 531, 662]. dilated [509]. dimension [123, 230, 613, 659]. dimensional [39, 136, 154, 212, 385, 505, 526, 613, 685]. directional [349]. Dirichlet [118]. disassembly [7]. disaster [294, 550]. discard [417]. discharges [1]. discover [610]. discovering [584, 698]. discovery [182]. discrete [7, 159, 529, 620]. discrete-time [159]. Discriminant [136, 404]. discrimination [717]. disease [411, 432]. disentangled [393]. disentanglement [540]. disjoint [48, 211, 273, 400]. Distance [173, 468, 511, 585, 636, 652, 715]. distance-based [715]. distance-induced [585]. DistilBERT [447]. DistilBERT-based [447]. distillation [56, 521, 696]. distinguishing [316]. distributed [11, 21, 57, 132, 228, 254, 322, 428, 443, 453, 591, 705]. distribution [195, 205, 257, 383, 607]. distribution-aware [383]. diverse [399]. divide [400, 536]. divide-and-swap [400, 536]. DML [711]. documents [163, 171]. Domain [114, 139, 306, 383, 522, 585, 588, 608, 688, 700, 711]. Domain-adaptive [688]. domains [249]. dominant [216]. dominating [511]. Double [76, 534]. double-DQN [76]. DPC [274]. DPC-empowered [274]. DPF [689]. DPF-Unet [689]. DPSMUNet [267]. DQN [76, 535]. DQN-MSRA [535]. DRCD [483]. driven [37, 174, 268, 272, 355, 390, 475, 483, 569, 634, 655, 701]. Driving [275, 288]. DRL [213]. drone [423, 621]. drowsiness [463]. DRSS [58]. drug [698]. drugs [53, 489]. DRX [336]. DSC [442]. DSC-RepVGG [442]. DSS [14]. DSSE [430]. DTCN [221]. Dual [8, 17, 19, 58, 68, 82, 127, 131, 157, 197, 267, 397, 409, 471, 474, 543, 604, 608, 636, 711]. Dual-branch [127, 604]. dual-channel [471]. dual-decoder [157]. dual-domain [608, 711]. dual-energy [8]. Dual-enhanced [543]. dual-hash [409]. dual-path [17]. dual-pooling [267]. dual-space [636]. dual-stream [19, 68]. dung [90]. dust [168]. DV [170]. DVTXAI [280]. dwarf [493]. dyed [239]. Dynamic [11, 18, 41, 44, 54, 63, 71, 77, 89, 159, 198, 348, 355, 378, 394, 396, 420, 449, 479, 491, 556, 587, 595, 597, 603, 607, 640, 687]. dynamical [35]. DYR [587]. DYR-SLAM [587].

e-commerce [229]. E-GRACL [42]. EA [424]. each [296, 691]. eagle [251]. EAR [495]. Early [401, 486, 631]. early-modern [631]. earthquake [191, 294]. echo [30, 608]. ecosystems [473, 617]. EDBLSD [366]. EDBLSD-IIoT [366]. Edge [15, 29, 35, 48, 76, 84, 100, 121, 164, 180, 192, 230, 246, 257, 270, 273, 278, 288, 317, 338, 341, 356, 363, 367, 445, 551, 565, 577, 593, 616, 625, 626, 653, 697]. edge-cloud [577]. edge-connectivity [180]. edge-independent [341]. education [91]. EEG [126, 414, 463, 605]. EEG-based [463]. effect [393]. Effective [6, 229, 288, 305, 503, 521, 532, 644]. efficiency [77, 109, 154, 184, 292, 494]. Efficient [4, 8, 15, 23, 38, 40, 50, 61, 103, 111, 113, 116, 123, 140, 149, 195, 209, 215, 237, 252, 264, 274, 298, 328, 350, 361, 362, 417, 433, 443, 463, 492, 506, 521, 549, 554, 584, 605, 650, 695, 719]. EFFN [221]. EFFN-Transformer [221]. elastic [426, 591]. elasticity [671]. elasticnet [538]. election [435]. electric [241, 321, 485]. Electrical [49]. electromagnetic [441]. electronic [403]. element [405]. elementary [633]. Ellipse [582]. elliptic [468]. embedded [356]. embedding [218, 555]. emergency [294, 358, 456]. EMGODV [140]. EMGODV-Hop [140]. emission [626]. emotion [31, 71, 573]. emphysema [315]. Empirical [642]. employing [553]. empowered [274]. Empowering [194]. emulators [528]. enabled [195, 455, 479]. encirclement [279]. encirclement-combat [279]. encoder [353, 480, 596, 658]. encoding [282]. encrypted [225]. encryption [12, 39, 63, 96, 108, 119, 181, 210, 464, 564, 665]. end [387, 438]. end-to-end [387, 438]. Energy [2, 4, 8, 15, 20, 52, 116, 123, 184, 209, 226, 237, 321, 366, 412, 436, 443, 470, 492, 494, 495, 554, 654, 661, 664]. Energy-aware [2, 495]. Energy-efficient [4, 8, 15, 116, 237, 443, 492, 554]. Energy-harvesting-aware [226]. engine [259]. engineering [148, 188, 251, 375, 493, 545, 643]. Enhanced [64, 75, 90, 110, 113, 117, 125, 140, 153, 155, 170, 220, 223, 300, 303, 326, 328, 334, 366, 377, 387, 395, 447, 459, 467, 468, 514, 525, 530, 543, 567, 587, 604, 622, 646, 679, 684, 710, 718]. enhancement [144, 160, 169, 244, 417, 422, 477, 519, 540]. Enhancing [10, 50, 72, 74, 77, 78, 178, 282, 292, 365, 401, 426, 550, 553, 573, 650, 673]. ensemble [110, 178, 613, 642, 647, 713]. entity [666]. entropy [65, 85, 167, 250]. entropy-based [65]. environment [142, 181, 237, 541, 586, 592, 601, 625, 655]. environmental [54, 626]. environments [11, 223, 384, 484, 571, 582, 629, 644, 650]. epileptiform [1]. epochs [265]. equation [49, 602, 709]. equations [609, 639, 681]. equilibrium [713]. equipment [7]. erasure [496]. ERBFT [496]. error [9, 62, 109, 487, 708]. error-resilient [109]. estate [118]. Estimating [393, 664]. Estimation [270, 438, 468, 481]. Evaluating [103, 288, 356, 643]. evaluation [247, 392, 478, 599, 642]. evasive [649]. event [126, 288, 411, 414]. event-related [126, 414]. evidence [205]. evolution [116, 569, 590, 615]. Evolutionary [54, 98, 395, 410, 449]. Evolutionary-enhanced [395]. examples [623]. execution [462]. expected [331]. Experience [214, 514]. Experience-guided [214]. experimentation [212]. experts [416]. explainable [62, 280]. explicit [709]. exploitation [462]. Exploiting [183]. exploration [357]. Exploring [227, 437, 686]. Exponential [685, 709]. Exponential-trigonometric [685]. Exponentially [204]. exponentiation [296]. exportable [475]. expression [238]. expressiveness [397]. extension [304, 405]. extensions [362]. external [281, 539]. Extra [164]. extracted [49]. extracting [299]. extraction [82, 240, 447, 634, 693]. extractive [171]. extrapolation [30].

F [661]. FaasFlows [206]. fabric

[239, 376, 707]. face [119, 540]. faced [236]. Facial [464]. facilitated [304]. factor [466, 640]. factoring [102]. Factorization [79, 87, 102, 292, 641]. factors [120, 281, 698]. failure [386]. fair [97, 379]. fake [177, 249, 299, 320]. fall [431]. FANET [274]. Fast [100, 182, 252, 478, 593]. fastener [504]. faster [534]. FastGA [505]. Fault [262, 273, 323, 372, 472, 539, 577, 640, 677, 716]. fault-tolerant [577, 640, 677]. faults [48, 183, 610]. FDPNet [645]. feasible [286]. Feature [17, 19, 28, 30, 40, 64, 106, 129, 147, 169, 193, 212, 220, 238, 254–256, 285, 299, 365, 377, 415, 422, 428, 447, 499, 501, 513, 525, 544, 563, 605, 622, 646, 647, 650, 654, 659, 668, 717]. feature-conditioned [30]. feature-guided [256]. features [36, 49, 150, 306, 353, 388, 392, 538, 573, 623]. FedBat [137]. federal [552]. Federated [26, 33, 63, 137, 158, 196, 226, 232, 246, 269, 324, 333, 385, 644, 702, 712]. feedback [89]. feedforward [346]. few [139, 596, 668]. few-shot [139, 596, 668]. Fibonacci [197, 681]. FIBS [478]. FIBS-fast [478]. field [440, 441, 475, 556, 586, 603]. file [260]. filter [532]. filtering [79, 522, 532]. financial [543]. Finding [36, 329, 363]. fine [44, 50, 368, 418, 490, 636]. fine-grained [44, 368, 418, 636]. fine-tuning [50]. finite [640]. finite-time [640]. firefighting [472]. firewalls [628]. first [23]. fisher [714]. fitness [449, 468]. fitness-distance-similarity [468]. fitting [230]. five [1, 39]. five-criterion [1]. five-dimensional [39]. flexible [293, 488, 564]. flight [488]. flight-follow [488]. floating [487]. floating-point [487]. flock [669]. flock-inspired [669]. flow [129, 198, 281, 437, 443]. flying [550]. FM [201]. FM/FM/1 [201]. FMO [714]. focal [16, 511]. focused [348]. fog [52, 116, 384, 386, 435, 518]. foggy [617]. folded [327, 536]. folded-concave [327]. follow [488]. FootprintNet [583]. footprints [583]. forecast [440]. forecasting [20, 49, 55, 68, 115, 281, 425, 498, 509, 579, 695]. foreign [646]. forensics [522]. forest [7]. forests [553]. forgery [645]. foundation [554]. Fourier [193, 425]. FPGA [121, 351, 463, 491]. FPGA-accelerated [463]. FPGA-based [121]. FPGAs [268]. fractal [613]. Fractional [90, 152, 425, 720]. frame [675]. framework [50, 69, 89, 101, 177, 234, 280, 319, 324, 338, 380, 432, 498, 533, 539, 598, 601, 613, 641, 658, 675, 712]. free [140, 521]. frequency [262, 349, 392, 522, 608]. frequent [578]. Frobenius [204]. frontier [452]. fronts [98]. fulfillment [1]. full [316]. fully [12, 296, 476]. fully-verifiable [296]. function [33, 90, 152, 327, 496, 574, 654]. functional [564]. functions [633]. Fusion [22, 28, 33, 43, 49, 51, 106, 129, 138, 158, 198– 200, 251, 262, 266, 306, 319, 352, 389, 394, 397, 402, 415, 416, 440, 477, 501, 557, 563, 604, 610, 654, 656, 659, 661, 685, 689, 690]. future [36, 291]. Fuzzy [29, 73, 103, 167, 173, 209, 294, 360, 411, 448, 640, 661, 701]. fuzzy-based [640].

GA [346]. GA-DE [346]. gain [325]. gait [380, 660]. Game [178, 228, 329, 410, 412, 497, 691]. game-based [329]. gaming [330]. GARCH [373]. gate [586]. gating [82, 125]. Gaussian [205, 269]. gazelle [140]. GBADroid [501]. GDNet [160]. GDRNet [509]. GDT [527]. GDT-IDS [527]. GEMM [614]. GEMV [506]. gene [238]. general [101, 531, 576]. generalization [114, 306, 547]. Generalized [102, 159, 217, 448, 681, 703]. generated [289, 426]. generating [645, 649]. generation [33, 132, 138, 151, 359, 386, 600, 610, 696]. Generative [239, 258, 289, 349]. generator [94]. GenerCTC [101]. genetic [88, 570, 609]. genome [362]. genotype [570]. geographical [103]. Geohash [85]. geolocation [223]. geometric [334]. geophysical [709]. Geospectra [223]. gesture [546]. GFIDF [319]. GHG [372]. Ghost [61, 160]. Ghost-Weight [61]. global [90, 158, 333, 384, 493]. GNN [215]. GNN-based [215]. Golden [545]. good [26]. goore [329, 691]. GPT [359]. GPT-based [359]. GPU [59, 270, 283, 287, 356, 368, 437, 462, 483, 549, 579]. GPU-based [270, 356]. GPU-centric [549]. GPUs [21, 580, 614]. GRACL [42]. gradient [3, 55, 328, 526, 683]. gradient-aware [3]. gradients [343]. gradual [319]. grained [44, 368, 418, 636]. granularity [46, 56, 122, 220]. granulator [523]. Graph [18, 27, 41–43, 76, 86, 107, 112, 125, 129, 138, 176, 193, 198, 215, 248, 293, 326, 348, 396, 415, 446, 480, 486, 498, 510, 520, 527, 543, 554, 585, 612, 641, 677, 693, 696]. graph-based [76, 520, 527, 677]. Graph-induced [43]. graphics [371]. graphs [48, 120, 145, 218, 245, 316, 329, 363, 466, 555]. graphyne [250]. Grassmann [404]. greedy [182]. grids [714]. Ground [276, 433]. Ground-based [433]. group [199, 227]. grouped [614]. grouping [509]. groups [197]. GRU [472, 676]. GRU-transformer [676]. GSW [222]. GUFORMER [3]. guided [80, 168, 214, 256, 355, 365, 374, 382, 480, 710]. Gumbel [33]. GWO [184].

Hadoop [655]. Hamacher [294]. handling [294]. handover [435, 606]. hardware [23, 252, 351]. harmonic [448]. Harris [148, 620, 657]. harvesting [8, 226]. harvesting-assisted [8]. hash [409]. hate [28]. hawks [148, 620, 657]. hazy [515]. head [2, 31]. Healthcare [141, 181, 285, 301, 598]. Hebei [358, 456]. herd [610]. heterogeneity [196]. heterogeneous [11, 38, 41, 112, 167, 268, 279, 283, 335, 340, 436, 443, 483, 491, 520, 529, 570, 601, 629, 632, 642, 647, 699, 710]. heuristic [257, 346, 381]. heuristics [381]. hexagonal [222]. HFSL [196]. hiding [210]. hierarchical [195, 365, 621]. hierarchy [597]. HIFNet [718]. High [81, 108, 154, 224, 260, 275, 277, 325, 351, 354, 405, 461, 492, 505, 526, 551, 576, 584, 615, 623, 662, 720]. high-accurate [354]. high-dimensional [154, 526]. high-efficiency [154]. high-level [461, 623]. high-performance [260, 351, 551, 615, 662]. High-precision [277]. high-security [108]. High-speed [492, 576]. High-utility [81]. higher [203]. higher-order [203]. highly [61]. Hilbert [605]. histopathological [402]. HMAC [534]. holes [389]. homomorphic [12, 63]. honey [181]. honey-based [181]. Hop [140, 170]. horse [188, 610]. hot [88]. HPC [194, 436, 454]. HSVDetector [520]. Human [12, 37, 170, 349, 481, 590, 680]. human-object [680]. human-powered [590]. hummingbird [110]. Hybrid [13, 19, 20, 68, 72, 83, 144, 210, 254, 324, 366, 380, 388, 428, 441, 457, 462, 463, 465, 472, 474, 492, 493, 497, 533, 638, 642, 695, 710, 717, 719, 720]. HydraGNN [554]. hydraulic [185]. hydrological [103]. hydropower [7]. hyper [381]. hyper-heuristic [381]. hyperchaotic [39, 94, 464]. hypercube [273]. hypergraph [627]. hyperledger [376, 707]. hyperspectral [193, 327, 519]. hypothesis [133].

I/O [630]. ICAT [191]. ICAT-net [191]. ICN [445]. identification [1, 56, 220, 370, 389, 525, 559, 583]. Identifying [263, 513]. identity [700, 707]. IDS [254, 428, 527]. II [575]. IID [63, 644]. IIoT [366]. image [3, 17, 26, 70, 90, 96, 100, 105, 108, 113, 138, 160, 168, 169, 193, 240, 244, 275, 338, 343, 345, 353, 374, 383, 402, 404, 440, 464, 490, 500, 565, 604, 616, 668]. image-set [404]. image-text [138, 374]. imagery [202, 303, 391]. images [22, 113, 267, 419, 515, 618, 645, 648, 689, 711]. imaging [432, 546, 565, 616]. imbalance [16, 530]. imbalanced [248, 452]. imitation [289]. Impact [183, 187, 265, 513, 662]. implementation [405, 506, 580, 639]. Implicit [102, 692]. Improve [57, 156, 408, 661]. Improved [34, 44, 66, 69, 116, 132, 177, 188, 230, 238, 251, 261, 266, 274–276, 290, 295, 330, 363, 389, 412, 443, 464, 481, 493, 496, 502, 507, 515, 571, 578, 595, 620, 638, 657, 701]. improvement [306]. Improving [157, 262, 444, 623]. in-context [607]. in-memory [190]. In-situ [225]. InceptionNet [353]. incorporating [52, 539]. increasing [2]. incremental [81]. independent [341, 542]. index [717]. indexing [328]. indices [250, 331, 489]. indispensable [207]. indistinguishability [376]. individual [355]. induced [43, 145, 585]. inductive [364]. industrial [84, 366, 424, 686]. InEPS [436]. infant [106]. Inference [60, 73, 183, 205, 209, 365, 506, 557, 602]. influence [21, 130, 569]. influential [513]. information [43, 56, 158, 255, 285, 364, 394, 477, 549, 572, 647, 676, 704]. informed [92, 401]. infrared [22, 56, 220]. infrastructure [455, 472, 574, 589]. infrastructures [15]. inheritance [158, 590]. initialization [710]. injection [544]. Innovations [347]. Innovative [20, 72, 386, 538]. Inqasm [451]. InQuIR [451]. insider [246]. insole [660]. inspection [561]. inspired [107, 669]. Instruction [405, 672]. Integrated [95, 198, 201, 300, 346, 544]. integrating [386]. integration [358, 456, 673]. integrity [534]. Intelligence [62, 270]. Intelligent [45, 76, 237, 360, 369, 411, 436, 472, 606, 651]. intensity [54]. intent [319, 690]. intent-based [690]. inter [283, 611]. inter-bot [611]. inter-GPU [283]. interaction [560, 680, 704]. interactions [698]. Interactive [189]. interconnected [112]. interconnection [537, 677]. interdependencies [673]. Interest [122, 146, 221]. interference [104]. interictal [1]. intermediate [427]. internally [211]. internet [84, 123, 271, 301, 339, 559, 613, 624, 667]. interpretable [623]. interruption [201]. intersample [484]. intersections [29]. intra [374]. intra-modal [374]. Introducing [295, 478]. introspection [260]. introspective [696]. intrusion [16, 42, 110, 231, 254, 277, 428, 527, 553, 650, 659, 713]. intrusions [646]. invariant [306]. invasive [674]. Inverse [159, 702]. Investigating [187]. Investigation [709]. IoE [617]. IoMT [219, 271, 650, 700]. IoT [6, 42, 61, 110, 133, 141, 149, 151, 195, 209, 231, 232, 246, 272, 366, 367, 369, 390, 399, 444, 455, 473, 541, 557, 611, 642, 665]. IoT-based [6, 209, 444]. IoT-cloud [141]. IoT-enabled [195, 455]. IPAQ [384]. irregular [98, 500].

IRS [589]. IRS-assisted [589]. isogeny [478]. isogeny-based [478]. itemsets [578]. iterated [182]. iteration [186]. iterative [532]. iterative-partitioning [532]. iterator [350].

Jackal [545]. Jacobsthal [197]. Japanese [631]. Jaya [443]. jitter [470]. job [293, 368, 488, 620]. job-shop [293]. jobs [368]. Joint [84, 469, 589]. jointly [46]. Josephus [180, 341]. Journal [308– 314, 406, 407, 458, 516, 517, 568, 619, 670, 721]. judgment [429]. juggling [682].

K-NN [66]. KdV [709]. Kepler [66].

kernel [24, 70, 684]. kernel-based [24]. Key [6, 9, 57, 149, 225, 259, 350, 378, 389, 409, 581, 678]. key-derivation [225]. key-value [57, 259]. keyframe [417]. keyphrase [693]. knowledge [33, 131, 158, 258, 416, 521, 590, 694]. Korean [163]. KSIPF [532]. KubePipe [461]. KV [156].

label [122, 607, 622]. label-aware [122]. labeled [717]. Lagrange [434]. landmarks [490]. landscape [354, 357]. language [50, 75, 163, 194, 258, 359, 506, 542]. Large [41, 50, 70, 91, 104, 145, 163, 194, 506, 526, 574, 578]. Large-scale [41, 91, 104, 163, 526, 574, 578]. latency [366]. latent [118, 168]. latent-content [168]. lattice [219]. lattice-based [219]. layer [31, 329, 389, 518, 557]. layout [340]. Ldstd [423]. leader [488]. leadership [630]. leakage [165]. leakage-resilient [165]. learn [207]. Learnable [585]. learned [328]. Learning [1, 9, 20, 33, 35, 37, 58, 63, 69, 72, 74, 92, 110, 113, 114, 130, 137, 139, 148, 153, 158, 161, 176, 178, 184, 185, 188, 196, 200, 207, 220, 223, 232, 241, 246, 254, 256, 263, 265, 269, 271, 284, 286, 288, 289, 293, 295, 315, 321, 322, 324, 327, 332, 335, 339, 344, 368, 377, 380, 385, 393, 395, 396, 401, 419, 428, 429, 431, 432, 447, 457, 461, 472, 481, 485, 488, 495, 500, 526, 535, 557, 567, 569, 573, 577, 588, 596, 602, 607, 609– 611, 613, 627, 628, 642, 644, 648, 653, 658, 674, 680, 683, 690, 691, 713]. learning [213, 333]. learning- [628]. learning-based [35, 113, 184, 241, 485, 495, 628]. learning-driven [569]. least [210]. legacy [494]. legal [429]. Leipnik [152]. length [666]. length-representation-based [666]. lesion [334]. leukemia [419]. level [46, 125, 183, 189, 200, 256, 325, 353, 381, 418, 453, 461, 541, 612, 623, 673]. leveling [185]. Levenberg [55]. Leveraging [1, 223, 272, 390, 435, 617, 645, 687, 713]. LEVYEFO [441]. LEVYEFO-WTMTOA [441]. LGASR [168]. library [283]. LiDAR [519, 656]. LIF [405]. lifetime [2]. light [27, 160, 275, 340]. LightUAV [105]. LightUAV-YOLO [105]. Lightweight [56, 105, 191, 199, 207, 224, 233, 261, 301, 305, 307, 372, 399, 503, 539, 558, 571, 593, 622]. like [48, 359]. limitations [643]. line [561]. linear [136, 218, 325, 363, 434, 681]. Linguistic [103, 360, 388]. link [364, 641]. lite [297, 413]. literature [253]. LiteYOLO [372]. LiteYOLO-GHG [372]. LM [55]. load [49, 52, 55, 257, 335, 479, 657]. local [132, 158, 182, 302, 333, 673]. locality [404, 682]. localization [1, 140, 170, 195, 249, 318, 481]. Location [85, 128, 221, 376, 584]. lock [206, 694]. lock-in [206]. log [382, 457]. log-based [457]. LogCTBL [457]. logic [701]. logistic [684]. LogSD [382]. long [68, 115, 117, 416, 509]. long-tailed [416]. long-term [68, 115, 509]. loop [67]. LoRA [50, 378]. LoRaWAN [378]. loss [16, 33, 217, 357, 652]. lossy [399]. low [6, 160, 325, 353, 381, 423, 521, 592]. low-altitude [423]. low-level [353, 381]. low-light [160]. low-power [6]. low-visibility [592]. LPQAA [233]. LSM [57, 350]. LSM-tree-based [57, 350]. LSPP [165]. LSTM [31, 142, 177, 265, 373, 412, 463, 544, 600, 642]. LSTM-based [142]. LU [87]. lung [338]. LVAST [305]. Lynxsight [715].

M [512, 519]. M-C&M-BL [512]. MAC [190]. MACAE [231]. Machine [1, 20, 69, 78, 110, 184, 217, 237, 254, 260, 263, 271, 321, 339, 377, 401, 428, 431, 461, 485, 526, 610, 628, 674]. machines [479]. maintaining [91]. maintenance [7]. making [29, 360]. malicious [447, 596, 707]. malware [151, 501, 649]. Mamba [235, 459, 695]. Mamba-Enhanced [459]. management [103, 166, 232, 251, 278, 324, 378, 606, 678]. Manifold [404]. manipulators [640]. mantis [714]. manufacturing [135]. Many [80, 124, 449]. Many-objective [80, 124, 449]. map [440, 579, 587]. MAPER [15]. mapping [193, 533, 607, 617]. MAPPO [467]. MapsTSF [695]. MAQT [438]. Maritime [22]. marketing [229]. Markov [89]. Marmara [265]. Marquardt [55]. mash [285]. mash-up [285]. mask [96]. Masked [239, 596]. masking [500]. massive [584]. matching [24, 138, 362, 374, 585]. MATD3 [279]. material [358, 456]. materials [13, 494, 554]. mathematical [201, 347, 633]. mating [449]. MATLAB [489]. matrices [159]. matrix [79, 175, 204, 214, 292, 506, 641]. MAVs [491]. max [47]. max-min [47]. maximization [21, 130, 146, 364, 569]. maximum [363]. Mayfly [633]. MC [153]. MC-CRS [153]. MCAN [299]. MCGDM [173]. Mdcsnet [394]. MDH [19]. MDH-Net [19]. means [65, 274, 448, 532, 712]. measure [204]. measurement [230, 582]. measures [173, 250, 665]. MEC [469]. MECG [326]. mechanism [18, 28, 31, 40, 82, 104, 124, 126, 338, 343, 410, 414, 447, 449, 597, 636]. mechanisms [191, 282, 652]. media [272, 284, 390]. Median [522]. Medical [26, 39, 70, 90, 96, 111, 271, 315, 324, 385, 403, 490, 559, 565, 604, 616]. medium [485]. medium-term [485]. MeghMesa [671]. MEKF [416]. memory [115, 117, 189, 190, 231, 521, 595, 597, 649, 682]. memory-based [595]. memristive [204, 434]. Menger [180]. meniscus [538]. merging [672]. mesoscale [525]. message [364]. Meta [207, 346, 357, 365, 653]. meta-heuristic [346]. meta-learning [207]. Meta-path-guided [365]. metaheuristic [201, 236, 637]. metal [40, 155]. metaphor [477]. Method [2, 11, 62, 83, 103, 113, 221, 238, 249, 255, 262, 264, 337, 357, 358, 377, 389, 392, 420, 434, 440, 453, 456, 486, 497, 500, 501, 520, 532, 535, 571, 575, 576, 583, 594, 610, 631, 634, 654, 656, 683, 701, 706, 709, 720]. methodology [475]. methods [204, 322, 408, 555]. metric [66]. Mfe [525]. Mfe-net [525]. MFFCNN [425]. MFS [622]. MGFormer [676]. MGSF [504]. microscopic [419]. microservices [104]. migration [184]. migrations [15]. Millimeter [270, 656]. millimeter-wave [656]. mills [88]. MIMO [270]. mine [51]. minimal [225]. minimization [306]. minimize [620]. minimizing [184]. Minimum [218, 511, 576]. mining [56, 81, 579]. minority [248, 706]. MIoT [559]. missile [381]. mitigate [183]. Mitigating [104, 455, 552]. mitigation [484, 559]. Mixed [87, 190, 214]. Mixed-precision [87, 214]. mixed-signal [190]. mixer [707]. mixture [13]. Mixup [382]. MLDDoS [453]. MLS [533]. Mobile [9, 35, 119, 123, 128, 257, 286, 442, 467, 653, 687, 707]. mobility [15]. mobility-aware [15]. modal [107, 127, 177, 374, 656, 675]. modality [86, 326]. modality-enhanced [326]. modality-squeeze [86]. model [30, 34, 39, 40, 65, 75, 105, 106, 143, 162, 163, 178, 179, 200, 201, 222, 228, 237, 254, 265, 273, 304, 307, 321, 344, 352, 370, 373, 385, 388, 395, 401, 403, 405, 420, 424, 428, 446, 457, 484, 505, 512, 524, 551, 556, 558, 560, 593, 600, 603, 632, 647, 648, 655, 666, 719, 720]. Modeling [46, 122, 171, 247, 300, 347, 374, 421, 489, 494, 554, 695]. modelling [664]. models [20, 50, 187, 194, 258, 288, 345, 359, 363, 503, 506, 547, 554]. modern [268, 631]. modified [99, 164, 186]. modular [296]. modulation [683]. module [231, 267]. module-assisted [231]. modules [503, 571]. moduli [102]. MOEAD [710]. Mondrian [553]. mongoose [493]. monitoring [260, 339, 444, 523, 626, 674, 707]. monocular [459]. Monomm [459]. Monte [381, 575]. morphological [1]. most [23]. motion [297, 413]. mountain [140]. move [182]. movements [106]. movie [78, 354]. moving [392]. MQTT [533]. MR [648]. MRI [19, 352, 432, 689]. MSBES [251]. MSRA [535]. MSS [635]. MST [86]. MST-ARGCN [86]. MT [38]. MT-3000 [38]. Multi [4, 7, 10, 17, 19, 31, 32, 40, 46, 49, 54, 56, 59, 60, 70, 74, 84, 88, 89, 96, 98, 107, 108, 113, 116, 122, 124, 135, 153, 156, 177, 185, 186, 199, 216, 220, 251, 255, 266, 277, 279, 287, 293, 321, 329, 355, 379, 384, 388, 394, 418, 421, 425, 438, 440, 441, 445, 453, 459, 465, 468, 469, 497, 501, 512, 519, 545, 564, 566, 570, 572, 589, 612, 621, 622, 635, 644–646, 649, 652, 654, 656– 658, 675, 676, 685, 699, 710]. multi- [251]. multi-access [445]. multi-agent [279, 388]. multi-BiLSTM [512]. multi-branch [519]. multi-carrier [589]. multi-cell [469]. multi-channel [465]. multi-class [74, 644]. multi-client [564]. multi-CNN [512]. multi-compactions [156]. multi-contrastive [153]. multi-core [59, 287]. multi-criteria [89]. multi-cylinder [185]. multi-feature [654]. multi-genotype [570]. Multi-granularity [46, 56, 122, 220]. multi-head [31]. multi-hierarchical [621]. multi-information-based [676]. multi-label [622]. multi-layer [329]. Multi-level [418, 453, 612]. multi-mechanism [124]. Multi-medical [96]. Multi-modal [107, 177, 656, 675]. multi-model [321]. Multi-objective [54, 59, 88, 98, 116, 255, 287, 293, 355, 384, 622, 657, 699, 710]. multi-party [379]. multi-person [7]. multi-process [135]. Multi-queue-based [4]. Multi-robot [497]. Multi-scale [10, 17, 32, 40, 49, 70, 199, 277, 394, 425, 438, 440, 459, 519, 645, 646, 652]. multi-scenario [566]. multi-secret [635]. multi-server [84]. multi-sources [649]. multi-stage [19]. multi-step [186]. Multi-strategy [266, 468, 545, 685]. Multi-stream [572]. Multi-task [658]. multi-threshold [113]. Multi-time-scale [421]. multi-TPU [60]. multi-tracker [441]. multi-valued [216]. multi-view [501]. multi-wing [108]. multichannel [200]. multiclass [432]. multicore [548]. multidisk [259]. multilayer [353]. multilayered [700]. multilevel [90, 671]. Multimodal [28, 46, 58, 64, 71, 86, 284, 299, 326, 477, 560, 636, 658, 683]. multinational [442]. multiobjective [636]. multiparty [507]. multiple [36, 39, 72, 83, 249, 279, 298, 416, 553, 691]. multiplication [214]. multiplier [109]. multiscale [525]. multisense [661]. multisensor [661]. multistep [535]. Multivariate [162, 282, 425, 498]. multizone [529]. music [594]. Mutual [242, 255, 301, 364, 696]. MVS [371].

N [694]. N-Lock [694]. named [666]. Namib [412]. nanotechnology [253]. natural [248, 359]. navigation [107, 308–314, 406, 407, 418, 458, 516, 517, 568, 617, 619, 670, 721]. NBO [412]. nearest [23, 79, 647]. negative [79, 138]. negotiation [228]. Neighbor [79, 316]. neighborhood [248, 513, 647]. neighbors [23]. nested [666]. Net [19, 70, 191, 525, 604]. NetQASM [451]. nets [247]. Network [2, 3, 16, 17, 19, 27, 36, 56, 68, 71, 82, 86, 100, 125, 129, 133, 139, 144, 152, 157, 160, 182, 191, 212, 233, 239, 242, 250, 267, 275, 291, 293, 298, 299, 315, 318, 332, 334, 340, 348, 373, 402, 405, 412, 417, 418, 421, 425, 446, 459, 465, 467, 471, 491, 495, 498, 509, 514, 519, 525, 527, 553, 563, 566, 567, 572, 583, 604, 625, 649, 659, 663, 669, 676, 679, 688, 689, 692, 704, 711]. network-assisted [315]. networked [550]. networking [435, 690]. networks [6, 8, 18, 25, 36, 42, 64, 97, 127, 131, 145, 146, 164, 170, 175, 180, 183, 189, 195, 204, 209, 211, 216, 231, 246, 254, 277, 279, 335, 346, 349, 366, 386, 389, 415, 428, 434, 460, 469, 474, 476, 490, 513, 537, 543, 562, 569, 611, 677, 693, 703]. Networks-based [97]. Neural [3, 6, 25, 42, 49, 97, 129, 183, 191, 204, 212, 277, 315, 332, 346, 402, 405, 415, 421, 425, 434, 465, 474, 498, 679]. NeuralWiGait [380]. neurodynamics [159]. neuromorphic [405]. neutron [575]. news [122, 177, 249, 299, 388]. Newton [152]. Next [151, 386]. Next-generation [151, 386]. NLP [177]. NN [66]. no [620]. no-wait [620]. Node [140, 389, 400, 410, 672]. Node-to-set [400]. nodes [57, 513]. noise [530, 532, 547]. NOMA [589]. non [63, 79, 644, 674]. non-IID [63]. non-invasive [674]. non-negative [79]. nonlinear [92, 152]. nonnegative [641]. nonterrestrial [567]. norm [204]. normalization [114]. normalized [652]. NoSQL [664]. NoSQL-based [664]. Note [228, 250]. Novel [6, 27, 65, 67, 69, 83, 94, 128, 166, 179, 219, 262, 271, 278, 280, 322, 323, 435, 444, 512, 533, 637, 661, 677, 717]. NPGPT [359]. NR [336]. NR-unlicensed [336]. nuclei [174, 199]. number [94, 250]. number-based [250]. numbers [576]. numerical [188, 203, 720].

O [630]. object [105, 157, 202, 303, 391, 426, 459, 465, 476, 515, 571, 572, 585, 593, 618, 623, 646, 680, 697, 704, 718]. objective [54, 59, 80, 88, 98, 116, 124, 255, 287, 293, 355, 384, 449, 590, 622, 657, 699, 710]. objects [392, 652]. oblivious [678]. obstacle [617]. OCB [580]. occlusion [370]. OCT [482]. ODD [202, 391]. ODD-YOLOv8 [202]. off [379, 403]. off-chain [379, 403]. offline [667]. offloading [29, 35, 45, 52, 76, 289, 367, 469, 514, 577, 625, 653]. on-chain [379, 403]. oneAPI [437]. Online [111, 286, 328, 470, 535, 613, 667]. online/offline [667]. Onsager [97]. Ontological [698]. Ontologies [243]. ontology [93]. open [426, 475]. open-world [426]. operators [43, 294, 591]. opportunities [291]. Opposition [148]. Opposition-based [148]. optical [581]. optimal [102, 212, 218, 384, 550]. optimisation [557]. Optimization [8, 13, 20, 41, 45, 54, 55, 59, 66, 72, 80, 90, 95, 98, 99, 124, 132, 143, 148, 150, 170, 188, 201, 209, 236, 237, 255, 264, 266, 287, 302, 320, 325, 337, 347, 355, 365, 374, 375, 412, 420, 434, 441, 470, 474, 479, 493, 497, 545, 549, 569, 575, 589, 595, 597, 620, 636, 638, 651, 655, 657, 672, 685, 690, 705, 714]. optimizations [188]. Optimized [6, 47, 191, 323, 340, 366, 398, 432, 438, 452, 650, 674, 702, 719]. optimizer [90, 140, 236, 264, 375, 468, 713]. Optimizing [29, 72, 76, 88, 134, 172, 257, 270, 289, 303, 346, 377, 409, 473, 693]. order [90, 152, 203, 247]. ordinary [49]. ordinates [529]. oriented [82, 200]. orthogonal [125, 212]. orthogonal-gating [125]. orthopair [103]. other [296]. outsourced [225]. Outsourcing [296, 507]. overhead [289, 350]. oversampling [248, 532, 706].

P2PPO [514]. Packing [211]. Padovan [197]. PageRank [186]. pages [320]. pair [358, 456, 530]. Paired [273]. pairing [436]. Pakistan [103]. Pallet [481]. palsy [106]. pancake [245]. pandemic [339]. PANs [26]. Parallel [7, 24, 38, 47, 73, 135, 226, 341, 352, 460, 514, 526, 529, 575, 580, 612, 643, 699, 705]. parallelization [461]. parameter [266, 468, 530]. parameterized [547].

Pareto [98]. parity [708]. ParsBERT [719]. part [634]. part-of-speech-driven [634]. partial [602, 609]. partially [717]. Particle [13, 255, 497]. particles [637]. partition [448]. partitioned [273, 712]. partitioning [532]. parts [106]. party [9, 379]. passing [364]. password [38]. patent [579]. Path [17, 48, 83, 95, 245, 273, 276, 329, 360, 365, 381, 497, 685, 701]. paths [211, 400]. patient [166]. patient-centric [166]. patients [339]. Pattern [81, 325, 682]. patterns [584, 715]. PB [678]. PB-UOKM [678]. PBFT [330, 716]. PCB [261]. PDCF [213]. PDCF-DRL [213]. PDPA [634]. pedestrian [370, 676]. perception [439]. Perceptual [474]. Performance [41, 57, 156, 187, 260, 265, 336, 351, 409, 420, 450, 478, 494, 508, 551, 601, 615, 639, 662, 682]. performance-to-cost [601]. periodic [129, 681, 695]. permutation [443]. Persian [719]. person [7, 56, 220]. Personalized [158]. perspective [682]. perturbed [203]. PESA [487]. Petri [247]. pharmaceutical [561]. phase [191]. phishing [69, 320]. photography [502]. photovoltaic [20, 468]. physical [51, 299, 559]. physics [92]. physics-informed [92]. picking [191]. PID [266]. pinball [217]. pipeline [156]. pipeline-based [156]. pipelines [461]. pivoting [87]. pixel [230]. placement [192, 257, 302, 444]. placer [268]. planar [316]. Planning [7, 83, 88, 95, 276, 360, 381, 497, 685, 701]. plants [401]. plaque [267]. platform [100, 219]. platforms [121, 550, 710]. play [75, 503]. plug [75, 503]. plug-and-play [75, 503]. PMCKV [156]. Pocket [235]. point [176, 221, 394, 487, 666, 715]. Polaris [405]. policy [107, 343, 365, 483, 678]. policy-based [678]. pollution [275]. polynomial [179, 531]. polynomials [252]. pooling [176, 240, 267]. popularity [552]. population [595, 637]. portability [639]. pose [438, 481]. position [348]. positional [282]. positioning [230, 582]. post [9, 233]. post-quantum [9, 233]. potentials [126, 414]. power [6, 20, 190, 336, 401, 452]. power-constrained [190]. powered [565, 590, 616]. PPLBB [219]. PPO [514]. PPTopicPLM [75]. practical [678, 716]. practice [26]. pre [75, 87, 111]. pre-diagnosis [111]. pre-pivoting [87]. pre-trained [75]. Precise [230, 604]. Precision [87, 109, 214, 249, 277]. predictability [548]. predicting [321]. prediction [18, 54, 92, 129, 162, 178, 198, 241, 297, 323, 355, 364, 373, 381, 401, 411–413, 429, 431, 485, 499, 641, 656, 669, 676, 679, 695, 703]. predictive [332, 554]. preference [27]. prefilling [77]. Preliminary [599]. Prescribed [640]. preserving [14, 111, 127, 128, 219, 246, 404, 430]. pressure [674]. pretraining [187]. prevention [611]. price [373, 498]. Pricing [278, 467]. principle [124]. principles [291]. printed [631]. prior [334]. PRIORITI [342]. prioritisation [97]. prioritization [342]. prioritized [514]. priority [52]. priority-aware [52]. Privacy [14, 85, 111, 127, 128, 134, 219, 246, 333, 376, 430, 507, 540, 567, 707]. privacy-preserving [14, 111, 127, 128, 219, 246, 430]. PRNU [25]. PRNU-based [25]. pro [47]. probabilistic [376]. probability [355]. probability-driven [355]. Problem [102, 161, 234, 400, 444, 488, 511, 569, 620]. problems [80, 132, 148, 150, 203, 236, 251, 375, 493]. procedure [294]. process [135]. processes [332]. processing [190, 271, 405, 411, 581]. processor [47]. processors [214, 371]. product [359, 424]. product-like [359]. production [88, 561]. profiling [649]. profit [84]. programmable [586]. programming [363, 643]. programs [487]. progressive [560]. project [323]. projection [404]. projects [263]. prompt [429, 634, 643, 658]. Prompt4LJP [429]. proof [97, 694]. proof-of-reputation [97]. propagation [77]. Proposal [256, 548]. Proposal-level [256]. Proposed [631, 708]. protection [85, 96, 333, 376, 626, 707]. protocol [9, 61, 149, 435, 635, 694]. protocols [317, 379]. prototype [431]. provably [667]. provider [278]. Province [358, 456]. provision [537]. proximal [197]. pseudo [94]. pseudo-random [94]. PSI [622]. PSI-MFS [622]. publication [143]. Publisher [287, 616]. pump [472]. purification [422, 476]. PV [401]. PVGwfa [612]. pyelonephritis [489].

qARM [578]. QbE [24]. QEKLR [684]. QENN [679]. QLW [224]. QML [431]. QNN [354]. QP [474]. QSPR [489]. QSVM [78]. qualitative [392]. Quality [144, 537, 673]. Quality-of-service [537]. quantisation [557]. quantized [506]. Quantum [9, 78, 117, 142, 147, 154, 223, 233, 253, 315, 361, 395, 398, 427, 431, 460, 465, 508, 547, 615, 679, 684, 686]. quantum-classical [465]. quantum-enhanced [679, 684]. Quantum-resistant [508]. quantum-SAR [223]. quasi [204]. quasi-synchronization [204]. quaternion [204, 434]. quaternion-valued [204, 434]. quaternions [681]. query [350, 438]. query-optimized [438]. queue [4, 201].

rabbit [320]. rabbits [150]. race [643]. radar [30]. radial [152]. radiation [475]. radius [120, 466]. railways [646]. ramification [713]. random [7, 94, 193, 331, 396, 488, 496, 555]. RandomForest [162]. range [140, 350, 576]. range-free-based [140]. rank [43]. rank-aggregation [43]. rate [265]. rating [410]. ratio [241, 601]. ratios [13, 452]. RAVDESS [31]. raw [13]. RBC [635]. RBC-MSS [635]. RBF [212]. re [56, 220, 370]. re-identification [56, 220, 370]. reaction [203]. reaction-diffusion [203]. read [57, 409]. real [3, 118, 121, 222, 226, 229, 260, 307, 344, 351, 459, 470, 482, 524, 674, 686]. real-time [222, 226, 229, 260, 307, 344, 351, 459, 470, 482, 524, 674]. real-world [686]. reasoning [50, 112, 205]. rechargeable [298]. recognition [10, 12, 31, 37, 275, 349, 353, 380, 416, 417, 442, 463, 524, 546, 558, 573, 609, 660, 666]. recognizers [631]. recommendation [79, 122, 208, 221, 234, 304, 354, 421, 446, 552, 594, 627]. recommendations [27, 78, 292]. recommender [153, 215, 227, 272, 390]. recommending [698]. reconfigurable [369, 548, 586]. reconfiguration [135, 591, 694]. Reconstruction [37, 62, 168, 482, 687]. Reconstruction-driven [37]. records [324, 403]. recovery [38, 550]. recurrent [86, 332, 421]. recursive [175]. redefining [354]. reduced [366]. reducing [206, 350, 538]. reduction [167, 526]. redundancy [183, 238]. Reference [80, 240, 360, 682]. Reference-based [240]. referring [174]. refinement [19, 215]. reflection [345]. regional [483]. regional-contention-driven [483]. regions [103]. registers [599]. registration [19, 70, 490]. registry [15]. regression [387, 684]. regularization [55, 327, 641]. reinforcement [35, 113, 130, 161, 184, 185, 188, 213, 286, 293, 295, 335, 488, 495, 535, 577, 609, 653]. related [126, 414]. relation [693]. relation-aware [693]. Relational [112, 364]. relationship [27, 486]. relationship-aware [27]. relay [77]. released [694]. relevance [238]. reliability [180, 205, 444, 562]. reliable [237, 256, 655]. ReLU [557]. remote [17, 303, 339, 377, 678, 711]. removal [345]. replay [514]. representation [58, 197, 353, 383, 393, 584, 666]. representations [326, 402, 427]. reputation [97, 330]. RepVGG [442]. requirements [570]. rescaled [217]. Research [28, 221, 228, 241, 306, 358, 370, 456, 552, 565, 616, 660]. reserve [358, 456]. reservoir [300]. reshaping [615]. residual [131, 509, 514, 576]. residuals [240, 522]. Resilience [51, 365, 550]. resilient [109, 165]. resistance [470]. resistant [508]. resisting [252]. resolution [17, 240]. Resource [135, 192, 278, 292, 340, 367, 398, 469, 473, 514, 570, 571, 629, 650, 705]. resource-aware [705]. resource-constrained [571, 650]. response [92, 206]. restricted [576]. Rethink [669]. Retraction [228, 250]. retrieval [127, 258, 343, 353, 549]. reuse [17]. reversed [125]. reversible [210]. Review [253, 427, 705]. reviews [624]. reward [343, 381]. Reynolds [669]. RFAConv [155]. RFAConv-CBM-ViT [155]. RFD [134]. RFD-based [134]. RGB [242, 572]. RGB-D [242, 572]. Riesz [720]. RIME [375]. ring [9, 14, 117]. RIOKV [350]. RISC [405]. RISC-V [405]. risk [698]. risks [275, 455]. RL [495]. RL-EAR [495]. RNN [184]. road [387, 439]. Roaen [125]. RoboSoccer [617]. Robot [95, 497]. robotic [640]. robots [687]. Robust [25, 84, 185, 244, 271, 415]. robustness [530]. Rock [431]. RocksDB [57]. root [195]. rotation [682]. rough [647]. roulette [99]. route [651]. route-based [651]. routing [8, 274, 352, 386, 435, 495, 541]. RSA [102]. RT [290, 307]. RT-DETRmg [307]. RTAB [587]. Rule [205, 243, 411]. rule-based [411]. rules [669]. rumor [75, 486]. rung [103]. RV [405]. RV-SNN [405]. RVV [172].

SA [710]. sacroiliitis [648]. safety [205, 288]. salesman [161]. salient [106, 704]. SALSTM [115]. salted [534]. sample [138, 289]. sampling [68, 145, 394, 486]. sampling-based [68]. sand [113, 168, 325]. sand-dust [168]. sandbox [151]. Sanitation [596]. SAPFIS [73]. SAR [34, 223]. sarcasm [560]. SatEC [514]. satellite [233]. saving [661]. SCA [373]. Scalable [259, 461, 524, 554]. scale [10, 17, 32, 40, 41, 49, 70, 91, 104, 163, 199, 277, 394, 421, 425, 438, 440, 459, 494, 519, 526, 574, 578, 645, 646, 652]. scaled [55]. scales [434]. scaling [640]. scenario [566]. scenarios [697]. Scene [107, 138, 515, 626, 687, 696]. scenes [67]. SCG [55]. scheduler [436]. schedules [298]. scheduling [4, 116, 226, 293, 295, 368, 384, 443, 488, 601, 617, 629, 633, 657, 699, 710]. scheme [14, 85, 96, 119, 203, 213, 233, 289, 301, 376, 378, 379, 439, 464, 531, 551, 667, 678, 700]. schemes [564, 665]. scientometric [624]. SCNet [566]. SCNet-YOLO [566]. scoring [104, 342, 716]. scoring-based [716]. script [69]. script-based [69]. SD [291]. SD-WAN [291]. SDHNet [68]. SDN [386, 445]. SDSMS [378]. SDSMS-LoRa [378]. Sea [265]. sealed [392]. search [124, 212, 216, 251, 302, 328, 511, 578, 595]. search-based [212, 302]. searchable [119]. secret [635]. Secure [9, 12, 94, 271, 296, 378, 534, 564, 667, 678]. Securing [181, 246]. security [69, 72, 108, 121, 165, 317, 366, 369, 471, 533, 567, 598, 650, 665]. segmentation [60, 74, 90, 113, 157, 174, 176, 199, 235, 242, 267, 338, 419, 604, 654, 689, 692]. segmented [115]. select [283]. selectable [39]. selected [500]. selection [2, 89, 134, 147, 150, 212, 238, 254, 255, 323, 358, 377, 428, 449, 456, 544, 577, 622, 636, 647, 650, 672, 717]. selections [605]. Selective [133]. selector [499]. Self [28, 56, 58, 115, 121, 137, 267, 343, 345, 397, 484, 491, 500, 573, 645, 647]. self-adapting [137]. self-adaptive [343, 491]. self-attention [28, 115, 267, 484, 573]. Self-aware [121]. self-distillation [56]. self-expressiveness [397]. self-forgery [645]. self-information-based [647]. self-supervised [58, 345, 500]. Semantic [64, 74, 242, 299, 353, 382, 520, 541, 585, 668, 687]. Semantic-aware [687]. semantic-physical [299]. Semi [169, 210, 269, 383]. semi-reversible [210]. Semi-supervised [169, 269, 383]. sensing [17, 303, 377, 707, 711]. sensitive [577]. sensitivity [487]. sensor [2, 8, 133, 170, 209, 298, 349, 495]. sensors [410, 546]. sentiment [46, 58, 64, 86, 125, 131, 189, 200, 348, 543, 634, 658, 673, 688, 719]. SeqMatcher [362]. sequence [7, 197, 362, 612, 682]. sequence-to-graph [612]. sequencing [88]. sequential [81, 421]. series [68, 114, 129, 179, 282, 425, 503, 509]. server [84, 257, 450, 470]. serverless [206]. servers [296]. service [84, 89, 201, 278, 445, 453, 537, 574, 625]. services [690]. session [292, 304, 378]. session-based [292, 304]. set [225, 304, 358, 360, 400, 404, 456, 511, 647]. sets [173, 578]. settings [297, 413]. SFC [535]. SFOD [256]. SFVE [477]. SGB [523]. SGB-YOLOv5 [523]. SGSLNet [176]. shallow [639]. Shannon [65]. shaped [3]. shard [694]. sharding [420]. shared [712]. sharing [368, 385, 403, 439, 599, 635, 678]. sharpness [306]. sharpness-aware [306]. SHAVER [599]. shifts [500]. ship [34, 566]. shop [293, 443, 488, 620]. short [55, 75, 115, 117, 350]. short-range [350]. short-term [55, 115, 117]. short-text [75]. shortest [329]. shot [139, 596, 668]. siamese [491, 583]. Siamsdt [491]. side [69, 252, 611]. sidelobe [325]. sign [542, 632]. signal [190]. signature [14, 44, 478, 531, 534, 621, 667]. signature* [478]. signer [542]. signer-independent [542]. signers [542]. significant [23, 210]. signing [379]. signs [307]. SIKE [473]. Sim [539]. Sim-ConvFormer [539]. SimAM [539]. SIMD [405]. SIMD-style [405]. Similar [594]. similarity [124, 150, 468, 605]. simplified [648]. Simulation [132, 201, 247, 300, 460, 586]. simulator [494]. simulators [528]. Simultaneous [95].

Single [169, 345, 590, 663, 666, 708]. single-cell [663]. Single-image [169, 345]. single-objective [590]. single-point [666]. Singularity [454]. singularly [203]. site [358, 456]. situ [225]. situation [73, 471]. size [662]. size-dependent [662]. skeleton [10, 37]. skeleton-based [10, 37]. sketch [21, 453]. sketch-based [21]. skin [334, 563]. skyline [216]. SLAM [587, 687]. slice [509]. sliding [328]. slopes [431]. small [202, 303, 307, 315, 391, 423, 424, 515, 618]. Smart [44, 97, 297, 379, 412, 413, 455, 518, 520, 600, 660, 714]. smartphone [272, 390]. smoke [626]. Smooth [217]. snake [638]. SNN [405]. SoC [491]. Social [27, 145, 146, 208, 215, 227, 272, 284, 390, 446, 569, 627]. social-curiosity-based [227]. social-media [284]. soft [294, 540]. soft-biometric [540]. SoftMax [33]. software [291, 435, 495, 499, 610]. software-defined [291, 435, 495]. solar [20, 468]. solution [151, 444, 574]. solutions [617]. solve [161, 720]. solver [602]. solving [102, 132, 236]. sort [164]. sound [288]. source [25, 323, 611]. source-side [611]. sources [649]. space [636, 668, 720]. spanning [341]. Spark [518, 526]. sparse [175, 476, 695]. spatial [169, 281, 394, 519, 522, 584, 604, 715]. spatial-channel [169, 604]. spatial-spectral [519]. spatial-temporal [281]. spatially [500]. Spatio [18, 129, 396, 703]. Spatio-temporal [18, 129, 396, 703]. Spatiotemporal [10, 332, 695, 704]. special [636]. specific [93, 218]. specifications [321]. Spectral [120, 466, 519]. spectrogram [392]. spectrum [14, 707]. speech [28, 573, 634]. speed [18, 241, 492, 576]. spherical [173, 595]. spiking [405]. spiral [39]. spirals [633]. spiro [331]. split [63, 196]. splitting [186, 203, 477]. Spotify [594]. Sql [544]. squaring [398]. squeeze [86]. SS [303]. SSA [212]. SSCENet [519]. SSD [528].

SSIM [500]. SSO [123]. SSO-based [123]. stability [434]. stacked [178]. Stackelberg [228]. stage [19, 62, 101, 440, 486, 524, 653, 680]. star [562, 677]. start [234]. state [185, 479]. Static [198, 546]. station [7]. Statistical [373]. STD [24]. STDP [405]. steepest [148]. Steiner [211]. step [186]. STGEN [703]. stochastic [329, 526]. Stock [373]. Storage [57, 328, 549, 664, 678]. storage-efficient [328]. store [259, 409]. stores [156, 350]. storytellers [263]. STPNet [332]. strategies [45, 204, 402, 467, 710]. strategy [4, 8, 54, 58, 80, 87, 125, 134, 138, 168, 170, 251, 266, 333, 468, 493, 510, 545, 577, 591, 614, 638, 685]. stratified [176]. straw [523]. stream [19, 68, 139, 318, 572, 591, 613]. streaming [437, 601, 671]. streamlines [300]. strengthened [449]. strengths [643]. strip [88, 711]. strongly [180]. structural [92, 176, 389, 431]. Structure [132, 199, 536]. structured [637]. study [152, 265, 292, 315, 333, 358, 410, 454, 456, 554, 618]. style [305, 405]. sub [230, 279]. sub-networks [279]. sub-pixel [230]. subgraph [248]. subspace [327, 397, 505]. substructure [536]. successful [263]. sum [316]. summarization [171]. Sunway [575]. super [17, 240]. super-resolution [17, 240]. supercomputer [575]. supercomputing [615]. supervised [58, 74, 169, 269, 318, 345, 383, 395, 500]. supervoxel [654]. support [78, 217, 229]. supported [192]. surface [40, 155]. surfaces [369, 424]. surrogate [300]. surveillance [297, 413]. survey [141, 291, 317, 528, 555, 628, 665]. sustainable [123]. SVM [271]. SW26010 [47]. SW26010-pro [47]. swap [400, 536]. Swarm [13, 95, 255, 325, 497]. swarms [621]. sweep [83]. swin [689]. switch [470]. Swpmmas [47]. SYCL [462, 639]. symmetric [510, 566]. synchronization [6, 204]. synchronous [526]. Syntactic [189, 374, 673]. Syntactic-guided [374]. syntax [200]. synthesis [325]. Synthetic [248]. system [14, 38, 42, 47, 51, 73, 94, 108, 110, 141, 152, 153, 156, 185, 209, 229, 241, 259, 272, 278, 289, 297, 337, 339, 344, 351, 369, 380, 390, 413, 430, 442, 463, 464, 479, 523, 527, 529, 559, 561, 626, 630, 633, 650, 651, 656, 682, 713]. Systematic [253]. systems [39, 45, 57, 112, 149, 192, 197, 215, 227, 254, 283, 328, 356, 428, 444, 450, 483, 525, 591, 664, 699]. SZ4IoT [399].

T [385, 596]. T-BFL [385]. T-Sanitation [596]. table [210]. table-based [210]. tabu [511]. tackling [452]. tags [598]. tailed [416]. tardiness [620]. Target [22, 95, 422, 423, 502, 592]. Task [11, 29, 35, 52, 76, 141, 192, 279, 289, 295, 384, 514, 570, 581, 617, 629, 633, 653, 657, 658]. tasks [4, 99, 121, 226, 577]. TC [163]. TC-BERT [163]. TCP [484]. TCP-based [484]. tears [538]. technique [248, 367, 424, 472, 677]. techniques [72, 247, 284, 347, 349, 567, 606, 611]. technologies [462, 581, 686]. technology [163, 167, 579, 660]. temperature [162, 401]. template [491]. temporal [18, 71, 129, 281, 318, 396, 498, 641, 703]. Tensor [327, 361, 460]. term [55, 68, 115, 117, 485, 509]. terminal [625]. terms [720]. ternary [190, 581]. terrestrial [567]. terrestrial-nonterrestrial [567]. test [610]. test-generation [610]. testing [133]. tethered [550]. text [75, 101, 138, 171, 374, 498, 510]. text-based [498]. texts [719]. textual [46, 393, 688]. texture [351]. their [53, 106, 173, 294, 584]. Theoretical [682]. Theory [228, 410, 412, 497, 556, 603]. therapeutics [5, 698]. Thermal [546, 565, 616]. thermoplastic [13]. things [123, 271, 301, 339, 559, 624, 667]. thoughts [50]. threat [142, 342, 642]. threats [246, 559]. Three [9, 139, 685]. three-dimensional [685]. Three-stream [139]. threshold [90, 113]. thresholds [435]. throughput [405]. Time [68, 114, 116, 121, 128, 129, 159, 179, 203, 206, 222, 226, 229, 260, 262, 282, 307, 344, 351, 384, 392, 396, 421, 425, 434, 459, 470, 482, 509, 524, 548, 608, 640, 674, 709, 720]. time-aware [384]. time-dependent [203]. Time-frequency [262, 392, 608]. time-series [282, 509]. time-slice [509]. time-space [720]. timed [247]. timing [268]. timing-driven [268]. Tiny [719]. Tiny-ParsBERT [719]. TLNN [96]. TMTC [675]. token [74, 417]. Tolerance [716]. tolerant [577, 640, 677]. tool [461, 487]. Topic [75, 171, 382, 579]. topic-enhanced [75]. topological [250, 331, 489]. topology [10]. TOPSIS [85, 103, 173, 358, 456]. torus [48]. torus-like [48]. total [316, 620]. touch [690]. Tp [558]. Tp-yolov8 [558]. TPU [60]. traceable [14]. track [504]. tracker [441]. tracking [71, 344, 491, 687]. trading [14]. Traffic [16, 18, 129, 198, 281, 307, 422, 558, 596, 611, 626, 632, 649, 695, 697, 703]. Train [504, 631]. trained [75]. training [101, 212, 323, 554, 602]. trajectory [143, 669, 676]. transaction [97, 694]. transaction-released [694]. transactions [77]. transfer [80, 305, 355, 481, 688]. transferability [623]. transform [39, 425, 718]. transform-enhanced [718]. transformation [662]. Transformer [3, 19, 86, 144, 155, 161, 191, 221, 262, 280, 281, 304, 305, 324, 357, 372, 417, 438, 484, 510, 675, 676, 689, 695]. transformer-based [357]. transformer-facilitated [304]. transformers [282, 356, 538]. translation [542]. transmission [241]. transport [455, 575]. transportation [45, 192, 651]. traveling [161]. Treating [243]. treatment [393]. tree [50, 57, 225, 350, 527]. trees [341]. tri [658]. tri-encoder [658]. triage [294].

Triangle [145]. Triangle-induced [145]. triangular [250]. trick [617]. trigonometric [685]. trip [292]. triple [486]. triple-relationship [486]. Triplet [82, 634]. truck [626]. truncated [269]. Trust [232, 385, 410, 655]. trust-driven [655]. trust-rating [410]. trusted [675]. trustworthy [554]. TS [710]. TS-guided [710]. TSCANet [139, 318]. TSESRec [304]. tubular [692]. tubular-aware [692]. tumor [235, 352, 512, 689]. tuning [50, 266, 600]. tuple [103]. turbines [32]. TVM [172, 214]. two [101, 136, 296, 318, 385, 389, 524, 613, 653, 680]. two-dimensional [136, 385, 613]. two-layer [389]. two-stage [101, 524, 653, 680]. two-stream [318]. type [53, 651]. types [399, 433].

U [3, 70]. U-Net [70]. U-shaped [3]. UAV [192, 202, 391, 502, 570, 685, 718]. UAVs [11, 83]. ultra [505]. ultra-high-dimensional [505]. ultrasound [267]. unbalanced [102, 326]. unbiased [696]. uncertainty [322]. Underlying [93]. understanding [112]. underwater [157, 244, 592, 652]. UNet [338, 689, 711]. unified [281, 641]. unique [160]. unlearning [702]. unlicensed [336]. Unlocking [452]. unmanned [105, 276, 302]. Unpaired [48]. Unsupervised [37, 588, 668]. unveiled [10]. Unveiling [20, 649]. UOKM [678]. updatable [678]. update [150, 225]. upon [386]. Urban [221, 223, 297, 413, 651]. URL [447]. use [686]. User [27, 72, 89, 93, 122, 335, 336, 376, 552]. user-specific [93]. user-to-multiple [72]. users [194]. Using [5, 23, 24, 33, 36, 43, 49, 77, 78, 88, 89, 100, 106, 113, 116, 117, 140, 177, 179, 181, 182, 209, 210, 230, 247, 257, 284, 285, 292, 321, 328, 335, 338, 349, 353, 354, 361, 371, 380, 397, 408, 410, 411, 419, 431, 432, 447, 453, 460, 465, 467, 484, 489, 507, 522, 580, 583, 586, 594, 610, 611, 637, 640, 648, 649, 651, 652, 658, 679, 699, 707, 709, 710, 714]. USV [701]. utility [81, 584]. utilization [289, 292]. utilizing [533]. uTransformer [281]. UWSoS [365].

V [405]. v1.1 [378]. VAE [16, 505]. VAES [580]. value [57, 259, 350, 409]. valued [204, 216, 434]. values [331]. VANETs [435]. variability [629]. variable [123, 152, 170]. variable-order [152]. variance [526]. variances [331]. variants [186]. Variational [5, 16, 97, 117, 154, 353]. various [317]. VASP [494]. VBATS [614]. vector [62, 78, 98, 217, 599]. vectorization [437]. vectors [80]. Vehicle [29, 105, 241, 275, 290, 302, 321, 386, 439, 485, 589, 669]. vehicle-road [439]. vehicle-to-infrastructure [589]. vehicles [276, 613]. vehicular [76, 577]. velocity [170, 435]. velocity-aware [435]. VEM [713]. vendor [206]. ventilation [51]. veracity [285]. verifiable [296, 430, 496]. verification [534, 621]. vertical [702]. vertically [712]. VFCkM [712]. VGG [714]. VGG-16 [714]. via [16, 56, 63, 131, 167, 180, 293, 345, 382, 393, 416, 462, 498, 588, 623, 643, 647, 695]. vibrant [297, 413]. vibrating [197]. video [144, 297, 413, 417, 534, 675, 704]. view [501]. viewpoint [682]. views [91]. virtual [142, 184, 237, 260, 479, 586, 637]. visibility [592]. visible [22, 56, 220, 340]. visible-infrared [56, 220]. Vision [155, 172, 262, 280, 305, 356, 357, 538, 561, 582, 656]. vision-based [561]. vista [297, 413]. vista-lite [297, 413]. visual [46, 107, 416, 418, 477, 491, 587, 683, 687, 688]. visual-textual [46, 688]. ViT [155, 370]. VITS [699]. VLIW [172, 672]. VNDN [651]. VolRec [482]. volumetric [482]. voxel [476]. VPSNet [476]. vulnerability [520]. VVC [474].

wait [620]. walk [555]. walk-based [555]. walks [396]. WAN [291]. warning [401]. Wasserstein [652]. water [195, 251, 639]. watershed [103]. Wave [270, 656]. WAVECAP [433]. wavelet [433, 718]. wavelet-capsule [433]. WBSN [661]. weakly [74, 318, 505]. weakly-supervised [318]. weather [515]. web [320]. weight [61, 85, 392, 426]. weighted [238, 620]. weighting [348, 706]. well [300]. well-based [300]. Wetland [524]. whale [8, 209, 714]. while [408]. White [236, 419]. White-faced [236]. who [263]. wide [6, 291]. Wiedemann [533]. WiFi [380]. WiFi-based [380]. wild [188]. wind [32]. window [213, 328]. wing [108]. wire [222]. wirelength [218]. wireless [2, 8, 170, 209, 298, 495]. wise [16, 145]. within [392]. WK [175]. WK-recursive [175]. WOA [714]. WOA-FMO [714]. words [382]. workflow [371]. workflows [206, 475]. world [426, 686]. worn [349]. WRF [454]. written [388]. WSN [140, 410]. WTDD [32]. WTMTOA [441]. WV [201].

XGBoost [16, 644]. XGBoost-based [16]. XIDINTFL [16]. XIDINTFL-VAE [16].

Yannakakis [507]. yarn [239]. yarn-dyed [239]. YOLO [34, 105, 303, 424, 503, 504, 515, 563, 566, 571, 592, 593, 632, 697]. YOLO-based [571]. YOLO-CE [592]. YOLO-EDGE [697]. YOLO-Fast [593]. YOLO-SS [303]. YOLO11 [592]. YOLOv5 [261, 523]. YOLOv7 [566]. YOLOv8 [32, 34, 202, 275, 372, 391, 515, 558, 587]. YOLOv8-based [372]. YOLOv8-WTDD [32]. YOLOv8s [222]. YOLOv8s-GSW [222]. YOLOv9 [502]. YOLOX [618].

zebra [266].  zero [690, 694].  zero-

knowledge [694]. zero-touch [690].

References

Trigui:2025:AIL

[1] Omar Trigui, Sawsan Daoud, Mohamed Ghorbel, Mariem Dammak, Chokri Mhiri, and Ahmed Ben Hamida. Automated identification and localization of interictal epileptiform discharges: leveraging morphological analysis, five-criterion fulfillment, and machine learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                             URL https://link.springer.com/
                                             article/10.1007/s11227-024-06558-
                                             z.

Alimohammadi:2025:EAC

[2] Edris Alimohammadi, Sajad Haghzad Klidbary, and Mohammad Javadian. Energy-aware clustering method for cluster head selection to increasing lifetime in wireless sensor network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                             tronic). URL https://link.springer.
                                             com/article/10.1007/s11227-024-
                                             06474-2.

Bai:2025:GGA

[3] Xuefei Bai, Yongsong Wan, Weiming Wang, and Bin Zhou. GU-FORMER: a gradient-aware U-shaped transformer neural network for real image denoising. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                             URL https://link.springer.com/
                                             article/10.1007/s11227-024-06576-

Duan:2025:MQB

[4] Lintao Duan and Haiying Wang. Multiqueue-based energy-efficient scheduling strategy for tasks with deadline constraints in cloud data center. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06580-1.

Wang:2025:CBC

[5] Chengling Wang, Yuexin Zhang, Yunru Ma, Peng Chen, and Yang Xiang. CNN-based continuous authentication for digital therapeutics using variational autoencoder. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06490-
2.

Dehghani:2025:NEK

[6] Abbas Dehghani, Sadegh Fadaei, and Resul Das. A novel effective key synchronization approach based on optimized deep neural networks for IoT-based lowpower wide area networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06571-
2.

Bailin:2025:ADR

[7] Li Bailin, Chen Ao, Wu Panqi, Zhang Chao, and Fu Wenlong. Application of discrete random forest algorithm in multi-person asynchronous parallel disassembly sequence planning for hydropower station equipment maintenance. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06540-
9.

Sheng:2025:EER

[8] Hao Sheng, Chen Jun, Cui Jianqun, Fan Xiying, and Li Zhen. An energy-efficient routing algorithm for dual-energy harvesting-assisted wireless sensor networks based on whale optimization strategy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06536-
5.

Kumar:2025:DAP

[9] Uddeshaya Kumar, Manish Garg, and Dharminder Chaudhary. Design and analysis of a post-quantum secure three party authenticated key agreement protocol based on ring learning with error for mobile device. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06467-
1.

Chen:2025:MSS

[10] Hongwei Chen, Jianpeng Wang, and Zexi Chen. Multi-scale spatiotemporal topology unveiled: enhancing skeletonbased action recognition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06531-
w.

Yan:2025:DTA

[11] Shaokun Yan and Yuanqing Xia. A distributed task allocation method for heterogeneous UAVs in dynamic and communication-constrained environments. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06517-
8.

Wang:2025:CBS

[12] Ruyan Wang, Qinglin Zeng, Zhigang Yang, and Puning Zhang. Cloudbased secure human action recognition with fully homomorphic encryption. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06512-
z.

Ozturk:2025:OMR

[13] Ercüment Öztürk, Ayfer Dönmez Çavdar, and Tugrul Çavdar. Optimization of mixture ratios of raw materials in thermoplastic hybrid composites based on particle swarm optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06555-
                                        2.

Yin:2025:CPP

[14] Luona Yin and Huaqun Wang. Conditional privacy-preserving spectrum trading scheme based on traceable ring signature for DSS system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06518-
                                        7.

Temp:2025:MMA

[15] Daniel C. Temp, Alexandre A. F. da Costa, Angelo N. C. Vieira, Ester S. Oribes, Ivan M. Lopes, Paulo Silas S. de Souza, Marcelo C. Luizelli, Arthur F. Lorenzon, and Fábio D. Rossi. MAPER: mobility-aware energy-efficient container registry migrations for edge computing infrastructures. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06516-
                                        9.

Abdulganiyu:2025:XVX

[16] Oluwadamilare Harazeem Abdulganiyu, Taha Ait Tchakoucht, Yakub Kayode Saheed, and Hilali Alaoui Ahmed. XIDINTFL-VAE: XGBoost-based intrusion detection of imbalance network traffic via class-wise focal loss variational autoencoder. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06552-
5.

Xiao:2025:DPF

[17] Huanling Xiao, Xintong Chen, Liuhui Luo, and Cong Lin. A dualpath feature reuse multi-scale network for remote sensing image superresolution. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06569-
w.

Chen:2025:TSP

[18] Hongwei Chen, Hui Han, Yifan Chen, Zexi Chen, Rong Gao, and Xia Li. A traffic speed prediction algorithm for dynamic spatio-temporal graph convolutional networks based on attention mechanism. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06539-
2.

Liu:2025:MNA

[19] Chenou Liu, Kangjian He, Dan Xu, and Hongzhen Shi. MDH-Net: advancing 3D brain MRI registration with multi-stage transformer and dual-stream feature refinement hybrid network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06470-
6.

Zhu:2025:IAS

[20] Chaoyang Zhu, Mengxia Wang, Mengxing Guo, Jinxin Deng, Qipei Du, Wei Wei, and Yunxiang Zhang. Innovative approaches to solar energy forecasting: unveiling the power of hybrid models and machine learning algorithms for photovoltaic power optimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06504-
                                        z.

Gokturk:2025:DDS

[21] Gökhan Göktürk and Kamer Kaya. Di-FuseR: a distributed sketch-based influence maximization algorithm for GPUs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06566-z.

Liu:2025:MTD

[22] Qinxiao Liu, Hangyu Chen, and Fen Zhao. Maritime target detection algorithm based on fusion of visible and infrared images. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06509-
                                        8.

Gorgin:2025:EHA

[23] Saeid Gorgin, Malik Zohaib Nisar, and Jeong-A Lee. Efficient hardware accelerators for k-nearest neighbors classification using most significant digit first arithmetic. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06466-
2.

Gaonkar:2025:PCA

[24] Manisha Naik Gaonkar, Veena Thenkanidiyoor, and Aroor Dinesh Dileep. A parallel computing approach to CNNbased QbE–STD using kernel-based matching. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06497-
9.

Nayerifard:2025:RPB

[25] Tahereh Nayerifard, Haleh Amintoosi, and Abbas Ghaemi Bafghi. A robust PRNU-based source camera attribution with convolutional neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06579-
8.

Makhanov:2025:TGP

[26] Nursultan Makhanov, Ho Duc Nhan, Kok-Seng Wong, and Nguyen Anh Tu. Towards good practice for convolution and attention with PANs in federated medical image classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06476-0.

Zhang:2025:UPS

[27] Hongxia Zhang, Hao Li, Zeya Li, and Pengyu Chen. User preference and social relationship-aware recommendations base on a novel light graph convolutional network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06583-
                                        y.

Mao:2025:RMH

[28] Junjie Mao, Hanxiao Shi, and Xiaojun Li. Research on multimodal hate speech detection based on self-attention mechanism feature fusion. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06602-
                                        y.

Zhang:2025:OVE

[29] Lei Zhang, Miao Wang, Liqiang Wang, Zijian Chen, and Hong Zhang. Optimizing vehicle edge computing task offloading at intersections: a fuzzy decision-making approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06477-
                                        z.

Qi-liang:2025:DFC

[30] Wu Qi-liang, Wang Xing, Zhang Tong, Miao Zi-shu, Ye Wei-liang, and Li Hao. DiffREE: feature-conditioned diffusion model for radar echo extrapolation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06577-
w.

Jin:2025:AER

[31] Zeyu Jin and Wenjiao Zai. Audiovisual emotion recognition based on bi-layer LSTM and multi-head attention mechanism on RAVDESS dataset. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06582-
z.

Yu:2025:YWM

[32] Xiaoyan Yu, Peng Yan, Shaokai Zheng, Qinghan Du, and Daolei Wang. YOLOv8-WTDD: multi-scale defect detection algorithm for wind turbines. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06487-x.

Izadi:2025:NFL

[33] Saadat Izadi and Mahmood Ahmadi. New fusion loss function based on knowledge generation using Gumbel–SoftMax for federated learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06593-
w.

Huang:2025:AYI

[34] Yuqin Huang, Dezhi Han, Bing Han, and Zhongdai Wu. ADV-YOLO: improved SAR ship detection model based on YOLOv8. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06527-
6.

Xie:2025:DRL

[35] Bo Xie and Haixia Cui. Deep reinforcement learning-based dynamical task offloading for mobile edge computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06603-
x.

Yadav:2025:FFA

[36] Rahul Kumar Yadav, Shashi Prakash Tripathi, and Abhay Kumar Rai. Finding future associations in complex networks using multiple network features. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06544-
5.

Liu:2025:RDC

[37] Xing Liu and Bo Gao. Reconstructiondriven contrastive learning for unsupervised skeleton-based human action recognition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06573-
0.

Luo:2025:EHP

[38] Yongtao Luo, Jie Liu, Chunye Gong, and Tun Li. An efficient heterogeneous parallel password recovery system on MT-3000. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06532-
9.

Li:2025:MME

[39] Jianing Li, Wenjing Zhang, and Bing Zhao. 3D medical model encryption based on five-dimensional hyperchaotic systems with 3D Arnold transform and selectable multiple spiral arrangements. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06483-
1.

Zhang:2025:EMM

[40] Heng Zhang, Wei Fu, Xiaoming Wang, Dong Li, Danchen Zhu, and Xingwang Su. An efficient model for metal surface defect detection based on attention mechanism and multiscale feature. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06591-
y.

Wang:2025:POH

[41] Haifeng Wang, Wenkang Guo, and Ming Zhang. Performance optimization of heterogeneous computing for large-scale dynamic graph data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06562-
3.

Lin:2025:GII

[42] Lieqing Lin, Qi Zhong, Jiasheng Qiu, and Zhenyu Liang. E-GRACL: an IoT intrusion detection system based on graph neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06471-
5.

Keyhanipour:2025:GIR

[43] Amir Hosein Keyhanipour. Graphinduced rank-aggregation using information fusion operators. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06595-

Xiangyu:2025:DFG

[44] Wu Xiangyu, Du Xuehui, Yang Qiantao, Liu Aodi, and Wang Wenjuan. Dynamic fine-grained access control for smart contracts based on improved attribute-based signature. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06570-
3.

Zhu:2025:DAC

[45] Sifeng Zhu, Zhaowei Song, Changlong Huang, Hai Zhu, and Rui Qiao. Dependency-aware cache optimization and offloading strategies for intelligent transportation systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06596-
7.

Chen:2025:MGV

[46] Yuzhong Chen, Liyuan Shi, Jiali Lin, Jingtian Chen, Jiayuan Zhong, and Chen Dong. Multi-granularity visualtextual jointly modeling for aspectlevel multimodal sentiment analysis. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06567-y.

Tian:2025:SOP

[47] Min Tian, Chaoshuai Xu, Xiaoming Wu, Jingshan Pan, Ying Guo, Wei Du, and Zhenguo Wei. Swpmmas: an optimized parallel max-min ant system algorithm based on the SW26010-pro processor. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06581-
0.

Park:2025:UDP

[48] Jung-Heum Park. Unpaired disjoint path covers in bipartite torus-like graphs with edge faults. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06572-
1.

Zhou:2025:ELF

[49] Heng Zhou, Qingguo Zhou, Xiaorun Tang, Jun Shen, Binbin Yong, and Yuanming Huang. Electrical load forecasting based on the fusion of multi-scale features extracted by using neural ordinary differential equation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06485-
z.

Chen:2025:ECC

[50] Songlin Chen, Weicheng Wang, Xiaoliang Chen, Maolin Zhang, Peng Lu, Xianyong Li, and Yajun Du. Enhancing Chinese comprehension and reasoning for large language models: an efficient LoRA fine-tuning and tree of thoughts framework. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06499-
7.

Du:2025:RAM

[51] Pan Du, Xinping Wang, Tiezhi Li, Chang Su, and Zhenyu Li. Resilience analysis of mine ventilation cyberphysical fusion system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06533-
8.

Panda:2025:EDP

[52] Sanjaya Kumar Panda, Thanmayee Pounjula, Bhargavi Ravirala, and David Taniar. An energy, delay and priority-aware task offloading algorithm for fog computing incorporating load balancing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06557-
0.

Rather:2025:AAT

[53] Bilal Ahmad Rather, Muhammad Imran, and Fozia Bashir Farooq. Algebraic analysis of p-type brooms and their application in allergic drugs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06600-

Wang:2025:DMO

[54] Yu Wang, Yongjie Ma, Quanxiu Li, and Yan Zhao. A dynamic multiobjective optimization evolutionary algorithm based on classification of environmental change intensity and collaborative prediction strategy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06480-
                                        4.

Uwimana:2025:STL

[55] Eustache Uwimana, Yatong Zhou, and Ndiaye Mareme Sall. A short-term load demand forecasting: Levenberg– Marquardt (LM), Bayesian regularization (BR), and scaled conjugate gradient (SCG) optimization algorithm analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06513-
                                        y.

Zhang:2025:LNV

[56] Hongying Zhang and Jiangbing Zeng. Lightweight network for visible-infrared person re-identification via self-distillation and multi-granularity information mining. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s11227-024-06543-6.

Huang:2025:CCD

[57] Feixiong Huang, Yubiao Pan, Huizhen Zhang, and Mingwei Lin. CD-NRocks: computable data nodes with RocksDB to improve the read performance of LSM-tree-based distributed key-value storage systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06526-
7.

Meng:2025:DMS

[58] Jing Meng, Zhenfang Zhu, Jiangtao Qi, and Huaxiang Zhang. DRSS: a multimodal sentiment analysis approach based on dual representation and self-supervised learning strategy. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06524-9.

Sun:2025:CCC

[59] Zhuoran Sun, Ying Ying Liu, and Parimala Thulasiraman. Cooperative, collaborative, coevolutionary multiobjective optimization on CPU–GPU multi-core. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06525-
8.

Villarrubia:2025:BSC

[60] Jorge Villarrubia, Luis Costero, Francisco D. Igual, and Katzalin Olcoz. Balanced segmentation of CNNs for multi-TPU inference. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06605-
                                        9.

Xiao:2025:GWP

[61] Zhengqing Xiao, Youliang Tian, Changgen Peng, Yangyang Long, and Chuanda Cai. Ghost-Weight protocol: a highly efficient blockchain consensus for IoT. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06565-0.

Li:2025:ASD

[62] Jinze Li, Xiangyu Meng, Zichen Qi, Dong Guo, and Cong Fu. Attack stage detection method based on vector reconstruction error autoencoder and explainable artificial intelligence. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06473-
                                        3.

Liang:2025:FSL

[63] Xingzhu Liang, Yachen Xu, Yu e Lin, and Chunjiong Zhang. Federated split learning via dynamic aggregation and homomorphic encryption on non-IID data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06612-
                                        w.

Zeng:2025:AAS

[64] Biqing Zeng, Liangqi Xie, Ruizhe Li, Yongtao Yao, Ruiyuan Li, and Huimin Deng. Aspect-aware semantic feature enhanced networks for multimodal aspect-based sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06472-4.

Anjali:2025:NSE

[65] Anjali and Anjana Gupta. A novel Shannon entropy-based backward cloud model and cloud K-means clustering. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06528-5.

Cai:2025:AKN

[66] Liang Cai, Shijie Zhao, Fanshuai Meng, and Tianran Zhang. Adaptive K-NN metric classification based on improved Kepler optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06559-
y.

Zhang:2025:NLC

[67] Longfei Zhang, Gang Wang, and Wei Zhou. A novel loop closure detection algorithm based on crossroad scenes. The Journal of Supercomputing, 81(1):??, January 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06488-
w.

Ma:2025:SSB

[68] Shichao Ma, Shengfa Miao, Shaowen Yao, Xin Jin, Xing Chu, Qian Yu, Yuling Tian, and Ruoshu Wang. SDHNet: a sampling-based dualstream hybrid network for long-term time series forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06495-
x.

Hong:2025:CPS

[69] Jiwon Hong, Hyeongmin Kim, Suhyeon Oh, Yerin Im, Hyeonseong Jeong, Hyunmin Kim, Eunkueng Jang, and Kyounggon Kim. Combating phishing and script-based attacks: a novel machine learning framework for improved client-side security. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06551-
6.

Chen:2025:MSL

[70] Yilin Chen, Xin Hu, Tao Lu, Lu Zou, and Xiangyun Liao. A multi-scale large kernel attention with U-Net for medical image registration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06489-
9.

Zhang:2025:MTC

[71] Xiufeng Zhang, Jinwei Zhou, and Guobin Qi. Multimodal temporal context network for tracking dynamic changes in emotion. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06484-
0.

Nooh:2025:EBC

[72] Sameer Abdullah Nooh. Enhancing beyond 5G connectivity and security: optimizing user-to-multiple AP associations with hybrid deep learning and innovative optimization techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06503-0.

Gao:2025:SPF

[73] Lei Gao, Jingfei Jiang, Jinwei Xu, Weijia Wang, and Pengbo Wu. SAPFIS: a parallel fuzzy inference system for air combat situation assessment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06521-
y.

Luo:2025:EWS

[74] Huilan Luo and Zhen Zeng. Enhancing weakly supervised semantic segmentation through multi-class token attention learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06618-
4.

Zeng:2025:PPP

[75] Jiangfeng Zeng, Xinyu Li, and Xiao Ma. PPTopicPLM: plug-and-play topicenhanced pre-trained language model for short-text rumor detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06549-0.

Ullah:2025:OVE

[76] Ihsan Ullah and Youn-Hee Han. Optimizing vehicular edge computing: graph-based double-DQN approaches for intelligent task offloading. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06599-4.

Chin:2025:DCE

[77] Zi Hau Chin, Vishnu Monn Baskaran, Chee Keong Tan, Ian K. T. Tan, and Timothy T. V. Yap. DAP-CBR: enhancing Bitcoin block propagation efficiency using dynamic compact block relay’s prefilling of transactions. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06468-
0.

Shahid:2025:EMR

[78] Maida Shahid, Muhammad Awais Hassan, Faiza Iqbal, Ayesha Altaf, Sayyed Wajihul Husnain Shah, Ana Visiers Elizaincin, and Imran Ashraf. Enhancing movie recommendations using quantum support vector machine (QSVM). The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06501-
2.

Sun:2025:CFR

[79] Yu Sun and Qicheng Liu. Collaborative filtering recommendation based on K-nearest neighbor and non-negative matrix factorization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06537-4.

Wang:2025:ATS

[80] Liangliang Wang, Lei Wang, Qiaoyong Jiang, Zhaoqi Wang, Wenqian Zhu, and Zhennan Wang. An adaptive transfer strategy guided by reference vectors for many-objective optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06547-
2.

Yan:2025:HUS

[81] Huizhen Yan, Fengyang Li, Ming-Chia Hsieh, and Jimmy Ming-Tai Wu. Highutility sequential pattern mining in incremental database. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06568-
                                        x.

Li:2025:TEN

[82] Dan Li, Hongbin Xia, and Yuan Liu. Triplet extraction network with dual gating mechanism and dependencyoriented attention. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06575-
                                        y.

Ozdag:2025:NHP

[83] Recep Özdag. A novel hybrid path planning method for sweep coverage of multiple UAVs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06574-
                                        z.

Chen:2025:RSD

[84] Yanping Chen, Feifan Ran, Xiaomin Jin, Haizhou Liu, and Zhongmin Wang. Robust service deployment for edge computing in industrial internet with joint profit awareness and multi-server collaboration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06609-
5.

Jing:2025:GCL

[85] Li Jing, Liu Ke, Zhang Lei, Yin Xiaoya, Jia Yuanyuan, and Jia Huinan. Geohash coding location privacy protection scheme based on entropy weight TOPSIS. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06511-
0.

Hu:2025:MAM

[86] Chengyu Hu, Jin Liu, Xingye Li, Meijing Li, and Huihua He. MST-ARGCN: modality-squeeze transformer with attentional recurrent graph capsule network for multimodal sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06588-
7.

Sahraneshinsamani:2025:MPP

[87] Nima Sahraneshinsamani, Sandra Catalán, and José R. Herrero. Mixed-precision pre-pivoting strategy for the LU factorization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06523-
w.

Fardad:2025:OPP

[88] Hamidreza Fardad, Faramarz Safi-Esfahani, and Behrang Barekatain. Optimizing production planning and sequencing in hot strip mills: an approach using multi-objective genetic algorithms. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06469-
                                        z.

Latifi:2025:MCB

[89] Faride Latifi, Ramin Nassiri, Mehran Mohsenzadeh, and Hamidreza Mostafaei. A Markov chain-based multi-criteria framework for dynamic cloud service selection using user feedback. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06508-
                                        9.

Xia:2025:FOC

[90] Huangzhi Xia, Yifen Ke, Riwei Liao, and Yunqiang Sun. Fractional order calculus enhanced dung beetle optimizer for function global optimization and multilevel threshold medical image segmentation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06592-
                                        x.

Wu:2025:CAM

[91] Xianglin Wu, Tianhao Meng, Jingwei Zhang, Qing Yang, and Jintao Chen. Consensus algorithm for maintaining large-scale access-control views of education data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06625-
5.

Wu:2025:API

[92] Zheqian Wu and Yingmin Li. An adaptive physics-informed deep learning approach for structural nonlinear response prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06586-
9.

Dalal:2025:DUO

[93] Sumit Dalal, Sarika Jain, and Mayank Dave. DepressionFeature: Underlying ontology for user-specific depression analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06585-
w.

Brahim:2025:NPR

[94] A. Hadj Brahim, H. Ali Pacha, M. Naim, and A. Ali Pacha. A novel pseudorandom number generator: combining hyperchaotic system and DES algorithm for secure applications. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06639-
                                        z.

Alqudsi:2025:IOS

[95] Yunes Alqudsi. Integrated optimization of simultaneous target assignment and path planning for aerial robot swarm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06620-
                                        w.

Tan:2025:MMI

[96] Linlin Tan, Yinghong Cao, Santo Banerjee, and Jun Mou. Multimedical image protection: compression– encryption scheme based on TLNN and mask cubes. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06624-
                                        6.

Singh:2025:VON

[97] Chandra Prakash Singh, Rohita Yamaganti, and Lokendra Singh Umrao. Variational Onsager Neural Networksbased fair proof-of-reputation consensus for blockchain with transaction prioritisation for smart cities. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06626-
                                        4.

He:2025:CVA

[98] Maowei He, Hongxia Zheng, Hanning Chen, Zhixue Wang, Xingguo Liu, Yelin Xia, and Haoyue Wang. A clustering and vector angle-based adaptive evolutionary algorithm for multi-objective optimization with irregular Pareto fronts. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06496-
w.

Chen:2025:MAR

[99] Lei Chen, Jieru Hou, Yunpeng Ma, and Yikai Zhao. A modified averageroulette cellular automaton algorithm for optimization tasks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06561-
4.

Bai:2025:CFC

[100] Xuefei Bai, Yongsong Wan, and Weiming Wang. CEPDNet: a fast CNNbased image denoising network using edge computing platform. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06646-
0.

Lei:2025:GGT

[101] Jianjun Lei, Sida Chen, and Ying Wang. GenerCTC: a general twostage contrastive training framework for text classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06628-
2.

Zhang:2025:OBFb

[102] Ran Zhang, Jingguo Bi, Lixiang Li, and Haipeng Peng. An optimal bound for factoring unbalanced RSA moduli by solving generalized implicit factorization problem. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06478-
y.

Naz:2025:ETL

[103] Sumera Naz, Aqsa Tasawar, Areej Fatima, Shariq Aziz Butt, and Zhoe Comas Gonzalez. An efficient 2-tuple linguistic cubic q-rung orthopair fuzzy CILOS–TOPSIS method: evaluating the hydrological geographical regions for watershed management in Pakistan. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06505-
y.

Yang:2025:MIM

[104] Dingyu Yang, Kangpeng Zheng, Shiyou Qian, Qin Hua, Kaixuan Zhang, Jian Cao, and Guangtao Xue. Mitigating interference of microservices with a scoring mechanism in largescale clusters. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06534-
7.

Lyu:2025:LYL

[105] Yifan Lyu, Tianze Zhang, Xin Li, Aixun Liu, and Gang Shi. LightUAV-YOLO: a lightweight object detection model for unmanned aerial vehicle image. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06611-
x.

Rajpopat:2025:CPD

[106] Subodh Rajpopat, Sunil Kumar, and Narinder Singh Punn. Cerebral palsy detection from infant using movements of their salient body parts and a feature fusion model. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06520-
z.

He:2025:MMS

[107] Yu He, Kang Zhou, and T. Lifang Tian.

Multi-modal scene graph inspired policy for visual navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06541-
8.

Ding:2025:NCM

[108] Pengfei Ding, Penghui Geng, and Weiwei Hu. A new controllable multiwing chaotic system: applications in high-security color image encryption. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06635-3.

Sayadi:2025:BPE

[109] Ladan Sayadi, Abdolah Amirany, Mohammad Hossein Moaiyeri, and Somayeh Timarchi. Balancing precision and efficiency: an approximate multiplier with built-in error compensation for error-resilient applications. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06563-2.

Shi:2025:ESM

[110] Leyi Shi, Qihang Yang, Luhan Gao, and Haowei Ge. An ensemble system for machine learning IoT intrusion detection based on enhanced artificial hummingbird algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06475-
                                        1.

Zhou:2025:EPP

[111] Sufang Zhou, Jianing Fan, Ke Yuan, Xiaoyu Du, and Chunfu Jia. Efficient privacy-preserving online medical pre-diagnosis based on blockchain. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06486-y.

Li:2025:AHG

[112] Bin Li, Haoyu Wang, Xaoyu Tan, Qiong Li, Jue Chen, and Xihe Qiu. Adaptive heterogeneous graph reasoning for relational understanding in interconnected systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06623-
7.

Hu:2025:EMT

[113] Kun Hu and Yuanbin Mo. An efficient multi-threshold image segmentation method for COVID-19 images using reinforcement learning-based enhanced sand cat algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06498-
8.

He:2025:CLN

[114] Qi-Qiao He, Xueyuan Gong, and Yain-Whar Si. Collaborative learning with normalization augmentation for domain generalization in time series classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06622-
                                        8.

Dai:2025:SSS

[115] Zhi-Qiang Dai, Jie Li, Yang-Jie Cao, and Yong-Xiang Zhang. SALSTM: segmented self-attention long short-term memory for long-term forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06493-z.

Ijaz:2025:EET

[116] Samia Ijaz, Saima Gulzar Ahmad, Kashif Ayyub, Ehsan Ullah Munir, and Naeem Ramzan. Energy-efficient time and cost constraint scheduling algorithm using improved multiobjective differential evolution in fog computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06550-
                                        7.

Qi:2025:EQL

[117] Han Qi, Xinyue Lv, Changqing Gong, and Abdullah Gani. Enhanced quantum long short-term memory by using bidirectional ring variational quantum circuit. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06636-
                                        2.

Gu:2025:ALD

[118] Runhe Gu and Luchun Lin. Application of latent Dirichlet allocation and autoencoder to real estate datasets. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06659-
9.

Yue:2025:FAB

[119] Xiaohan Yue, Gang Yi, Haoran Si, Haibo Yang, Shi Bai, and Yuan He. A face authentication-based searchable encryption scheme for mobile device. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06554-3.

Zhou:2025:SRC

[120] Sizhong Zhou. Spectral radius and component factors in graphs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06522-
x.

Guha:2025:SAD

[121] Krishnendu Guha. Self-aware decentralized security for real time approximate computing tasks in FPGAbased edge platforms. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06538-
                                        3.

Zheng:2025:MGL

[122] Jianxing Zheng, Min Li, Suge Wang, Jian Liao, and Xiaoya Wan. Multigranularity label-aware user interest modeling for news recommendation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06502-1.

Ranjan:2025:EES

[123] Rajeev Ranjan, Raj Anwit, and Prabhat Kumar. Energy efficient and sustainable mobile data collection in internet of things: a variable dimension SSO-based approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06630-
                                        8.

Gan:2025:MOO

[124] Wei Gan, Hongye Li, and Pengpeng Hao. Many-objective optimization algorithm based on the similarity principle and multi-mechanism collaborative search. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06553-
                                        4.

Chen:2025:RRD

[125] Jiang Chen and Weijie Ye. Roaen: reversed dependency graph and orthogonalgating strategy attention-enhanced network for aspect-level sentiment classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06542-
7.

Tang:2025:CEE

[126] Yiou Tang, Yan Ma, Chunling Xiao, Min Wu, and Guoyuan Zeng. Classification of EEG event-related potentials based on channel attention mechanism. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06627-

3. See correction [414].

Peng:2025:DBN

[127] Jianting Peng, Xuyu Xiang, Jiaohua Qin, and Yun Tan. Dual-branch networks for privacy-preserving crossmodal retrieval in cloud computing. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06643-3.

Merdassi:2025:NLT

[128] Imen Merdassi, Cherif Ghazel, and Leila Saidane. A novel location and time privacy-preserving approach for mobile cloud based on blockchain. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06604-
                                        w.

Chen:2025:STG

[129] Guihui Chen, Yuli Wei, Jiao Peng, Xinyu Zheng, Kai Lu, and Zhongbing Li. Spatio-temporal graph neural network based on time series periodic feature fusion for traffic flow prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06653-
                                        1.

Wang:2025:DIM

[130] Jingwen Wang, Zhoulin Cao, Chunzhi Xie, Yanli Li, Jia Liu, and Zhisheng Gao. DGN: influence maximization based on deep reinforcement learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06621-
                                        9.

Zhu:2025:ABS

[131] Chao Zhu and Qiang Ding. Aspectbased sentiment analysis via dual residual networks with sentiment knowledge. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06546-3.

Shi:2025:ILC

[132] Meifeng Shi, Guoyan Jia, and Makoto Yokoo. The improved local cost simulation algorithms based on coalition structure generation for solving distributed constraint optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06644-
2.

Jha:2025:SHT

[133] Vidyapati Jha and Priyanka Tripathi.

Selective hypothesis testing in cognitive IoT sensor network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06515-
w.

Sadeghi-Nasab:2025:ODP

[134] Alireza Sadeghi-Nasab and Mohsen Rahmani. Optimizing data privacy: an RFD-based approach to anonymization strategy selection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06642-
4.

Liang:2025:MPP

[135] Bo Liang and Jie Gao. A multiprocess parallel clustering algorithm for resource reconfiguration in cloud manufacturing. The Journal of Supercomputing, 81(1):??, January 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06607-
7.

Guo:2025:BTD

[136] Yan Ru Guo, Xiang Fei Yang, Xiang Yu Hua, and Jun Pan. Bilateral twodimensional linear discriminant analysis and its applications. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06660-
2.

Wang:2025:FSA

[137] Jie Wang, Chaochao Sun, and Yuan Peng. FedBat: a self-adapting bat algorithm-based federated learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06514-
x.

Wang:2025:SGF

[138] Liqin Wang, Pengcheng Yang, Xu Wang, Zhihong Xu, and Yongfeng Dong. Scene graph fusion and negative sample generation strategy for image-text matching. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06652-
2.

Bi:2025:TTS

[139] Yuandong Bi, Hong Zhu, Jing Shi, and Bin Song. TsCANet: Three-stream contrastive adaptive network for crossdomain few-shot learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06482-
2.

Mostafa:2025:EHE

[140] Reham R. Mostafa, Fatma A. Hashim, Ahmed M. Khedr, Zaher Al Aghbari, Imad Afyouni, Ibrahim Kamel, and Naveed Ahmed. EMGODV-Hop: an efficient range-free-based WSN node localization using an enhanced mountain gazelle optimizer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06616-
6.

Hussain:2025:SIC

[141] Adedoyin A. Hussain and Barakat A.

Dawood. A survey on IoT-cloud task in healthcare system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06629-
1.

Tripathi:2025:QLB

[142] Sarvapriya Tripathi, Himanshu Upadhyay, and Jayesh Soni. A quantum LSTM-based approach to cyber threat detection in virtual environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06615-
7.

Zhang:2025:DCB

[143] Qian Zhang, Xing Zhang, Zhiguang Chu, and Xiang Li. Density clusteringbased optimization model for trajectory data publication. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06617-
5.

Li:2025:CVQ

[144] Hao Li, Xiaohai He, Shuhua Xiong, Haibo He, and Honggang Chen. A compressed video quality enhancement algorithm based on CNN and transformer hybrid network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06654-
0.

Gavagsaz:2025:TID

[145] Elaheh Gavagsaz and Alireza Souri. Triangle-induced and degree-wise sampling over large graphs in social networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06613-
9.

Gautam:2025:IMS

[146] Rahul Kumar Gautam, Anjeneya Swami Kare, and S. Durga Bhavani. Interest maximization in social networks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06598-5.

Vlasic:2025:FST

[147] Andrew Vlasic, Hunter Grant, and Salvatore Certo. Feature selection through quantum annealing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06673-
x.

Zhao:2025:OBL

[148] Yanfen Zhao and Hao Liu. Oppositionbased learning Harris hawks optimization with steepest convergence for engineering design problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06649-
x.

Cao:2025:CED

[149] Zhengjun Cao and Huachen Ye. Comment on “Efficient design of an authenticated key agreement protocol for dewassisted IoT systems”. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06661-
1.

Hamdipour:2025:ARO

[150] Ali Hamdipour, Abdolali Basiri, Mostafa Zaare, and Seyedali Mirjalili. Artificial rabbits optimization algorithm with automatically DBSCAN clustering algorithm to similarity agent update for features selection problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06606-
8.

Tavares-Silva:2025:ASI

[151] Sthéfano Henrique Mendes Tavares-Silva, Sidney Marlon Lopes-Lima, Ricardo Paranhos-Pinheiro, Liosvaldo Mariano Santiago-Abreu, Rafael Diniz Toscano-Lima, and Sérgio Murilo Maciel Fernandes. Antivirus solution to IoT malware detection with authorial next-generation sandbox. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06506-
x.

Bashir:2025:CSF

[152] Zia Bashir, M. G. Abbas Malik, and Sadam Hussain. A computational study of fractional variable-order nonlinear Newton–Leipnik chaotic system with radial basis function network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06492-
0.

Li:2025:MCE

[153] Xiaohong Li, Jin Yao, Peng liu, and Yang Han. MC-CRS: enhanced conversational recommender system based on multi-contrastive learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06666-
w.

Qi:2025:HEV

[154] Han Qi, Sihui Xiao, Zhuo Liu, Changqing Gong, and Abdullah Gani. A high-efficiency variational quantum classifier for high-dimensional data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06676-8.

Wei:2025:RCV

[155] Hao Wei, Linchang Zhao, Ruiping Li, and Mu Zhang. RFAConv-CBM-ViT: enhanced vision transformer for metal surface defect detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06662-
0.

Pan:2025:PPB

[156] Yubiao Pan, Jianing Zhao, Yixiang Cai, Huizhen Zhang, and Mingwei Lin. PMCKV: pipeline-based multicompactions KV stores to improve the system performance. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06680-
y.

Liu:2025:IUC

[157] Yiwen Liu, Xiaoyu Zhang, Jinchao Zhu, and Panlong Tan. Improving underwater camouflage object segmentation with dual-decoder attention network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06584-
x.

Li:2025:PFL

[158] Hongjiao Li, Jiayi Xu, Ming Jin, and Anyang Yin. Personalized federated learning with global information fusion and local knowledge inheritance collaboration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06529-
4.

Xiang:2025:NDT

[159] Qiuhong Xiang, Hongfang Gong, and Cheng Hua. A new discrete-time denoising complex neurodynamics applied to dynamic complex generalized inverse matrices. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06601-
z.

Chang:2025:GLL

[160] Rui Chang, Gang Liu, Yao Qian, Haojie Tang, Gaoqiang Wang, and Durga Prasad Bavirisetti. GDNet: a low-light image enhancement network based on Ghost–Block and unique image decomposition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06683-
9.

Liu:2025:DRL

[161] Chang Liu, Xue-Feng Feng, Feng Li, Qing-Long Xian, Zhen-Hong Jia, Yu-Hang Wang, and Zong-Dong Du. Deep reinforcement learning combined with transformer to solve the traveling salesman problem. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06691-
9.

Bai:2025:MTP

[162] Xiangqi Bai, Lingtao Zhang, Yanyan Feng, Haoran Yan, and Quan Mi. Multivariate temperature prediction model based on CNN–BiLSTM and RandomForest. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06689-

Kim:2025:TBL

[163] Taero Kim, Changdae Oh, Hyeji Hwang, Eunkyeong Lee, Yewon Kim, Yunjeong Choi, Sungjin Kim, Hosik Choi, and Kyungwoo Song. TC-BERT: large-scale language model for Korean technology commercialization documents. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06597-
                                        6.

Zhang:2025:ECE

[164] Guozhen Zhang, Zhimin Yue, and Dajin Wang. 1-extra 3-component edge connectivity of modified bubblesort networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06610-
                                        y.

Bouleghlimat:2025:LLR

[165] Imene Bouleghlimat, Souheila Boudouda, and Salima Hacini. LSPP: a leakageresilient security approach for a cloudassisted big data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06657-
                                        x.

Mishra:2025:DNB

[166] Deepak Kumar Mishra and Pawan Singh Mehra. DiabeticChain: a novel blockchain approach for patient-centric diabetic data management. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06589-
6.

Liu:2025:NFE

[167] Suping Liu and Xiaomin Li. New fuzzy entropy via class-consistent technology with applications to attribute reduction for heterogeneous data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06631-
7.

Si:2025:LLC

[168] Yazhong Si, Chen Li, and Fan Yang.

LGASR: latent-content guided adversarial sand-dust image reconstruction strategy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06638-
0.

Zhang:2025:SSS

[169] Meijia Zhang, Junzheng Li, and Shengpeng Yu. Semi-supervised singleimage dehazing based on spatial-channel feature enhancement. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06665-
x.

Panda:2025:EDH

[170] Subrat Kumar Panda, Debasis Acharya, Dushmanta Kumar Das, and R. Kumar Rajagopal. An enhanced DV– Hop localization algorithm in wireless sensor networks with variable velocity strategy and human conception optimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06641-
                                        5.

AbdelAziz:2025:CBT

[171] Nabil M. AbdelAziz, Aliaa A. Ali, Soaad M. Naguib, and Lamiaa S. Fayed. Clustering-based topic modeling for biomedical documents extractive text summarization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06640-
                                        6.

Yu:2025:OCV

[172] Meng-Shiun Yu, Hao-Chun Chang, Chong-Teng Wang, Yu-Wei Tien, Tai-Liang Chen, and Jenq-Kuen Lee. Optimizing computer vision algorithms with TVM on VLIW architecture based on RVV. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06530-

Karamaz:2025:DMR

[173] Fatih Karamaz and Faruk Karaaslan.

Distance measures of r, s, t-spherical fuzzy sets and their applications in MCGDM based on TOPSIS. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06560-5.

Yuan:2025:CCD

[174] Ruosong Yuan, Wenwen Zhang, Xiaokang Dong, and Wanjun Zhang. Crns: CLIP-driven referring nuclei segmentation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06692-
8.

Nathiya:2025:BEW

[175] R. Nathiya, D. Meiyappan, Savari Prabhu, and Sudeep Stephen. Bandwidth of wk-recursive networks and its sparse matrix computation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06633-
5.

Zhao:2025:SSC

[176] Xu Zhao, Xiaohong Wang, and Bingge Cong. SGSLNet: stratified contextual graph pooling for point cloud segmentation with graph structural learning. The Journal of Supercomputing, 81(1):??, January 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06656-
y.

Alsuwat:2025:IMM

[177] Emad Alsuwat and Hatim Alsuwat. An improved multi-modal framework for fake news detection using NLP and Bi–LSTM. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06671-
z.

Guo:2025:EGC

[178] Rui Guo, Wen Xiong, Yungang Zhang, and Yanfang Hu. Enhancing game customer churn prediction with a stacked ensemble learning model. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06675-
9.

Diao:2025:CNT

[179] Jiarong Diao, Kai Cui, Yuling Huang, Chujin Zhou, Jianqing Li, and Haoyan Song. ChebyshevNet: a novel time series analysis model using Chebyshev polynomial. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06672-
y.

Huang:2025:ARC

[180] Zhaoman Huang, Yayu Yang, Mingzu Zhang, and Xing Yang. Assessing reliability in Complete Josephus Cube networks via strongly Menger edgeconnectivity. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06564-
1.

Kapil:2025:SBH

[181] Gayatri Kapil, Neeraj Kumar, Ashish Kumar Mourya, and Vijay Kumar. Securing big healthcare data using attribute and honey-based encryption in cloud environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06535-
6.

Taibi:2025:CNC

[182] Salaheddine Taibi, Lyazid Toumi, and Salim Bouamama. Complex network community discovery using fast local move iterated greedy algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06614-8.

Catalan:2025:ENN

[183] Izan Catalán, José Flich, and Carles Hernández. Exploiting neural networks bit-level redundancy to mitigate the impact of faults at inference. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06693-
7.

Parsafar:2025:RLB

[184] Parsa Parsafar. A reinforcement learning-based GWO–RNN approach for energy efficiency in data centers by minimizing virtual machine migration. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06510-
1.

Jia:2025:RRL

[185] Chao Jia, Tao Yu, and ZiJian Song. Robust reinforcement learning with augmented state for leveling control of multi-cylinder hydraulic system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06681-x.

Meng:2025:MMS

[186] Guang-Cong Meng, Yong-Xin Dong, and Yue-Hua Feng. A modified multistep splitting iteration and its variants for computing PageRank. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06669-
7.

Alammary:2025:IIP

[187] Ali Saleh Alammary. Investigating the impact of pretraining corpora on the performance of Arabic BERT models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06698-
2.

Xi:2025:IWH

[188] Mengyao Xi and Hao Liu. An improved wild horse optimization algorithm based on reinforcement learning for numerical and engineering optimizations. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06651-
3.

Wu:2025:IMN

[189] Danqing Wu and Chao Zhu. Interactive memory networks based on syntactic dependencies for aspect-level sentiment classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06594-
9.

Dabbagh:2025:CMM

[190] Arsalan Dabbagh, Mehrzad Karamimanesh, Kourosh Hassanli, and Ebrahim Abiri. CD-MAC: mixed-signal binary/ternary in-memory computing accelerator for power-constrained MAC processing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06700-
x.

Li:2025:INL

[191] Xue-Ning Li, Fang-Jiong Chen, Ye-Ping Lai, Peng Tang, and Xiao-Jun Liang. ICAT-net: a lightweight neural network with optimized coordinate attention and transformer mechanisms for earthquake detection and phase picking. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06664-
y.

Du:2025:TPR

[192] Jianbo Du, Jianjun Zhang, Jie Li, Jiaju Lv, Aijing Sun, Jing Jiang, Pengfei Du, and Jing Bai. Task placement and resource allocation for UAV and edge computing supported transportation systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06647-
z.

Li:2025:GCR

[193] Xingyu Li and Jinglei Liu. Graph convolutional and random Fourier feature mapping for hyperspectral image clustering. The Journal of Supercomputing, 81(1):??, January 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06696-
4.

Yin:2025:CEH

[194] Junqi Yin, Jesse Hines, Emily Herron, Tirthankar Ghosal, Hong Liu, Suzanne Prentice, Vanessa Lama, and Feiyi Wang. chatHPC: Empowering HPC users with large language models. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06637-1.

Mirzaie:2025:ERC

[195] Sara Mirzaie and Omid Bushehrian. Efficient root cause localization in IoTenabled water distribution networks by hierarchical anomaly analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06716-3.

Wu:2025:HHS

[196] Nengwu Wu, Wenjie Zhao, Yuxiang Chen, Jiahong Xiao, Jin Wang, Wei Liang, Kuan-Ching Li, and Nitin Sukhija. HFSL: heterogeneity split federated learning based on client computing capabilities. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06632-

Erdag:2025:JPF

[197] Özgür Erdag, James F. Peters, and Ömür Deveci. The Jacobsthal– Padovan–Fibonacci p-sequence and its application in the concise representation of vibrating systems with dual proximal groups. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06608-
                                        6.

Che:2025:ISD

[198] Xingliang Che, Wen Xiong, Xian Zhang, and Xitong Zhang. An integrated static and dynamic graph fusion approach for traffic flow prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06670-
                                        0.

Zhang:2025:LMS

[199] Xingpeng Zhang, Jing Xu, Dong He, Kaixin Wang, and Liping Wang. Lightweight multi-scale attention group fusion structure for nuclei segmentation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06710-9.

He:2025:MFL

[200] Jiangtao He. A multichannel fusion learning model with syntax for Chinese-oriented aspect-level sentiment classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06674-
w.

Dhibar:2025:ASI

[201] Sibasish Dhibar and Madhu Jain. AN-FIS simulation integrated in FM/FM/1(CV + WV) queue with Bernoulli service interruption and metaheuristic optimization for mathematical model. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06481-3.

Zhang:2025:OYA

[202] Yunjie Zhang, Guofeng Gao, Yadong Chen, and Zhenjian Yang. ODD-YOLOv8: an algorithm for small object detection in UAV imagery. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06703-

8. See correction [391].

Mohapatra:2025:SBH

[203] J. Mohapatra, L. Govindarao, and S. Priyadarshana. A splitting based higher-order numerical scheme for 2D time-dependent singularly perturbed reaction-diffusion problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06479-x.

Liu:2025:EQS

[204] Yutang Liu, Qin Zhang, and Ruoxia Li. Exponentially quasi-synchronization control of quaternion-valued memristive neural networks: matrix measure strategies and Frobenius norm methods. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06699-
                                        1.

Zhang:2025:IER

[205] Yunyi Zhang, Ye Du, Wei He, and Yu Tang. Inference of evidence reasoning rule with Gaussian distribution reliability and its application in safety assessment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06648-
                                        y.

Jahrmoi:2025:FAR

[206] Mohammad Amin Ghasvari Jahrmoi, Mehrdad Ashtiani, and Fatemeh Bakhshi. FaasFlows: an approach for reducing vendor lock-in and response time in serverless workflows. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06706-
                                        5.

Tiwari:2025:LLL

[207] Sambhavi Tiwari, Manas Gogoi, Shekhar Verma, and Krishna Pratap Singh.

Learning to learn: a lightweight meta-learning approach with indispensable connections. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06701-
w.

Zang:2025:DSA

[208] XiuBo Zang, HongBin Xia, and Yuan Liu. Diffusion social augmentation for social recommendation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06695-
5.

Saeedi:2025:EEC

[209] Ahmad Saeedi, Marjan Kuchaki Rafsanjani, and Samaneh Yazdani. Energy efficient clustering in IoT-based wireless sensor networks using binary whale optimization algorithm and fuzzy inference system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06556-
1.

Huang:2025:HCT

[210] Cheng-Ta Huang, Njabulo Sinethemba Shongwe, Hao-Yu Weng, Chi-Yao Weng, and Shiva Prasad Sirmulwar. Hybrid coding table-based semi-reversible data hiding using least significant bits and encryption. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06712-
                                        7.

Zhu:2025:PID

[211] Wen-Han Zhu, Rong-Xia Hao, Jou-Ming Chang, and Jaeun Lee. Packing internally disjoint Steiner paths of data center networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06545-
                                        4.

Mahapatra:2025:ADS

[212] Ajit Kumar Mahapatra, Nibedan Panda, and Binod Kumar Pattanayak. Adaptive dimensional search-based orthogonal experimentation SSA (ADOX-SSA) for training RBF neural network and optimal feature selection. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06507-w.

Zuo:2025:PDC

[213] Zhibin Zuo, Demin Wang, Xiaowei Nie, Xiaoduo Pan, Miaolei Deng, and Mimi Ma. PDCF-DRL: a contention window backoff scheme based on deep reinforcement learning for differentiating access categories. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06634-
                                        4.

Castello:2025:EGM

[214] Adrián Castelló, Héctor Martínez, Sandra Catalán, Francisco D. Igual, and Enrique S. Quintana-Ortí. Experienceguided, mixed-precision matrix multiplication with Apache TVM for ARM processors. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06720-
7.

Ga:2025:EGB

[215] Sangmin Ga, Paul Hyunbin Cho, Gordon Euhyun Moon, and Sungwon Jung. Efficient GNN-based social recommender systems through social graph refinement. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06682-
w.

Guo:2025:DSC

[216] Zhifeng Guo, Zijun Chen, Xue Sun, and Wenyuan Liu. H-dominant skyline community search in multi-valued networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06679-
5.

Suppalap:2025:SSV

[217] Siwakon Suppalap, Dawrawee Makmuang, Vipavee Damminsed, and Rabian Wangkeeree. Smooth support vector machine with rescaled generalized pinball loss for classification. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06697-3.

Vincy:2025:MLA

[218] G. Caroline Vincy and M. David Raj. Minimum linear arrangement and embedding of (K_{11} − C_{11})^{n} into specific graphs with optimal wirelength calculation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06519-
6.

Sezer:2025:PNP

[219] Bora Bugra Sezer and Sedat Akleylek. PPLBB: a novel privacypreserving lattice-based blockchain platform in IoMT. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06650-
4.

Liu:2025:MGE

[220] Huilin Liu, Yuhao Wu, Zihan Tang, Xiaolong Li, Shuzhi Su, Xingzhu Liang, and Pengfei Zhang. Multi-granularity enhanced feature learning for visibleinfrared person re-identification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.

com/article/10.1007/s11227-024-
06731-4.

Zhang:2025:RRM

[221] Jing Zhang, Bing Li, Yao Zhang, Yuguang Xu, and Hongan Li. Research on the recommendation method of urban location point of interest based on DTCN–EFFN-Transformer. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06742-1.

Song:2025:YGR

[222] Limei Song, Shikun Lu, Yu Tong, and Fengyi Han. YOLOv8s-GSW: a realtime detection model for hexagonal barbed wire breakpoints. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06738-
x.

Sarin:2025:GLQ

[223] Saket Sarin, Sunil K. Singh, Sudhakar Kumar, and Shivam Goyal. Geospectra: leveraging quantum-SAR and deep learning for enhanced geolocation in urban environments. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06619-
3.

Yue:2025:QLB

[224] Xingqi Yue, Lang Li, Qiuping Li, Jiahao Xiang, and Zhiwen Hu. QLW: a lightweight block cipher with high diffusion. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06707-
4.

Ren:2025:SKU

[225] Zhengwei Ren, Xiaojuan Li, Pei He, Rongwei Yu, Shiwei Xu, Yan Tong, and Jinshan Tang. In-situ key update and minimal key set of encrypted outsourced data under binary key-derivation tree. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06684-
8.

Mohammadi:2025:EHA

[226] Jamal Mohammadi, Mahmoud Shirazi, and Mehdi Kargahi. Energyharvesting-aware federated scheduling of parallel real-time tasks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06685-
7.

Tseng:2025:ESC

[227] Tzu-Lan Tseng, Wen-Yau Liang, and Hung-Lin Huang. Exploring a socialcuriosity-based algorithm for group recommender systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06590-
z.

Tian:2025:RNR

[228] Ran Tian, Shanwei Li, and Guoying Yang. Retraction note: Research on a distributed auto-negotiation model based on Stackelberg game theory. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06777-4.

Pu:2025:CBD

[229] Luoxi Pu and Zhi Quan. A CIMSbased decision support system for ecommerce: boosting real-time and effective digital marketing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06730-
5.

Xiong:2025:PZB

[230] Jie Xiong, Dongsheng Wang, Jian Yin, and Runfang Wu. Precise Z-Block positioning and dimension measurement using improved Canny edge detection and sub-pixel contour fitting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06769-
4.

Gao:2025:MMM

[231] Jiaqi Gao, Mingrui Fan, Yaru He, Daoqi Han, Yueming Lu, and Yaojun Qiao.

MACAE: memory module-assisted convolutional autoencoder for intrusion detection in IoT networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06704-
7.

Wang:2025:TMI

[232] Liang Wang, Yilin Li, and Lina Zuo. Trust management for IoT devices based on federated learning and blockchain. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06715-
4.

Wang:2025:LLP

[233] Shanshan Wang, Guofeng Zhao, Chuan Xu, Zhenzhen Han, and Shui Yu. LPQAA: a lightweight post-quantum access authentication scheme for satellite network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06687-
5.

Zhou:2025:ABR

[234] Wang Zhou, Ying Tian, Amin Ul Haq, and Sultan Ahmad. An autoencoderbased recommendation framework toward cold start problem. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06721-
6.

Zhang:2025:PCM

[235] Hao Zhang, Jiashu Wang, Yunhao Zhao, Lianjie Wang, Wenyin Zhang, Yeh-Cheng Chen, and Neal Xiong. Pocket convolution Mamba for brain tumor segmentation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06732-
3.

Wang:2025:WFC

[236] Yinuo Wang, Huanqi Zheng, Qiang Wu, Shengkun Yang, and Yucheng Zhou. White-faced capuchin optimizer: a new bionic metaheuristic algorithm for solving optimization problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06713-
6.

Swain:2025:IVM

[237] Smruti Rekha Swain, Anshu Parashar, Ashutosh Kumar Singh, and Chung Nan Lee. An intelligent virtual machine allocation optimization model for energy-efficient and reliable cloud environment. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06734-
1.

Qin:2025:ICR

[238] Xiwen Qin, Siqi Zhang, Xiaogang Dong, Tingru Luo, Hongyu Shi, and Liping Yuan. An improved conditional relevance and weighted redundancy feature selection method for gene expression data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06714-
                                        5.

Zhang:2025:MCG

[239] Hongwei Zhang, Zhidong Lu, Xiwei Chen, Shuai Lu, and Le Yao. Masked contrastive generative adversarial network for defect detection of yarn-dyed fabric. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06711-
                                        8.

Du:2025:RBI

[240] Junkun Du, Mingqing Wang, Xin Wang, Zhipeng Yang, Xiaojie Li, and Xi Wu. Reference-based image superresolution with attention extraction and pooling of residuals. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06587-
                                        8.

He:2025:DLB

[241] Lian ge He, Jun Song, Yan Zhang, and Xin yang Wu. Deep learning-based electric vehicle transmission system speed ratio prediction research. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06748-
9.

Ge:2025:CCD

[242] Mengjiao Ge, Wen Su, Jinfeng Gao, and Guoqiang Jia. CDMANet: central difference mutual attention network for RGB-D semantic segmentation. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06760-z.

Boujelben:2025:TAR

[243] Abir Boujelben and Ikram Amous.

Treating anomalies in rule bases associated to ontologies. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06688-
4.

Hu:2025:RUI

[244] Kuo-Jui Hu, Yi-Tsung Pan, Li-Wei Jiang, Sin-Der Lee, and Sheng-Long Kao. A robust underwater image enhancement algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06719-
0.

Wang:2025:PCP

[245] Jiaqi Wang and Dongqin Cheng. The 3-path-connectivity of pancake graphs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06702-9.

Kamatchi:2025:SEP

[246] K. Kamatchi and E. Uma. Securing the edge: privacy-preserving federated learning for insider threats in IoT networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06752-
                                        z.

NamvariTazehkand:2025:MSE

[247] Leila NamvariTazehkand and Saeid Pashazadeh. Modeling, simulation, and evaluation of causal order assurance techniques in causal broadcast algorithms using timed colored Petri nets. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06754-
                                        x.

Zhao:2025:SMO

[248] Ming Zhao. Synthetic minority oversampling technique based on natural neighborhood graph with subgraph cores for class-imbalanced classification. The Journal of Supercomputing, 81(1):??, January 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06655-
z.

Li:2025:PLM

[249] Xuefeng Li, Chen Chen, Jian Wei, Chensu Zhao, and Xiaqiong Fan. Precision localization method for fake news detection across multiple domains. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06776-5.

Huang:2025:RNC

[250] Rongbing Huang, Muhammad Farhan Hanif, Muhammad Kamran Siddiqui, Mazhar Hussain, and Muhammad Faisal Hanif. Retraction note: On connection number-based topological indices and entropy measures for triangular γ-graphyne network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06781-
8.

Wang:2025:MIB

[251] Wen-Chuan Wang, Wei-Can Tian, Kwok-Wing Chau, and Hongfei Zang. MSBES: an improved bald eagle search algorithm with multi- strategy fusion for engineering design and water management problems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06727-
                                        0.

Madani:2025:FEH

[252] Bachir Madani, Mohamed Salah Azzaz, Said Sadoudi, and Redouane Kaibou. Fast and efficient hardware architecture of Chebyshev polynomials algorithm for resisting to side channel attacks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06761-
                                        y.

Pacheco:2025:SLR

[253] A. Lopez Pacheco and J. Aguilar. Systematic literature review on quantum applications in nanotechnology. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06747-w.

Khah:2025:HML

[254] Yashar Pourardebil Khah, Mirsaeid Hosseini Shirvani, and Homayun Motameni. A hybrid machine learning approach for feature selection in designing intrusion detection systems (IDS) model for distributed computing networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06677-

7. See correction [428].

Ling:2025:MOP

[255] Qinghua Ling, Zexu Li, Wenkai Liu, Jinlong Shi, and Fei Han. Multiobjective particle swarm optimization based on particle contribution and mutual information for feature selection method. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06762-
x.

Wei:2025:PLR

[256] Xing Wei, Jiong Xia, Cang Liu, Qi wen He, Jun Chen, Zhen Wei, Chong Zhao, Fan Yang, and Yang Lu. Proposallevel reliable feature-guided contrastive learning for SFOD. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06773-
8.

Zarei:2025:OES

[257] Sevda Zarei, Sadoon Azizi, and Awder Ahmed. Optimizing edge server placement and load distribution in mobile edge computing using ACO and heuristic algorithms. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06780-
9.

Chen:2025:CCR

[258] Fei Chen, Carter Zhang, and Bo Ning.

CRKG: combining retrieval knowledge with generative language models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06728-z.

Nie:2025:CSK

[259] Shiqiang Nie, Jie Niu, Fangxing Yu, Jianqiang Ma, Xingxing Zhu, and Weiguo Wu. Constructing a scalable key-value store engine on multidisk system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06725-
2.

Tan:2025:HPR

[260] Kai Tan, Dongyang Zhan, Lin Ye, Hongli Zhang, Binxing Fang, and Zhihong Tian. A high-performance real-time container file monitoring approach based on virtual machine introspection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06779-
2.

Xie:2025:LIY

[261] Yinggang Xie and Yanwei Zhao. Lightweight improved YOLOv5 algorithm for PCB defect detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.

com/article/10.1007/s11227-024-
06739-w.

Wang:2025:IBF

[262] Jingyuan Wang, Yuan Zhao, Wenyan Wang, and Ziheng Wu. Improving bearing fault diagnosis method based on the fusion of time-frequency diagram and a novel vision transformer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06793-
4.

Ardakani:2025:ICS

[263] Saeid Pourroostaei Ardakani, Jianwei Hu, Jing Zhang, Kaifeng Jin, Tianhong Cai, Anthony Graham Bellotti, and Xiuping Hua. Identifying crowdfunding storytellers who deliver successful projects: a machine learning approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06785-
4.

Zhao:2025:AOE

[264] Yongpeng Zhao, Shengwei Fu, Langlang Zhang, and Haisong Huang. Aitken optimizer: an efficient optimization algorithm based on the Aitken acceleration method. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06709-
2.

Shirini:2025:ILR

[265] Kimia Shirini, Meysam Balaneshin Kordan, and Sina Samadi Gharehveran. Impact of learning rate and epochs on LSTM model performance: a study of chlorophyll-a concentrations in the Marmara Sea. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06806-
                                        2.

Ren:2025:PPT

[266] Qingxin Ren and Feng Feng. PID parameter tuning optimization based on multi-strategy fusion improved zebra optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06548-
                                        1.

Wang:2025:DNN

[267] Xiaolong Wang, Hedi An, Jinsong Zhang, Dongya Huang, and Junxian Wen. DPSMUNet: a new network based on a dual-pooling self-attention module for carotid artery plaque segmentation in ultrasound images. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06770-
                                        x.

Lin:2025:ATD

[268] Zhifeng Lin, Yilu Chen, Yanyue Xie, Chuandong Chen, Jun Yu, and Jianli Chen. An analytical timingdriven placer for modern heterogeneous FPGAs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06755-
w.

Zhu:2025:FSS

[269] Suxia Zhu, Yunmeng Wang, and Guanglu Sun. Federated semi-supervised learning based on truncated Gaussian aggregation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06798-
z.

Lloria:2025:OMW

[270] Diego Lloria, Sandra Roger, Germán León, José M. Badía, Carmen Botella-Mascarell, and Jose A. Belloch. Optimizing millimeter wave MIMO channel estimation through GPU-based edge artificial intelligence. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06795-
2.

Khan:2025:BIN

[271] Abdullah Ayub Khan, Asif Ali Laghari, Abdullah M. Baqasah, Rex Bacarra, Roobaea Alroobaea, Majed Alsafyani, and Jamil Abedalrahim Jamil Alsayaydeh. BDLT-IoMT — a novel architecture: SVM machine learning for robust and secure data processing in Internet of Medical Things with blockchain cybersecurity. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06782-
                                        7.

Yu:2025:ADS

[272] Dongxian Yu, Xiaoyu Zhou, Ali Noorian, and Mehdi Hazratifard. An AI-driven social media recommender system leveraging smartphone and IoT data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06722-

5. See correction [390].

Li:2025:PDP

[273] Yifan Li, Shuming Zhou, and Eddie Cheng. Paired 2-disjoint path covers of balanced hypercube under the partitioned edge fault model. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06736-
                                        z.

Zaminkar:2025:IKM

[274] Mina Zaminkar. An improved Kmeans and DPC-empowered clustering approach for efficient routing in the FANET. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06724-
3.

Zhang:2025:DRL

[275] Lili Zhang, Ke Zhang, Kang Yang, Wei Wei, Jing Li, Hongxin Tan, Pei Yu, Yucheng Han, and Xudong Yang. Driving risks from light pollution: an improved YOLOv8 detection network for high beam vehicle image recognition. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06809-
z.

Tang:2025:PPU

[276] Can Tang, Tao Peng, Xingxing Xie, and Junhu Peng. 3D path planning of unmanned ground vehicles based on improved DDQN. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06690-
w.

Yang:2025:HPI

[277] Hao Yang, Junyang Yu, and Rui Zhai. High-precision intrusion detection for cybersecurity communications based on multi-scale convolutional neural networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06737-
y.

Li:2025:EBN

[278] Shiyong Li, Huan Liu, Wenzhe Li, and Wei Sun. Edge bank: a novel resource pricing and management system for edge service provider. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06578-
                                        9.

Yuxin:2025:MMH

[279] Zhang Yuxin, Zhao Enjiao, Liang Hong, and Zhou Wentao. MATD3 with multiple heterogeneous sub-networks for multi-agent encirclement-combat task. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06756-9.

Kamal:2025:EDN

[280] Sadia Kamal, Parth Sharma, P. K. Gupta, Mohammad Khubeb Siddiqui, Ankush Singh, and Abhijit Dutt. DVTXAI : a novel deep vision transformer with an explainable AI-based framework and its application in agriculture. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06494-
                                        y.

Li:2025:UUS

[281] Junyan Li, Wenyong Dong, and Xuewen Gui. uTransformer: unified spatialtemporal transformer with external factors for traffic flow forecasting. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06774-7.

Alioghli:2025:EMT

[282] Abdul Amir Alioghli and Feyza Yildirim Okay. Enhancing multivariate timeseries anomaly detection with positional encoding mechanisms in transformers. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06694-
6.

Camara:2025:AAS

[283] Jesús Cámara, Javier Cuenca, Victor Galindo, Arturo Vicente, and Murilo Boratto. An autotuning approach to select the inter-GPU communication library on heterogeneous systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06794-
3.

MohammedJany:2025:DMC

[284] Shaik MohammedJany, Chandra Bhushana Rao Killi, Shaik Rafi, and Syed Rizwana. Detecting multimodal cyber-bullying behaviour in social-media using deep learning techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06772-
                                        9.

Saini:2025:VAA

[285] Jatinderkumar R. Saini and Shraddha Vaidya. A veracity assessment algorithm for classification of healthcare information using feature bag mash-up approach. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06500-
                                        3.

Zhang:2025:OBFa

[286] Bolei Zhang and Lifa Wu. Online budget-feasible mobile crowdsensing with constrained reinforcement learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06767-
                                        6.

Sun:2025:PCC

[287] Zhuoran Sun, Ying Ying Liu, and Parimala Thulasiraman. Publisher correction: Cooperative, collaborative, coevolutionary multi-objective optimization on CPU–GPU multi-core. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06803-
                                        5.

Castorena:2025:ECD

[288] Carlos Castorena, Jesus Lopez-Ballester, Juan A. De Rus, Maximo Cobos, and Francesc J. Ferri. Edge computing for driving safety: evaluating deep learning models for cost-effective sound event detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06796-
1.

Huang:2025:GAI

[289] Haojing Huang, Jiajun Li, Fei Lu, and Jianxin Li. Generative adversarial imitation learning computing task offloading scheme for optimizing of generated sample utilization and system overhead. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06744-
z.

Wang:2025:VDA

[290] Yuhai Wang, Shuobo Xu, Peng Wang, Lele Liu, YanShun Li, and Ze Song. Vehicle detection algorithm based on improved RT–DETR. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06766-
7.

Ouamri:2025:CSS

[291] Mohamed Amine Ouamri, Turki Alharbi, Daljeet Singh, and Zenadji Sylia. A comprehensive survey on software-defined wide area network (SD-WAN): principles, opportunities and future challenges. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06718-
1.

Mat:2025:ESB

[292] Abdullah Ugur Mat and Ayse Nurdan Saran. Enhancing session-based trip recommendations using matrix factorization: a study on algorithm efficiency and resource utilization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06726-
1.

Li:2025:MOF

[293] Yuanhe Li, Wenjian Zhong, and Yuanqing Wu. Multi-objective flexible jobshop scheduling via graph attention network and reinforcement learning. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06741-2.

Ahmad:2025:BFS

[294] Waseem Ahmad, Aurang Zeb, Muhammad Asif, and Muzhou Hou. Bipolar fuzzy soft hamacher aggregations operators and their application in triage procedure for handling emergency earthquake disaster. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06757-
8.

Salari-Hamzehkhani:2025:IID

[295] Behnam Salari-Hamzehkhani, Mehdi Akbari, and Faramarz Safi-Esfahani. Introducing an improved deep reinforcement learning algorithm for task scheduling in cloud computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06668-8.

Torki:2025:SFV

[296] Omid Torki, Hamid Mala, and Maede Ashouri-Talouki. Secure fully-verifiable outsourcing of modular exponentiation: two servers check each other. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06745-y.

Alasiry:2025:SVL

[297] Areej Alasiry and Mohammed Qayyum.

A smart vista-lite system for anomaly detection and motion prediction for video surveillance in vibrant urban settings. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06753-

y. See correction [413].

Ghosh:2025:ECS

[298] Saurav Ghosh, Kanyaka Chakraborty, Piyali Bagchi Khatua, and Utpal Biswas. Efficient charging schedules in a rechargeable wireless sensor network with multiple chargers. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06804-
                                        4.

Zhang:2025:MMC

[299] Yaozeng Zhang, Jing Ma, and Yuguang Jia. MCAN: multimodal cross-aware network for fake news detection by extracting semantic-physical feature consistency. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06815-
                                        1.

Saberali:2025:EWB

[300] Behzad Saberali, Kai Zhang, Farzad Saberali, Fatna Adinani Said, and Lu Yang. Enhanced well-based surrogate reservoir modeling with integrated streamlines simulation data. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06686-6.

Trivedi:2025:EAA

[301] Chandan Trivedi, Keyur Parmar, and Udai Pratap Rao. ALMASH : an anonymity-based lightweight mutual authentication scheme for Internet of Healthcare Things. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06801-
7.

Emami:2025:ALS

[302] Hojjat Emami. An adaptive local search-based arithmetic optimization algorithm for unmanned aerial vehicle placement. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06812-
4.

Tang:2025:YSO

[303] Qiang Tang, Chang Su, Yuan Tian, Shibin Zhao, Kai Yang, Wei Hao, Xubin Feng, and Meilin Xie. YOLO-SS: optimizing YOLO for enhanced small object detection in remote sensing imagery. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06765-
8.

Liu:2025:TTF

[304] Chen Liu, Tianhao Yu, Xianghong Zhou, Lixin Zhou, and Xiaoyu Gong. TSESRec: A transformer-facilitated set extension model for session-based recommendation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06814-
                                        2.

Yang:2025:LLV

[305] Gaoming Yang, Chenlong Yu, Xiujun Wang, Xianjin Fang, and Ji Zhang. LVAST: a lightweight vision transformer for effective arbitrary style transfer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06787-
                                        2.

Yang:2025:RID

[306] Yixuan Yang, Mingrong Dong, Kai Zeng, and Tao Shen. Research on the improvement of domain generalization by the fusion of invariant features and sharpness-aware minimization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06797-
                                        0.

Wang:2025:RDL

[307] Yiqiao Wang, Jinling Chen, Bo Yang, Yu Chen, Yanlin Su, and Rong Liu. RT-DETRmg: a lightweight realtime detection model for small traffic signs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06800-
                                        8.

Anonymous:2025:JNa

[308] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNb

[309] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNc

[310] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNd

[311] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNe

[312] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNf

[313] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNg

[314] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Oviesi:2025:QNN

[315] Safura Oviesi, Mohamad Jafar Tarokh, and Mohamad kazem Momeni. Quantum neural network-assisted learning for small medical datasets: a case study in emphysema detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06740-
3.

Yue:2025:NFS

[316] Zhongzheng Yue, Fei Wen, and Zhijun Li. Neighbor full sum distinguishing total coloring of planar graphs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06825-z.

Bhattacharya:2025:SVS

[317] Tathagata Bhattacharya, Adithya Vardhan Peddi, Srikanth Ponaganti, and Sai Teja Veeramalla. A survey on various security protocols of edge computing. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06678-
6.

Zhang:2025:TTS

[318] Haiping Zhang, Haixiang Lin, Dongjing Wang, Dongyang Xu, Fuxing Zhou, Liming Guan, Dongjing Yu, and Xujian Fang. TSCANet: a twostream context aggregation network for weakly-supervised temporal action localization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06810-
6.

Yang:2025:GGF

[319] Qimeng Yang, Yi Liu, Lanlan Lu, and Lei Liu. GFIDF: gradual fusion intent detection framework. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06708-
3.

Shahba:2025:DFW

[320] Leyla Shahba, Ahmad Heidary-Sharifabad, and Mohammadreza Mollahoseini Ardakani. Detection of fake web pages and phishing attacks with rabbit optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06658-
w.

Khan:2025:MMA

[321] Ajmal Khan, Naveed Iqbal, Zeeshan Kaleem, Zul Qarnain, and Mohammed M. Bait-Suwailam. A multimodel approach for predicting electric vehicle specifications and energy consumption using machine learning. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06820-4.

Guerrero-Pantoja:2025:AUM

[322] David Guerrero-Pantoja, Erick Pautsch, Clara Almeida, Silvio Rizzi, George K. Thiruvathukal, and Maria Pantoja. Accelerating uncertainty methods for distributed deep learning on novel architectures. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06818-
y.

Manchala:2025:NSP

[323] Pravali Manchala and Manjubala Bisi. A novel source project and optimized training data selection approach for cross-project fault prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06750-1.

Mnasri:2025:HBF

[324] Sami Mnasri, Dorsaf Salah, and Hanen Idoudi. A hybrid blockchain and federated learning attention-based BERT transformer framework for medical records management. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06816-
0.

Mou:2025:PSL

[325] Jianhui Mou, Jian Wang, Yangwei Wang, Ming Xiang, and Lihua Zhang. Pattern synthesis of linear antennaarray for high gain and low sidelobe level based on sand cat swarm optimization algorithm. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06763-
w.

Tang:2025:MME

[326] Jiajia Tang, Binbin Ni, Yutao Yang, Yu Ding, and Wanzeng Kong. MECG: modality-enhanced convolutional graph for unbalanced multimodal representations. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06729-
y.

Ma:2025:TSL

[327] Fei Ma, Aihua Hou, Feixia Yang, and Guangxian Xu. Tensor subspace learning and folded-concave function regularization for hyperspectral anomaly detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06791-
                                        6.

Asiamah:2025:SEL

[328] Emmanuel Acheampong Asiamah, Nana Kwadwo Akrasi-Mensah, Prince Odame, Eliel Keelson, Andrew Selasi Agbemenu, Eric Tutu Tchao, Mohammed Al-Khalidi, and Griffith Selorm Klogo. A storage-efficient learned indexing for blockchain systems using a sliding window search enhanced online gradient descent. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06805-
                                        3.

Khomami:2025:CGGa

[329] Mohammad Mehdi Daliri Khomami, Mohammad Reza Meybodi, and Alireza Rezvanian. A cellular Goore gamebased algorithm for finding the shortest path in stochastic multi-layer graphs. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06786-3.

Li:2025:IPC

[330] Zhe Li, Jinsong Wang, and Yi Li. An improved PBFT consensus algorithm based on reputation and gaming. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06822-2.

Zhang:2025:EVV

[331] Weilin Zhang, Lihua You, Hechao Liu, and Xiaona Fang. The expected values and variances for degreebased topological indices in random spiro chains. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06808-
0.

Chen:2025:SRN

[332] Zeqiang Chen, Zhiqing Li, Xu Tang, Lai Chen, and Nengcheng Chen. STP-Net: a recurrent neural network for spatiotemporal processes predictive learning. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06823-
1.

Zhu:2025:PPF

[333] Libo Zhu and Xiang Chen. Privacy protection in federated learning: a study on the combined strategy of local and global differential privacy. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06845-
9.

Wang:2025:GAE

[334] Fei Wang, Ming Ju, Xianxun Zhu, Qiuyu Zhu, Haiquan Wang, Chunhua Qian, and Rui Wang. A geometric algebra-enhanced network for skin lesion detection with diagnostic prior. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06833-
z.

Ramesh:2025:UAB

[335] Parameswaran Ramesh, P. T. V. Bhuvaneswari, V. S. Dhanushree, G. Gokul, and S. Sahana. User associationbased load balancing using reinforcement learning in 5G heterogeneous networks. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06788-
1.

Rastogi:2025:PAU

[336] Eshita Rastogi, Mukesh Kumar Maheshwari, and Ayush Rastogi. Performance analysis of user power consumption in NR-unlicensed DRX. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06839-
7.

Wang:2025:OMC

[337] Jian wei Wang, Qing Zhang, and Cheng sheng Pan. Optimization method of C2 system architecture based on ALCARO. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-024-06768-
5.

Li:2025:AUC

[338] Hongzhi Li, Zhanghao Ren, Guoqing Zhu, and Jiaxi Wang. AE-UNet: a composite lung CT image segmentation framework using attention mechanism and edge detection. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06874-
4.

Krishna:2025:CCI

[339] Charu Krishna, Divya Kumar, and Dharmender Singh Kushwaha. CovMed-Care: confluence of internet of things, blockchain and machine learning for remote monitoring system of pandemic patients. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06751-
0.

Du:2025:RAB

[340] Yu Du, Liwei Yang, and Yuchuan Luo. Resource allocation based on optimized cellular network AP layout for visible light communication heterogeneous network. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06834-
y.

He:2025:PCE

[341] Qi He, Yan Wang, Jianxi Fan, and Baolei Cheng. Parallel construction of edge-independent spanning trees in complete Josephus cubes. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06784-
                                        5.

Patil:2025:PSC

[342] Rajendra Patil, Sivaanandh Muneeswaran, Vinay Sachidananda, Peng Hongyi, and Mohan Gurusamy. PRIOR-ITI: scoring and categorization-based threat prioritization. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06465-
                                        3.

Ahmad:2025:DDP

[343] Farooq Ahmad, Xinfeng Zhang, Zifang Tang, Fahad Sabah, Muhammad Azam, and Raheem Sarwar. Deep deterministic policy gradients with a self-adaptive reward mechanism for image retrieval. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06764-
                                        9.

Cakici:2025:DRT

[344] Muhammed Emir Çakici, Feyza Yildirim Okay, and Suat Özdemir. Deepat: a real-time deep learning based model for aircraft tracking system. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06759-
6.

Lu:2025:SIR

[345] Zhengyang Lu, Weifan Wang, Tianhao Guo, and Feng Wang. Single-image reflection removal via self-supervised diffusion models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06837-
9.

Shang:2025:GIM

[346] Mengying Shang, Mengnan Tian, and Xinduan Wang. GA-DE: an integrated meta-heuristic approach for optimizing feedforward neural networks. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06799-y.

Ohue:2025:IMM

[347] Masahito Ohue, Nobuaki Yasuo, and Masami Takata. Innovations in mathematical modeling, AI, and optimization techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06861-

Yu:2025:DPW

[348] Bengong Yu, Chengwei Cao, and Ying Yang. Dynamic position weighting aspect-focused graph convolutional network for aspect-based sentiment analysis. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06783-
                                        6.

Kia:2025:HAR

[349] Zohre Kia, Meisam Yadollahzaeh-Tabari, and Homayun Motameni. Human activity recognition by body-worn sensor data using bi-directional generative adversarial networks and frequency analysis techniques. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06743-
                                        0.

Lin:2025:RRI

[350] Xinwei Lin, Yubiao Pan, Wenjuan Feng, Huizhen Zhang, and Mingwei Lin. RIOKV: reducing iterator overhead for efficient short-range query in LSM-tree-based key–value stores. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06735-
                                        0.

Zhang:2025:NHA

[351] Yanjun Zhang, Xin Guo, Hongchen Guo, and Yichen Zhang. A new hardware architecture of high-performance real-time texture classification system based on FPGA. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06705-
6.

Liu:2025:CPF

[352] Shiguo Liu, Dejian Wei, Junzhong Zhang, Xurui Ji, and Hui Cao. ConvBiFuseNet: a parallel fusion model with routing attention for MRI brain tumor classification. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06758-
7.

Giveki:2025:SIR

[353] Davar Giveki and Sajad Esfandyari. Semantic image representation for image recognition and retrieval using multilayer variational auto-encoder, InceptionNet and low-level image features. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06792-5.

Sinha:2025:BCA

[354] Bam Bahadur Sinha, Ramnish Sinha, and Vishnu Priye. Beyond classical approaches: redefining the landscape of high-accurate movie recommendation using QNN. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06746-
                                        x.

Ge:2025:DMO

[355] Fangzhen Ge, Xuan Zhao, Debao Chen, Longfeng Shen, and Huaiyu Liu. A dynamic multi-objective optimization algorithm based on probabilitydriven prediction and correlation-guided individual transfer. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06832-
                                        0.

Martin-Salinas:2025:EAV

[356] Ignacio Martin-Salinas, Jose M. Badia, Oscar Valls, German Leon, Rocio del Amor, Jose A. Belloch, Adrian Amor-Martin, and Valery Naranjo. Evaluating and accelerating vision transformers on GPU-based embedded edge AI systems. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06807-
                                        1.

Zhang:2025:VTB

[357] Enzhi Zhang, Rui Zhong, Xingbang Du, Mohamed Wahib, and Masaharu Munetomo. Vision transformer-based meta loss landscape exploration with actor-critic method. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06867-
3.

Kong:2025:RSS

[358] Dekun Kong and Wengaung Yang.

Research on site selection of emergency material reserve based on set pair analysis and TOPSIS integration method: a case study of Hebei Province, China. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06813-

3. See correction [456].

Sakano:2025:NNP

[359] Koh Sakano, Kairi Furui, and Masahito Ohue. NPGPT: natural product-like compound generation with GPT-based chemical language models. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06860-
w.

Gan:2025:IDM

[360] Lian Gan, YuHong Du, Shuai Wang, WeiJia Ren, ZiQi Rong, and XinLong Li. Intelligent decision making algorithm for path planning based on reference linguistic fuzzy set. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06842-
y.

Lopez-Oliva:2025:EQC

[361] Vicente Lopez-Oliva, Jose M. Badia, and Maribel Castillo. Efficient quantum circuit contraction using tensor decision diagrams. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06836-
                                        w.

Espinosa:2025:SEG

[362] Elena Espinosa, Ricardo Quislant, Rafael Larrosa, and Oscar Plata. SeqMatcher: efficient genome sequence matching with AVX-512 extensions. The Journal of Supercomputing, 81(1): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06789-0.

Ghadiri:2025:IBL

[363] Mohammad Javad Ghadiri and Mehri Bagherian. Improved binary linear programming models for finding maximum edge bi-clique in bipartite graphs. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06733-
                                        2.

Liang:2025:RMP

[364] Xinyu Liang, Guannan Si, Jianxin Li, Pengxin Tian, Zhaoliang An, and Fengyu Zhou. Relational message passing with mutual information maximization for inductive link prediction. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06749-
8.

Wang:2025:MPG

[365] Kexin Wang, Dingrui Xue, Yingdong Gou, Wanlong Qi, Bo Li, Jiancheng Liu, Yinglong Feng, and Yuqing Lin. Meta-path-guided causal inference for hierarchical feature alignment and policy optimization in enhancing resilience of UWSoS. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06848-
6.

Dehkordi:2025:EIC

[366] Afsaneh Banitalebi Dehkordi. EDBLSD-IIoT: a comprehensive hybrid architecture for enhanced data security, reduced latency, and optimized energy in industrial IoT networks. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06872-
6.

Jha:2025:AOR

[367] Mukesh Kumar Jha and Mohit Kumar.

An autonomic offloading and resource allocation technique for IoT applications in edge computing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06491-
                                        1.

Chung:2025:FGG

[368] Wu-Chun Chung, Jyun-Sen Tong, and Zhi-Hao Chen. A fine-grained GPU sharing and job scheduling for deep learning jobs on the cloud. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06849-
                                        5.

Abideen:2025:AIS

[369] Syed Zain Ul Abideen, Abdul Wahid, Mian Muhammad Kamal, Nouman Imtiaz, Nabila Sehito, Yousef Ibrahim Daradkeh, Mahmoud Ahmad Al-Khasawneh, Abdullah Alwabli, and Inam Ullah. Advancements in IoT system security: a reconfigurable intelligent surfaces and backscatter communication approach. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06819-
                                        x.

Guo:2025:ROP

[370] Yuepeng Guo, ZhenPing Lan, Yanguo Sun, Yuheng Sun, Xinxin Li, Yuru Wang, and Bo Li. Research on occlusion pedestrian re-identification based on ViT model. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06831-
1.

Diaz-Cano:2025:AMW

[371] Roberto Díaz-Cano, Francesc Folch, Enrique S. Quintana-Ortí, and Pedro Alonso-Jordá. Acceleration of the MVS workflow using graphics processors. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06835-
x.

Xiao:2025:LGL

[372] Senyue Xiao, Jianhua Liu, Zeming Pan, Shaoze Wang, Yang Yang, Zilong Song, and Anni Fan. LiteYOLO-GHG: a lightweight YOLOv8-based algorithm for transformer bushing fault detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06852-
w.

Mehtarizadeh:2025:SPP

[373] Homa Mehtarizadeh, Najme Mansouri, Behnam Mohammad Hasani Zade, and Mohammad Mehdi Hosseini. Stock price prediction with SCA–LSTM network and statistical model ARIMA–GARCH. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06775-
6.

Wu:2025:SGO

[374] Di Wu, Le Zhang, and Yao Chen. Syntactic-guided optimization of imagetext matching for intra-modal modeling. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06840-0.

Wang:2025:CCCb

[375] Zhen Wang, Dong Zhao, Ali Asghar Heidari, Huiling Chen, and Guoxi Liang. CGWRIME: collaboration and competition-boosted RIME optimizer for engineering optimization problems. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06817-
                                        z.

Lei:2025:LPP

[376] Lei Zhang, Yongbo Bai, Shiyi Lin, Shuaishuai Lian, Yijia Geng, and Zhili Liu. Location privacy protection scheme of user collaborative probabilistic indistinguishability based on Hyperledger Fabric. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06854-
                                        8.

Ewees:2025:OFS

[377] Ahmed A. Ewees, Mohammed M. Alshahrani, Abdullah M. Alharthi, and Marwa A. Gaheen. Optimizing feature selection and remote sensing classification with an enhanced machine learning method. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06790-
7.

Sravan:2025:SLS

[378] S. S. Sravan, Susmita Mandal, and P. J. A. Alphonse. SDSMS-LoRa: secure dynamic session key management scheme for LoRaWAN v1.1. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06802-
6.

Jiang:2025:FMP

[379] Rong Jiang, Yulin Li, Xuetao Pu, Xueke Wang, Wenyu Niu, and Zhiming Song. A fair multi-party contract signing scheme based on off-chain protocols and onchain smart contracts. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06844-
w.

Wang:2025:NAW

[380] Chenlu Wang, Xiaoyi Fu, Ziyi Yang, and Shenglin Li. NeuralWiGait: an accurate WiFi-based gait recognition system using hybrid deep learning framework. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06878-
                                        0.

Xu:2025:MCH

[381] Shuangfei Xu, Zhanjun Huang, Wenhao Bi, and An Zhang. A Monte Carlo hyper-heuristic algorithm with low-level heuristics reward prediction for missile path planning. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06771-
                                        w.

Wang:2025:LLA

[382] Peipeng Wang, Xiuguo Zhang, and Zhiying Cao. LogSD: log anomaly detection via topic words awareness semantic augmentation and category-guided Mixup data augmentation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06850-
                                        y.

Wu:2025:DDA

[383] Heng Wu, Zijun Zheng, Laishui Lv, Changchun Zhang, Dalal Bardou, Shanzhou Niu, and Gaohang Yu. Dara: distribution-aware representation alignment for semi-supervised domain adaptation in image classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06886-0.

Qi:2025:IMO

[384] Mingjun Qi, Xiaochun Wu, Keke Li, and Fenghao Yang. IPAQ: a multiobjective global optimal and time-aware task scheduling algorithm for fog computing environments. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06853-
9.

Jiang:2025:BMB

[385] Rong Jiang, Hejiao Zhang, Zhiming Song, Shenghu Tian, and Wenlu Lou. T-BFL model based on two-dimensional trust and blockchain-federated learning for medical data sharing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06873-
5.

Darabkh:2025:NGR

[386] Khalid A. Darabkh, Mamoun F.

Al-Mistarihi, and Mera Ismail Al-Maaitah. Next-generation routing for autonomous vehicle networks based on innovative clustering: integrating SDN and fog computing along with AODV upon failure. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06880-
6.

Xing:2025:EEE

[387] Hongjia Xing, Feng Yang, Xu Qiao, Fanruo Li, and Xinxin Huang. Enhanced end-to-end regression algorithm for autonomous road damage detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06871-
7.

Lin:2025:HMD

[388] Ching-Sheng Lin. A hybrid model for the detection of multi-agent written news articles based on linguistic features and BERT. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06882-
4.

Chen:2025:KNI

[389] Bo Chen, Yulin Zhang, Yunming Wang, Rui Tong, Yufeng Chen, Lingdong Sun, and Wenxue Xie. Key node identification method for twolayer fusion networks based on improved structural holes. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06843-
x.

Yu:2025:CAD

[390] Dongxian Yu, Xiaoyu Zhou, Ali Noorian, and Mehdi Hazratifard. Correction to: An AI-driven social media recommender system leveraging smartphone and IoT data. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06857-

5. See [272].

Zhang:2025:COY

[391] Yunjie Zhang, Guofeng Gao, Yadong Chen, and Zhenjian Yang. Correction: ODD–YOLOv8: an algorithm for small object detection in UAV imagery. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06829-

9. See [202].

Geng:2025:QEM

[392] Renxuan Geng, Yuang Guo, Guotao Wang, Yuansong Liu, Bingze Lv, Hui Wang, and Songyi Yu. A qualitative evaluation method for the weight of moving objects within a sealed cavity based on time-frequency spectrogram features. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06717-
2.

Yang:2025:ETT

[393] Zhimi Yang and Bo Shen. Estimating textual treatment effect via causal disentangled representation learning. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06887-z.

Xia:2025:MMS

[394] Pusen Xia, Shengwei Tian, Long Yu, Xin Fan, Zhezhe Zhu, Hualong Dong, Na Qu, Tong Liu, and Xiao Yuan. Mdcsnet: multi-scale dynamic spatial information fusion with criticality sampling for point cloud classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06838-
                                        8.

Simen:2025:EEQ

[395] Anton Simen, Rodrigo Bloot, Otto M. Pires, and Erick G. Sperandio Nascimento. Evolutionary-enhanced quantum supervised learning model. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06875-
                                        3.

Sheng:2025:CTD

[396] Jinfang Sheng, Yifan Zhang, and Bin Wang. Continuous-time dynamic graph learning based on spatiotemporal random walks. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06881-
                                        5.

Li:2025:DSC

[397] Meng Li, Bo Yang, Tao Xue, and Shaowei Han. Deep subspace clustering using dual self-expressiveness and convolutional fusion. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06885-
1.

Sultana:2025:ROQ

[398] Afrin Sultana and Edgard Muñoz-Coreas. Resource optimized quantum squaring circuit. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06876-
2.

Idrees:2025:SAL

[399] Sara Kadhum Idrees, Joseph Azar, Raphaël Couturier, Ali Kadhum Idrees, and Franck Gechter. SZ4IoT: an adaptive lightweight lossy compression algorithm for diverse IoT devices and data types. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06667-
9.

Zhang:2025:NSD

[400] Yunsong Zhang, Yuejuan Han, Jianfeng Jiang, and Lantao You. Nodeto-set disjoint paths problem in divideand-swap cube. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06895-
z.

Khala:2025:EML

[401] Mohamed Khala, Naima El Yanboiy, Ismail Elabbassi, Omar Eloutassi, Mohammed Halimi, Youssef El Hassouani, and Choukri Messaoudi. Enhancing machine learning model for early warning in PV plants: air temperature prediction informed by power temperature coefficient. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06909-
                                        w.

Osmani:2025:FSD

[402] Nooshin Osmani, Erfan Esmaeeli, and Sorayya Rezayi. Fusion strategies for deep convolutional neural network representations in histopathological image classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06663-
                                        z.

Wang:2025:CCCa

[403] Bo Wang, Rong Jiang, Xuetao Pu, and Hejiao Zhang. An on-chain and off-chain collaborative data sharing and access control model for electronic medical records. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06884-

Li:2025:DLP

[404] Benchao Li, Ting Wang, and Ruisheng Ran. Discriminant locality preserving projection on Grassmann Manifold for image-set classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06904-
1.

Zong:2025:PHT

[405] Jixiang Zong, Jiulong Wang, Guirun Li, Ruopu Wu, and Di Zhao. Polaris 23: a high throughput neuromorphic processing element by RISC-V customized instruction extension for spiking neural network (RV-SNN 2.0) and SIMDstyle implementation of LIF model with backpropagation STDP. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06826-
y.

Anonymous:2025:JNh

[406] Anonymous. Journal navigation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNi

[407] Anonymous. Journal navigation. The Journal of Supercomputing, 81(1):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rafieinejad:2025:IAC

[408] Mohammadreza Rafieinejad, Mohammadreza Binesh Marvasti, Seyyed Amir Asghari, and Kimiya Shahbakhti. Improve accuracy in CNNs while using approximate computing methods. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06901-4.

Lu:2025:DDH

[409] Jinkang Lu, Meng Lv, Peixuan Li, Zhu Yuan, and Ping Xie. Dhcache: a dual-hash cache for optimizing the read performance in key–value store. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06828-w.

Tourani:2025:STR

[410] Azadeh Navaei Tourani, Hamid Haj Seyyed Javadi, Hamidreza Navidi, and Arash Sharifi. A study on trust-rating mechanism for WSN node sensors using evolutionary game theory. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06824-
0.

Kumar:2025:FRB

[411] Shashi Shekhar Kumar, Ritesh Chandra, Anurag Harsh, and Sonali Agarwal. Fuzzy rule-based intelligent cardiovascular disease prediction using complex event processing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06911-
2.

Chahardoli:2025:ECP

[412] Meysam Chahardoli, Nafiseh Osati Eraghi, and Sara Nazari. An energy consumption prediction approach in smart cities by CNN–LSTM network improved with game theory and Namib Beetle Optimization (NBO) algorithm. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06811-
5.

Alasiry:2025:CSV

[413] Areej Alasiry and Mohammed Qayyum.

Correction: A smart vista-lite system for anomaly detection and motion prediction for video surveillance in vibrant urban settings. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06877-

1. See [297].

Tang:2025:CCE

[414] Yiou Tang, Yan Ma, Chunling Xiao, Min Wu, and Guoyuan Zeng. Correction: Classification of EEG eventrelated potentials based on channel attention mechanism. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06830-

2. See [126].

Jin:2025:RGN

[415] Yan Jin, Haoyu Shi, and Huaiye Meng. Robust graph neural networks based on feature fusion. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06917-
                                        4.

Zhang:2025:MLT

[416] Qian Zhang, Chenghao Ji, Mingwen Shao, and Hong Liang. MEKF: longtailed visual recognition via multiple experts with knowledge fusion. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        06920-9.

Zhang:2025:EVT

[417] Qian Zhang, Zuosui Yang, Mingwen Shao, and Hong Liang. An efficient video transformer network with token discard and keyframe enhancement for action recognition. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06927-
                                        2.

Liang:2025:MLN

[418] Hong Liang, Xian Li, Mingwen Shao, and Qian Zhang. Multi-level navigation network: advancing fine-grained visual classification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06933-
4.

Ferreira:2025:UDL

[419] Fernando Rodrigues Trindade Ferreira and Loena Marins do Couto. Using deep learning on microscopic images for white blood cell detection and segmentation to assist in leukemia diagnosis. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06903-
2.

Wang:2025:DSM

[420] Yan Wang, Zheng Gong, Dayu Jia, Aiping Tan, and Minchao Liu. Dynamic sharding model and performance optimization method for consortium blockchain. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06870-
8.

Huang:2025:MTS

[421] Nana Huang, Hongwei Ding, Ruimin Hu, Pengfei Jiao, Zhidong Zhao, Bin Yang, and Qi Zheng. Multi-time-scale with clockwork recurrent neural network modeling for sequential recommendation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06925-
4.

Liu:2025:TTD

[422] Tao Liu, Chenyoukang Lin, Yunteng Hu, Ruyi Cao, and Wendong Zhang. Traffic target detection based on context enhancement and feature purification. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06944-
1.

Sun:2025:LLA

[423] Yuheng Sun, Zhenping Lan, Yanguo Sun, Yuepeng Guo, Xinxin Li, Yuru Wang, and Bo Li. Ldstd: low-altitude drone aerial small target detector. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06950-
3.

Li:2025:SDD

[424] Biao Li, Bing Wang, Xiong Hu, Jianhui Zhai, and Changping Ji. A small defect detection technique for industrial product surfaces based on the EA–YOLO model. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06929-
0.

Chen:2025:MMS

[425] Wuqi Chen, Junjie Ye, Chunna Zhao, and Yaqun Huang. MFFCNN: multiscale fractional Fourier transform convolutional neural network for multivariate time series forecasting. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06888-
y.

Xue:2025:EOW

[426] Wenjin Xue, Guowei Xu, Nan Yang, and Jian Liu. Enhancing openworld object detection with AIGCgenerated datasets and elastic weight consolidation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06910-
3.

Cardama:2025:RIR

[427] F. Javier Cardama, Jorge Vázquez-Pérez, César Piñeiro, Juan C. Pichel, Tomás F. Pena, and Andrés Gómez. Review of intermediate representations for quantum computing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06892-

Khah:2025:CHM

[428] Yashar Pourardebil Khah, Mirsaeid Hosseini Shirvani, and Homayun Motameni. Correction: A hybrid machine learning approach for feature selection in designing intrusion detection systems (IDS) model for distributed computing networks. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06869-

1. See [254].

Huang:2025:PPL

[429] Qiongyan Huang, Yuhan Xia, Yunfei Long, Hui Fang, Ruiwei Liang, Yin Guan, and Ge Xu. Prompt4LJP: prompt learning for legal judgment prediction. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06945-
                                        0.

Peng:2025:TPP

[430] Tianqi Peng, Bei Gong, and Pengxuan Sun. Toward privacy-preserving verifiable DSSE for attribute-based cloud computing system. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06912-
                                        1.

Eufracio:2025:RBF

[431] Alfredo Cisneros Eufracio, Roberth Saenz Perez Alvarado, Jimmy Aurelio Rosales Huamani, Uwe Rojas Villanueva, Jose Luis Castillo Sequera, and Jose Manuel Gomez Pulido. Rock block fall prediction prototype by structural control applied to slopes using quantum machine learning (QML). The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06913-
0.

Alahmed:2025:ADL

[432] Hiba A. Alahmed and Ghaida A. Al-Suhail. AlzONet: a deep learning optimized framework for multiclass Alzheimer’s disease diagnosis using MRI brain imaging. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06924-
5.

Mishra:2025:CWG

[433] Sanjukta Mishra, Samarjit Kar, and Parag Kumar Guhathakurta. Cloud-WAVECAP: Ground-based cloud types detection with an efficient waveletcapsule approach. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06941-
4.

Li:2025:LSQ

[434] Ruoxia Li, Linli Si, and Jinde Cao.

Lagrange stability of quaternion-valued memristive neural networks on time scales: linear optimization method.

The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06914-z.

Darabkh:2025:LFC

[435] Khalid A. Darabkh, Mamoun F. Al-Mistarihi, and Bayan Abdallah Odat. Leveraging fog computing and softwaredefined networking for a novel velocityaware routing protocol with election and handover thresholds in VANETs. The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06883-3.

Lopez:2025:IEP

[436] Marta López, Esteban Stafford, and Jose Luis Bosque. Intelligent energy pairing scheduler (InEPS) for heterogeneous HPC clusters. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06907-
                                        y.

Campos:2025:EDF

[437] Cristian Campos, Rafael Asenjo, and Angeles Navarro. Exploring data flow design and vectorization with oneAPI for streaming applications on CPU+GPU. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06891-
3.

Liang:2025:MMS

[438] Hong Liang, Cuiping Wang, Mingwen Shao, and Qian Zhang. MAQT: multiscale attention and query-optimized transformer for end-to-end pose estimation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06923-
6.

Liu:2025:CPS

[439] Jianhang Liu, Chunxing Xia, Xuerong Cui, and Haibo Wu. CPCS: a perception sharing scheme of vehicle-road cooperation based on cybertwin. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06939-
y.

Li:2025:DDM

[440] Pei Li, Tong Bai, Xiaoying Pan, and Chengyu Zuo. Defocus deblur method of multi-scale depth-of-field cross-stage fusion image based on defocus map forecast. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06934-
3.

Safi-Esfahani:2025:LWH

[441] Faramarz Safi-Esfahani, Leili Mohammadhoseini, Habib Larian, and Seyedali Mirjalili. LEVYEFO-WTMTOA: the hybrid of the multi-tracker optimization algorithm and the electromagnetic field optimization. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06856-
6.

Yang:2025:MRS

[442] Xiaonan Yang, Zuoxi Zhao, Kai Yuan, Can Xiao, and YangFan Luo. Mobile recognition system for multinational currencies based on CA–DSC-RepVGG algorithm. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06936-
1.

Zhang:2025:IJA

[443] Qiwen Zhang and Tian Zhen. Improved Jaya algorithm for energy-efficient distributed heterogeneous permutation flow shop scheduling. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06938-
z.

Zangeneh:2025:NSC

[444] Iman Zangeneh, Amir Massoud Bidgoli, and Ardeshir Dolati. Novel solution of the controller and backup controller placement problem for improving reliability in IoT-based data monitoring systems. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06846-
8.

Nascimento:2025:AAS

[445] Erick Nascimento, Luan Lins, Eduardo Tavares, Paulo Pereira, Jamilson Dantas, Sokol Kosta, and Paulo Maciel. Availability assessment of SDN–ICN service for multi-access edge computing. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06879-
z.

Prasad:2025:BDG

[446] Chandrabhushan Prasad, Sri Khetwat Saritha, and Sweta Jain. Bilinear diffusion graph convolutional network model for social recommendation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06930-
7.

Zaimi:2025:EMM

[447] Rania Zaimi, Khouloud Safi Eljil, Mohamed Hafidi, Mahnane Lamia, and Farid Nait-Abdesselam. An enhanced mechanism for malicious URL detection using deep learning and DistilBERTbased feature extraction. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06908-
                                        x.

Wu:2025:GHF

[448] Chengmao Wu and Siyu Zhou. Generalized harmonic fuzzy partition Cmeans clustering. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06723-
                                        4.

Li:2025:MOE

[449] Wei Li, Wenhao Tang, and Lei Wang. Many-objective evolutionary algorithm based on dynamic mating and strengthened fitness selection mechanism. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-024-
                                        06821-3.

Dong:2025:APB

[450] Jian Dong, Wei Bao, Xiaoqi Cao, Yang Xu, Yuze Yang, Binbin Li, Qi Zhang, and Heng Ye. AIS-Bench: an performance benchmark for AI server systems. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06778-
                                        3.

Vazquez-Perez:2025:IIC

[451] Jorge Vázquez-Pérez, F. Javier Cardama, César Piñeiro, Juan C. Pichel, Tomás F. Pena, and Andrés Gómez. Inqasm: InQuIR compiler to NetQASM.

The Journal of Supercomputing, 81(2): ??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
06955-y.

Aymaz:2025:UPO

[452] Samet Aymaz. Unlocking the power of optimized data balancing ratios: a new frontier in tackling imbalanced datasets. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06919-
2.

Xiao:2025:MDD

[453] Junbi Xiao, Ruifeng Sun, and Jianhang Liu. MLDDoS: a distributed denial of service attack detection method using multi-level sketch. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06942-
3.

Tondreau:2025:SDH

[454] Pierre-Simon Callist Yannick Tondreau, Juan C. Perez, Juan P. Díaz, Vicente Blanco Pérez, and Jonatan Felipe. Singularity to deploy HPC applications: a study case with WRF. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06893-

Alkhudhayr:2025:MCR

[455] Hanadi Alkhudhayr and Hanin Ardah. Mitigating cyberphysical risks in IoT-enabled smart transport infrastructure. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06948-
                                        x.

Kong:2025:CRS

[456] Dekun Kong and Wenguang Yang. Correction: Research on site selection of emergency material reserve based on set pair analysis and TOPSIS integration method: a case study of Hebei Province, China. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06971-

y. See [358].

Huang:2025:LHD

[457] Hong Huang, Wengang Luo, Yunfei Wang, Yinghang Zhou, and Weitao Huang. LogCTBL: a hybrid deep learning model for log-based anomaly detection. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06926-
                                        3.

Anonymous:2025:JNj

[458] Anonymous. Journal navigation. The Journal of Supercomputing, 81(2):??, January 2025. CODEN JOSUED.

ISSN 0920-8542 (print), 1573-0484 (electronic).

Fu:2025:MMS

[459] Youjia Fu, Zihao Xu, Junsong Fu, Huixia Xue, Shuqiu Tan, Lei Li, and Shaoxun Qing. Monomm: a Multi-scale Mamba-Enhanced network for real-time monocular 3D object detection. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
06946-z.

Pastor:2025:CDB

[460] Alfred M. Pastor, Jose M. Badia, and Maribel Castillo. A community detection-based parallel algorithm for quantum circuit simulation using tensor networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06918-
3.

Suarez:2025:KCB

[461] Daniel Suárez, Francisco Almeida, Vicente Blanco, and Pedro Toledo. KubePipe: a container-based high-level parallelization tool for scalable machine learning pipelines. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06956-

Raul:2025:CGC

[462] Nozal Raúl and Jose Luis Bosque. CPU-GPU co-execution through the exploitation of hybrid technologies via SYCL. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06963-
                                        y.

Yanamala:2025:FAH

[463] Rama Muni Reddy Yanamala and Muralidhar Pullakandam. FPGAaccelerated hybrid CNN–LSTM system for efficient EEG-based drowsiness recognition. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06947-
                                        y.

Zhang:2025:FIE

[464] Xuncai Zhang, Mengrui Liu, and Ying Niu. Facial image encryption scheme based on improved 4-D hyperchaotic system. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06943-
                                        2.

Roh:2025:HQC

[465] Emily Jimin Roh, Joo Yong Shim, Joongheon Kim, and Soohyun Park. Hybrid quantum-classical 3D object detection using multi-channel quantum convolutional neural network. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06968-
7.

Zhou:2025:SRF

[466] Sizhong Zhou, Zhiren Sun, and Yuli Zhang. Spectral radius and kfactor-critical graphs. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06902-
3.

Zhao:2025:PSM

[467] Shengsheng Zhao, Yantao Yu, Tiancong Huang, Guojin Liu, and Yucheng Wu. Pricing strategies in mobile crowdsensing: an enhanced MAPPO approach using a behavior network. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06957-
w.

Liu:2025:MSE

[468] Tianbao Liu and Zhe Feng. Multistrategy enhanced dandelion optimizer based on elliptic approximation strategy and adaptive fitness-distance-similarity balance for solar photovoltaic parameter estimation. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06899-
9.

Xiao:2025:JCO

[469] Qimu Xiao and Mingyu Xiao. Joint computation offloading and resource allocation in multi-cell MEC networks. The Journal of Supercomputing, 81(3): ??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        06921-8.

Xiong:2025:ORT

[470] Zhi Xiong, Linhui Tan, Jianlong Xu, and Lingru Cai. Online realtime energy consumption optimization with resistance to server switch jitter for server clusters. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06827-
                                        x.

Zhang:2025:NSS

[471] Shengcai Zhang, Zhiying Fu, Dezhi An, and Huiju Yi. Network security situation assessment based on BKA and cross dual-channel. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06932-
                                        5.

Nguyen:2025:CBA

[472] Da-Thao Nguyen, Thanh-Phuong Nguyen, and Ming-Yuan Cho. Cloud-based AIoT intelligent infrastructure for firefighting pump fault diagnosis-based hybrid CNN–GRU deep learning technique. The Journal of Supercomputing, 81(3):

??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
06965-w.

Ismail:2025:OSB

[473] Nabil A. Ismail, Shaimaa Abu Khadra, Gamal M. Attiya, and Salah Eldin S. E. Abdulrahman. Optimizing SIKE for blockchain-based IoT ecosystems with resource constraints. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06906-
z.

Atencia:2025:PQO

[474] Javier Ruiz Atencia, Otoniel Mario López Granado, Manuel Pérez Malumbres, and Miguel Onofre Martínez-Rach. Perceptual QP optimization for VVC with dual hybrid neural networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06954-
z.

Nunez-Chongo:2025:CDD

[475] Osiris Núñez-Chongo, Hernán Asorey, Antonio Juan Rubio-Montero, Mauricio Suárez-Durán, Rafael Mayo-García, and Manuel Carretero. Convergent data-driven workflows for open radiation calculations: an exportable methodology to any field. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-024-06894-
                                        0.

Wen:2025:VOD

[476] Jia Wen, Qi Zhang, and Guanghao Zhang. VPSNet: 3D object detection with voxel purification and fully sparse convolutional networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06890-
                                        4.

Yang:2025:SVI

[477] Qimeng Yang, Hao Meng, Yuanbo Yan, Shisong Guo, and Qixing Wei. SFVE: visual information enhancement metaphor detection with multimodal splitting fusion. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06958-
                                        9.

Kim:2025:PEI

[478] Suhri Kim, Youngdo Lee, and Kisoon Yoon. Performance evaluation of isogeny-based digital signature algorithms: introducing FIBS-fast isogenybased digital signature*. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06970-
                                        z.

Roselin:2025:DDL

[479] J. Roselin and Israelin J. Insulata. Decentralized dynamic load balancing for virtual machines in cloud computing: a blockchain-enabled system with state channel optimization. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06922-
7.

Zheng:2025:CGG

[480] Zhiwen Zheng, Xiaoyun Chen, and Musheng Huang. Cluster-guided graph attention auto-encoder. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06953-
0.

Zhou:2025:PLA

[481] Zhuguo Zhou, Yujun Lu, and Liye Lv. Pallet localization algorithm based on improved human pose estimation with transfer learning. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06973-
w.

Vicente-Jaen:2025:VRT

[482] Arturo Vicente-Jaén, Juan Mompeán, Juan L. Aragón, and Pablo Artal. Vol-Rec: 4D real-time volumetric reconstruction of OCT data. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06969-

Fang:2025:DRC

[483] Juan Fang, Haoyu Cheng, Yuening Wang, and Ran Zhai. DRCD: a regional-contention-driven arbitration policy for CPU–GPU heterogeneous systems. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07001-
                                        7.

Kirubavathi:2025:DMT

[484] G. Kirubavathi, I. R. Sumathi, J. Mahalakshmi, and Durgesh Srivastava. Detection and mitigation of TCP-based DDoS attacks in cloud environments using a self-attention and intersample attention transformer model. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06940-
                                        5.

Tolun:2025:CBM

[485] Ömer Can Tolun, Kasim Zor, and Onder Tutsoy. A comprehensive benchmark of machine learning-based algorithms for medium-term electric vehicle charging demand prediction. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06975-
                                        8.

Wang:2025:ERD

[486] Youwei Wang, Lizhou Feng, and Yan Zhang. Early rumor detection method based on stage sampling and triplerelationship graph. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06959-
8.

Cui:2025:PES

[487] Mengqi Cui, Jinchen Xu, Yuchang Zhou, Hongru Yang, Liguang Ji, and Bei Zhou. PESA: error sensitivity analysis tool for floating-point computational programs. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06962-
z.

Shao:2025:RFF

[488] Changshun Shao, Zhenglin Yu, Hongchang Ding, Guohua Cao, Jingsong Duan, and Bin Zhou. A random flightfollow leader and reinforcement learning approach for flexible job shop scheduling problem. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06935-
2.

Hasani:2025:MQP

[489] Mehri Hasani, Masoud Ghods, Sourav Mondal, Muhammad Kamran Siddiqui, and Imran Zulfiqar Cheema. Modeling QSPR for pyelonephritis drugs: a topological indices approach using MATLAB. The Journal of Supercomputing, 81(3):??, February 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06967-
                                        8.

Cao:2025:CFM

[490] Zhipeng Cao, Nianmin Yao, Linqi Meng, Jingyi Fang, and Jian Zhao. Coarse-to-fine medical image registration with landmarks and deformable networks. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07000-
                                        8.

Zhang:2025:SSA

[491] Yuxin Zhang, Jiazheng Wen, Ran Wu, Huanyu Liu, and Junbao Li. Siamsdt: a self-adaptive dynamic template siamese network for airborne visual tracking of MAVs on heterogeneous FPGA–SoC. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06928-
                                        1.

Mashayekhi:2025:DEE

[492] Negin Mashayekhi, Ghassem Jaberipur, Mohammad Reza Reshadinezhad, and Shekoofeh Moghimi. Design of energyefficient and high-speed hybrid decimal adder. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-024-06897-
x.

He:2025:IDM

[493] Fuchun He, Chunming Fu, Youwei He, Shaoyong Huo, Jiachang Tang, and Xiangyun Long. Improved dwarf mongoose optimization algorithm based on hybrid strategy for global optimization and engineering problems. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06931-
6.

Nieves-Perez:2025:CEE

[494] Isidoro Nieves-Pérez, Alfonso Muñoz, Francisco Almeida, and Vicente Blanco. Correction: Energy efficiency and performance analysis of a legacy atomic scale materials modeling simulator (VASP). The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06976-
7.

Narwaria:2025:RER

[495] Abhishek Narwaria, Varsha Kumari, and Arka Prokash Mazumdar. RL-EAR: reinforcement learning-based energyaware routing for software-defined wireless sensor network. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06998-
1.

Lan:2025:EIA

[496] Yu Lan, Hui Huang, Zhenjie Huang, Qunshan Chen, and Shuaike Wu. ERBFT: improved asynchronous BFT with erasure code and verifiable random function. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06995-
                                        4.

Qiu:2025:MRC

[497] Hong Qiu, Wentao Yu, Gan Zhang, Xuan Xia, and Kun Yao. Multi-robot collaborative 3D path planning based on game theory and particle swarm optimization hybrid method. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06960-
                                        1.

Zhang:2025:TBF

[498] Dabin Zhang, Zehui Yu, Zhimei Zeng, Boting Zhang, Ruibin Lin, and Huanling Hu. A text-based framework for carbon price forecasting via multivariate temporal graph neural network. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06974-
                                        9.

Tang:2025:CFS

[499] Yu Tang, Qi Dai, Ye Du, Tian shuai Zheng, and Mei hong Li. Capsule feature selector for software defect prediction. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06949-
w.

Shao:2025:SSS

[500] Yunxue Shao, Zhiyang Wang, and Lingfeng Wang. SSIM: self-supervised learning method based on spatially selected shifts and irregular image masking. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07013-
3.

Meng:2025:GAM

[501] Yi Meng, Nurbol Luktarhan, Xiaotong Yang, and Guodong Zhao. GBADroid: an Android malware detection method based on multi-view feature fusion. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06977-
6.

Zhang:2025:UAP

[502] Heng Zhang, Yang Peng, and Yan li Liu. UAV aerial photography target detection based on improved YOLOv9. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-025-06991-
                                        8.

Dang:2025:EPP

[503] Lanxue Dang, Zan Li, Shilong Li, Baojun Qiao, and Liming Zhou. Effective plug-and-play lightweight modules for YOLO series models. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07026-
                                        y.

Ma:2025:TTF

[504] Siwei Ma, Ronghua Li, and Henan Hu. Train track fastener defect detection algorithm based on MGSF– YOLO. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07024-
                                        0.

Wan:2025:VFA

[505] Junhang Wan, Yanping Chen, and Cong Gao. The VAE–FastGA anomaly detection model based on subspace and weakly correlated ultra-high-dimensional data. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06988-
                                        3.

Zhang:2025:EQG

[506] Yu Zhang, Lu Lu, Rong Zhao, Yijie Guo, and Zhanyu Yang. An efficient quantized GEMV implementation for large language models inference with matrix core. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06993-
6.

Chen:2025:OCA

[507] Zigang Chen, Zhenjiang Zhang, Tao Leng, Haihua Zhu, and Yuhong Liu. Outsourcing collaboration analysis of multiparty privacy data using the improved Yannakakis. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06994-
5.

Wu:2025:QRB

[508] Faguo Wu, Bo Zhou, Jiale Song, and Lijia Xie. Quantum-resistant blockchain and performance analysis. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07018-
y.

Bao:2025:GCG

[509] Qingda Bao, Shengfa Miao, Yulin Tian, Xin Jin, Puming Wang, Qian Jiang, Shaowen Yao, Da Hu, and Ruoshu Wang. GDRNet: a channel grouping based time-slice dilated residual network for long-term timeseries forecasting. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07011-
                                        5.

Ma:2025:GAT

[510] Minyi Ma, Hongfang Gong, and Yingjing Ding. Graph affine transformer with a symmetric adaptation strategy for text classification. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06992-
                                        7.

Luo:2025:AFD

[511] Mao Luo, Xianhong Liu, Xinyun Wu, Caiquan Xiong, and Yuanzhi Ke. An adaptive focal distance tabu search approach for the minimum 2-connected dominating set problem. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07014-
                                        2.

Basarslan:2025:MCB

[512] Muhammet Sinan Basarslan. M-C&M-BL: a novel classification model for brain tumor classification: multi-CNN and multi-BiLSTM. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-06964-
                                        x.

Esfandiari:2025:IIN

[513] Shima Esfandiari and Seyed Mostafa Fakhrahmad. Identifying influential nodes in complex networks by adjusted feature contributions and neighborhood impact. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06645-
1.

Zhong:2025:PPR

[514] Jie Zhong, Juan Chen, Zongling Wu, Peng Chen, Di Tian, and Xi Li. P2PPO: parallel residual network and prioritized experience replay enhanced PPO for task offloading and resource allocation in SatEC. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06900-
5.

Zhou:2025:AYS

[515] Wentao Zhou, Chengtao Cai, Sutthiphong Srigrarom, Pengfei Wang, Zijian Cui, and Chenming Li. ADH-YOLO: a small object detection based on improved YOLOv8 for airport scene images in hazy weather. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06999-
0.

Anonymous:2025:JNk

[516] Anonymous. Journal navigation. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Anonymous:2025:JNl

[517] Anonymous. Journal navigation. The Journal of Supercomputing, 81(3):??, February 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

AlDawi:2025:ABA

[518] Abdelaziz Al Dawi, Necmi Serkan Tezel, Javad Rahebi, and Ayhan Akbas. An approach to botnet attacks in the fog computing layer and Apache Spark for smart cities. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06915-
                                        y.

Yu:2025:MSM

[519] Changhong Yu and Mingxuan Zhang. M^{2}SSCENet: a multi-branch multi-scale network with spatial-spectral crossenhancement for hyperspectral and Li-DAR data classification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07096-
                                        y.

Zhu:2025:HHS

[520] Heming Zhu, Hao Li, and Gehao Lu. HSVDetector: a heterogeneous semantic graph-based method for smart contract vulnerability detection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/

article/10.1007/s11227-025-06951-
2.

Jiang:2025:EEC

[521] Chenyang Jiang, Zhendong Li, Jun Yang, Yiqiang Wu, and Shuai Li. Effective and efficient conditional contrast for data-free knowledge distillation with low memory. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07115-
y.

Niu:2025:MFF

[522] Yakun Niu, Xiangru Chen, and Hongjian Yin. Median filtering forensics using spatial and frequency domain residuals. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07099-
9.

Tong:2025:SYS

[523] Haoyang Tong, Dongyang Gao, Zhixu Wang, Longlong Feng, Yue Li, and Xuewei Bai. SGB-YOLOv5: straw granulator blockage monitoring system. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07079-z.

Xia:2025:STS

[524] Wenyuan Xia, Qing Zhou, Dayu Wu, Siyuan Wang, and Mengshuang Zhou. A scalable two-stage model for real-time wetland bird recognition. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07061-
9.

Li:2025:MNM

[525] Peng Li, Zhanao Huang, Yan Li, Yi Kou, and Xiaojie Li. Mfe-net: a multiscale feature enhanced network for mesoscale convective systems identification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07065-
5.

Qin:2025:LSM

[526] Chuandong Qin, Yiqing Zhang, and Yu Cao. Large-scale machine learning with synchronous parallel adaptive stochastic variance reduction gradient descent for high-dimensional blindness detection on Spark. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07046-
8.

Ye:2025:GIG

[527] Pengdong Ye, Yanhua Liang, Yutao Bie, Guihe Qin, Jiaru Song, Yingqing Wang, and Wanning Liu. GDT-IDS: graph-based decision tree intrusion detection system for controller area network. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07116-
x.

Gheibi-Fetrat:2025:SSS

[528] Atiyeh Gheibi-Fetrat, Fatemeh Serajeh Hassani, Masoud Mohammadi-Lak, Amir Mirzaei, Negar Akbarzadeh, Mahmoud Reza Kheyrati-Fard, Mohammad Hosseini, Ahmad Javadi Nezhad, Arash Tavakkol, Jeong A. Lee, and Hamid Sarbazi-Azad. A survey of SSD simulators and emulators. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07078-
0.

Li:2025:HPA

[529] Runhua Li, Qinglin Wang, and Jie Liu.

A heterogeneous parallel algorithm for the Cartesian discrete ordinates for multizone heterogeneous system. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07087-
z.

Wang:2025:AAP

[530] Ziheng Wang, Zixuan Shao, Baowei Wang, and Xu Cheng. APPBoost: an adaptive parameter pair boosting algorithm for enhanced robustness against noise and imbalance. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-025-07053-
                                        9.

Li:2025:DSS

[531] Shouliang Li, Rudong Min, Jilong Zhang, Jiale Han, Yulin Shen, Zhen Yang, and Yi Yang. A digital signature scheme based on general Chebyshev polynomial. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07074-
                                        4.

Sun:2025:KEN

[532] Pengfei Sun, Zhiping Wang, Liyan Jia, and Xiaoxi Wang. KSIPF: an effective noise filtering oversampling method based on k-means and iterative-partitioning filter. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07081-
                                        5.

Rezaei:2025:MNH

[533] Azita Rezaei, Ali Broumandnia, and Seyed Javad Mirabedini. MLS: a novel hybrid security framework utilizing the Wiedemann algorithm and chaotic mapping for MQTT. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06858-
                                        4.

Lawrence:2025:DSH

[534] Linju Lawrence and R. Shreelekshmi.

Double salted HMAC signature with blockchain for faster and secure video integrity verification. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06996-
3.

Zhang:2025:DMO

[535] Zhibo Zhang, Huiqiang Wang, Rongqiang Li, Hongwu Lv, and Dongmiao He. DQN-MSRA: an online SFC deployment method based on multistep reinforcement learning. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07042-
y.

Turkmen:2025:SSC

[536] Muhammed Türkmen, Canan Çiftçi, and Gülnaz Boruzanli Ekinci. Structure and substructure connectivity of folded divide-and-swap cube. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07100-
5.

delaRosa:2025:QSP

[537] Miguel Sánchez de la Rosa, Gabriel Gomez-Lopez, Francisco J. Andújar, Jesús Escudero-Sahuquillo, José L. Sánchez, Francisco J. Alfaro-Cortés, and Pierre-Axel Lagadec. Quality-of-service provision for BXIv3-based interconnection networks. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07069-
1.

Genc:2025:IAC

[538] Hasan Genç, Canan Koç, Esra Yüzgeç Özdemir, and Fatih Özyurt. An innovative approach to classify meniscus tears by reducing vision transformers features with elasticnet approach. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07103-
2.

Gao:2025:SCL

[539] Jianbang Gao, Yuxiao Guo, and Guowang Gao. Sim-ConvFormer: a lightweight fault diagnosis framework incorporating SimAM and external attention. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07075-
3.

Huang:2025:CFS

[540] Weidi Huang, Zhiqiang Yao, Biao Jin, Zheyu Chen, and Yue Wang. Controllable face soft-biometric privacy enhancement based on attribute disentanglement. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07134-
9.

Yan:2025:SCB

[541] Xiao Yan, Fan Xiumei, Kok-Lim Alvin Yau, Xie Zhixin, and Men Rui. Semantic communication based on bi-level routing attention in IoT environment. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07062-
8.

Fu:2025:DAD

[542] Honghao Fu and Yidong Chen. Data augmentation and debiasing for signers in signer-independent sign language translation. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07119-
8.

Yao:2025:DEG

[543] Ruiyang Yao. Dual-enhanced graph convolutional networks for aspect-based financial sentiment analysis. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06972-
x.

Qin:2025:SID

[544] Qiurong Qin, Yueqin Li, Yajie Mi, Jinhui Shen, Kexin Wu, and Zhenzhao Wang. Sql injection detection algorithm based on Bi–LSTM and integrated feature selection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07109-
w.

Yang:2025:MSG

[545] Wenbiao Yang, Tingfeng Lai, and Yuhui Fang. Multi-Strategy Golden Jackal Optimization for engineering design. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07106-z.

Zhang:2025:SGR

[546] Zhi-Yuan Zhang, Hao Ren, Hao Li, Kang-Hui Yuan, and Chu-Feng Zhu. Static gesture recognition based on thermal imaging sensors. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07140-
x.

Khanal:2025:DDG

[547] Bikram Khanal and Pablo Rivas. Datadependent generalization bounds for parameterized quantum models under noise. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06966-
9.

Siqueira:2025:APM

[548] Hadley Siqueira and Marcio Kreutz. AccPRET: a proposal of a multicore architecture with reconfigurable accelerators and time predictability. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07131-
y.

Shrestha:2025:SAO

[549] Susav Shrestha, Aayush Gautam, and Narasimha Reddy. Storage access optimization for efficient GPU-centric information retrieval. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07118-
9.

Saif:2025:ERC

[550] Abdu Saif, Nor Shahida Mohd Shah, Yahya M. Al-Moliki, Weiwei Jiang, and Saeed Hamood Alsamhi. Enhancing resilience communication in B5G: optimal deployment of tethered networked flying platforms for disaster recovery. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07098-
w.

Li:2025:AMA

[551] Yi Li, Wenying Zheng, and Sai Ji.

AI model auditing scheme towards cloud-edge high-performance computing. The Journal of Supercomputing, 81 (4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07121-0.

Li:2025:RMP

[552] Peng Li, Xinru Zhu, Xiaoshan Li, and Baofeng Huo. Research on mitigating popularity bias in federal recommendation based on users’ behavior. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07144-
7.

Chuang:2025:ENI

[553] Po-Jen Chuang and Pang-Yu Huang. Enhancing network intrusion detection by employing Mondrian forests to achieve multiple attack classification. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07123-y.

Pasini:2025:STT

[554] Massimiliano Lupo Pasini, Jong Youl Choi, Kshitij Mehta, Pei Zhang, David Rogers, Jonghyun Bae, Khaled Z. Ibrahim, Ashwin M. Aji, Karl W. Schulz, Jordà Polo, and Prasanna Balaprakash. Scalable training of trustworthy and energy-efficient predictive graph foundation models for atomistic materials modeling: a case study with HydraGNN. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07029-
9.

Bozorgi:2025:SRR

[555] Elika Bozorgi, Sakher Khalil Alqaaidi, Afsaneh Shams, Hamid Reza Arabnia, and Krzysztof Kochut. A survey on the recent random walk-based methods for embedding graphs. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07019-
x.

Anjali:2025:DCM

[556] Anjali Anjali and Anjana Gupta. Dynamic cloud model based on decision field theory. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-06989-

2. See correction [603].

Mestre:2025:DLI

[557] Jose I. Mestre, Sergio Barrachina, Darwin Quezada, and Manuel F. Dolz. Deep learning inference optimisation for IoT: Conv2D-ReLU-BN layer fusion and quantisation. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07107-
y.

Ning:2025:TYL

[558] Zhaole Ning, Tianze Zhang, Xin Li, Aiying Wu, and Gang Shi. Tp-yolov8: a lightweight and accurate model for traffic accident recognition. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07129-
                                        6.

Tang:2025:ADA

[559] Yayuan Tang, Suchi Mishra, Noha Alduaiji, Piyush Kumar Shukla, Mohammad Yahya, and Tao Pang. An advanced data analytics approach to a cognitive cyber-physical system for the identification and mitigation of cyber threats in the medical internet of things (MIoT). The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07093-
                                        1.

Zhang:2025:PIM

[560] Yulei Zhang, Guangli Zhu, Yuanyuan Ding, Zhongliang Wei, Lei Chen, and Kuan-Ching Li. A progressive interaction model for multimodal sarcasm detection. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07110-
                                        3.

Xu:2025:VBI

[561] Haixia Xu, Yuting Xu, and Kaiyu Hu. A vision-based inspection system for pharmaceutical production line. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07135-
8.

Wang:2025:ACR

[562] Zhihang Wang, Jiafei Liu, Chia-Wei Lee, Jingli Wu, and Gaoshi Li. An analysis on component reliability of (n, k)-star networks. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07128-
7.

Wu:2025:AFF

[563] Jun Wu, Yajing Zhang, Tengfei Shan, Zhiwei Xing, Jiusheng Chen, and Runxia Guo. An additive feature fusion attention based on YOLO network for aircraft skin damage detection. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07148-3.

Rafiee:2025:TMS

[564] Mojtaba Rafiee. Toward more secure constructions of flexible multiclient functional encryption schemes. The Journal of Supercomputing, 81(4): ??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07112-1.

Hoorfar:2025:TIE

[565] Hamid Hoorfar, Adam C. Puche, and Istvan Merchenthaler. Thermal image edge detection for AI-powered medical research imaging. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07050-

y. See publisher correction [616] about missing third author.

Zhou:2025:SYS

[566] Weina Zhou, Yuqi Yang, Ming Zhao, and Wenhua Hu. SCNet-YOLO: a symmetric convolution network for multi-scenario ship detection based on YOLOv7. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07120-
                                        1.

Khalid:2025:DLT

[567] Maira Khalid, Jehad Ali, Ahmed Raza Mohsin, Byeong hee Roh, and Mohammed J. F. Alenazi. Deep learning techniques for enhanced security and privacy in 6G terrestrial-nonterrestrial network architecture. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07097-
                                        x.

Anonymous:2025:JNm

[568] Anonymous. Journal navigation. The Journal of Supercomputing, 81(4):??, March 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Chai:2025:CCL

[569] Baoqiang Chai, Ruisheng Zhang, Xinyue Li, and Jianxin Tang. CLDE: a competitive learning-driven differential evolution optimization for the influence maximization problem in social networks. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07175-
0.

Zhang:2025:CTA

[570] Xiaoyong Zhang and Wei Yue. Collaborative task allocation for heterogeneous UAV coalitions with resource requirements based on multigenotype genetic algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07193-
y.

Song:2025:IYB

[571] Jian Song, Jie Xie, Qingwang Wang, and Tao Shen. An improved YOLObased method with lightweight C3 modules for object detection in resourceconstrained environments. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07187-
w.

Ying:2025:MSI

[572] Chenghao Ying, Zhiping Zhou, Kewei Li, Zhaozhong Zhang, and Qingshuang Yang. Multi-stream information complementarity network for RGB-D camouflaged object detection. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07201-
                                        1.

Aghajani:2025:ESE

[573] Khadijeh Aghajani and Mahbanou Zohrevandi. Enhancing speech emotion recognition: a deep learning approach with self-attention and acoustic features. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07166-
                                        1.

Bezerra:2025:CBS

[574] Diego de Freitas Bezerra, Elisson da Silva Rocha, Guto Leoni Santos, André Luis Cavalcanti Moreira, Judith Kelner, Djamel F. H. Sadok, Glauco Estácio Gonçalves, Maria Valéria Marquezini, and Patricia Takako Endo. A cluster-based solution for service function chain allocation in large-scale infrastructure. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07218-
                                        6.

Zhang:2025:POM

[575] Zhongliang Zhang, Tao Liu, Chengzhi Wang, Ying Guo, Jingshan Pan, Dawei Zhao, Xiaoming Wu, and Meihong Yang. Parallel optimization of Monte Carlo neutron transport method based on Sunway Bluelight II supercomputer. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07190-
1.

Chiyonobu:2025:HSC

[576] Miho Chiyonobu, Masami Takata, Kinji Kimura, and Yoshimasa Nakamura. High-speed computation method for condition numbers in the range restricted general minimum residual method. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-024-06868-2.

Babaiyan:2025:DRL

[577] Vahide Babaiyan and Omid Bushehrian.

A deep-reinforcement-learning-based strategy selection approach for faulttolerant offloading of delay-sensitive tasks in vehicular edge-cloud computing. The Journal of Supercomputing, 81 (5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07196-9.

Qi:2025:IQA

[578] Han Qi, Liyuan Wang, Dianshuo Fu, Abdullah Gani, and Changqing Gong. Improved qARM algorithm for frequent itemsets search in largescale data sets. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07209-
7.

Hou:2025:TMF

[579] Ke Hou, Mengying Wu, Wenjun Wu, and Linhao Huang. Topic mining and forecasting on patent map for GPU technology. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07215-
9.

Perez-Sarmiento:2025:PIO

[580] Luis A. Pérez-Sarmiento and Cuauhtemoc Mancillas-López. Parallel implementation of OCB using VAES and GPUs. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07179-
w.

Li:2025:KTT

[581] Shuang Li, Wen Yang, Luqun Li, Wenjing Li, Dazhi Li, and Zhehe Wang. Key technologies for task processing in ternary optical computer. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-025-07210-
0.

Song:2025:EDP

[582] Jin Song, Bowei Zeng, and Jun Ma. Ellipse detection and positioning in complex vision measurement environments. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07213-
x.

Ibrahimoglu:2025:FSN

[583] Nadir İbrahimoğlu, Amjad Osmani, Ali Ghaffari, Faruk Baturalp Günay, Tuğrul Çavdar, and Furkan Yıldız. FootprintNet: a Siamese network method for biometric identification using footprints. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07170-
5.

Bui:2025:CEA

[584] Thiloan Bui, Vanha Tran, Thaigiang Do, Hoangan Le, and Truongminh Ngo. A common and efficient algorithm for discovering high utility co-location patterns and their concise representation from massive spatial data. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07203-
z.

Xu:2025:LSC

[585] Zhijie Xu, Yuanyuan Qiu, and Jianqin Zhang. Learnable semantic completion and distance-induced graph matching for domain adaptive object detection. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07208-
                                        8.

Kawai:2025:SER

[586] Shunya Kawai, Eriko Maeda, Kazuki Yaguchi, Yasunori Osana, Takefumi Miyoshi, and Hironori Nakajo. Simulation environment for reconfigurable virtual accelerators using a field programmable gate array development environment. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06865-
                                        5.

Li:2025:DSE

[587] Cun Li, Shuhai Jiang, and Kangqian Zhou. DYR-SLAM: enhanced dynamic visual SLAM with YOLOv8 and RTAB–Map. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07191-
                                        0.

Wei:2025:UDA

[588] Xing Wei, Wenhao Jiang, Fan Yang, Chong Zhao, Yang Lu, Benhong Zhang, and Xiang Bi. Unsupervised domain adaptation via causalcontrastive learning. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07202-
0.

Ahmed:2025:JOB

[589] Manzoor Ahmed, Wali Ullah Khan, Mohammad Alamgeer, Eatedal Alabdulkreem, Shouki A. Ebad, Ali M. Al-Sharafi, Ashit Kumar Dutta, and Tahir Khurshaid. Joint optimization for 6G beyond diagonal IRS-assisted multicarrier NOMA vehicle-to-infrastructure communication. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07181-
2.

Zhong:2025:CDE

[590] Rui Zhong, Yang Cao, Enzhi Zhang, and Masaharu Munetomo. Competitive differential evolution with knowledge inheritance for single-objective humanpowered aircraft design. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06859-
3.

Sun:2025:ERS

[591] Dawei Sun, Yinuo Fan, Chengjun Guan, Jia Rong, Shang Gao, and Rajkumar Buyya. An elastic reconfiguration strategy for operators in distributed stream computing systems. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07216-
                                        8.

Chen:2025:YCU

[592] Ruolan Chen, Huibo Zhou, Hui Xie, and Bingyang Wang. YOLO-CE: an underwater low-visibility environment target detection algorithm based on YOLO11. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07205-
                                        x.

Song:2025:YFL

[593] Zijing Song, Xiaoyv Zhang, and Panlong Tan. YOLO-Fast: a lightweight object detection model for edge devices. The Journal of Supercomputing, 81(5): ??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07172-3.

Takata:2025:SMR

[594] Masami Takata and Miho Chiyonobu. Similar music recommendation method using Spotify API. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-024-06864-
                                        6.

Liu:2025:ISS

[595] Sicheng Liu, Sichen Tao, Kaiyu Wang, Zhenyu Lei, and Shangce Gao. Improved spherical search algorithm with memory-based dynamic population for optimization. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07095-
z.

Sun:2025:SCM

[596] Jianwen Sun, Bin Zhang, Hongyu Li, Lu Yuan, and Heyu Chang. T-Sanitation: contrastive masked autoencoder-based few-shot learning for malicious traffic detection. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07104-
1.

Fang:2025:DDH

[597] Juan Fang, Jingjing Li, and Ziyi Teng. Dhcm: a dynamic hierarchy coordination mechanism for memory optimization. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07224-
8.

Kumar:2025:BBF

[598] Ashish Kumar and Kakali Chatterjee. A blockchain-based framework with cryptographic tags for healthcare security. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07183-
0.

Tanaka:2025:PES

[599] Tomoaki Tanaka, Michiya Kato, Yasunori Osana, Takefumi Miyoshi, Jubee Tada, Kiyofumi Tanaka, and Hironori Nakajo. Preliminary evaluation of SHAVER: sharing vector registers with an accelerator. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-024-06863-
7.

Yong:2025:SCG

[600] Chen Yong, Hu Defeng, Xu Chao, Chen Nannan, and Liu Jianbo. Smart contract generation model based on code annotation and AST–LSTM tuning. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07186-
x.

Li:2025:ASF

[601] Hongjian Li, Gangfan Tan, Chenzi Wang, Yuan Gao, Tao Zhou, and Xiaolin Duan. Adaptive scheduling framework of streaming applications based on performance-to-cost ratio in heterogeneous cloud environment. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-025-07207-
9.

Wang:2025:ITA

[602] Xun Wang, Xianxi Zhu, Xiangyu Meng, Zeyang Zhu, Siyu Zhang, and Tao Song. Inference and training acceleration of deep learning partial differential equation solver. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07204-
y.

Anjali:2025:CDC

[603] Anjali and Anjana Gupta. Correction: Dynamic cloud model based on decision field theory. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07286-

8. See [556].

Wang:2025:DND

[604] Nianhao Wang and Han Wang. Dcsca-Net: a dual-branch network with enhanced cross-fusion and spatialchannel attention for precise medical image segmentation. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07239-
1.

Baawi:2025:HSC

[605] Salwa Shakir Baawi, Ekram Hakem, Abdulkareem A. Al-Hamzawi, Dhiah Al-Shammary, Ayman Ibaida, and Ahmed M. Mahdi. Hilbert similarity convex for efficient EEG feature selections. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07199-
6.

Topazal:2025:IDD

[606] S M Topazal, Shayla Islam, Raenu A./L. Kolandaisamy, Mohammad Kamrul Hasan, Ahmad Fadzil Ismail, Nur Hanis Sabrina Suhaimi, Huda Saleh Abbas, Muhammad Attique Khan, and Kamal Ali Alezabi. Intelligent device to device handover management techniques for 5G/ 6G and beyond. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07182-
1.

Du:2025:DDC

[607] Yongping Du, Qi Zhang, Shuyi Fu, Ying Hou, and Honggui Han. DCLMD: dynamic clustering and label mapping distribution for constructing in-context learning demonstrations. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07214-
w.

Huang:2025:TFD

[608] Yibo Huang, Weidong Qin, Zhiyong Li, and Qiuyu Zhang. Timefrequency dual-domain attention for acoustic echo cancellation. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07200-
2.

Du:2025:DRP

[609] Jinyang Du, Renyun Liu, Du Cheng, Qingliang Li, and Fanhua Yu. Deep recognition of partial differential equations based on reinforcement learning and genetic algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07162-
5.

Arasteh:2025:AST

[610] Bahman Arasteh, Keyvan Arasteh, and Ali Ghaffari. An automatic software test-generation method to discover the faults using fusion of machine learning and horse herd algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07219-
5.

Malekzadeh:2025:DPI

[611] Saba Malekzadeh, Saleh Yousefi, and Mir Saman Tajbakhsh. DDos prevention in IoT networks by analyzing source-side inter-bot traffic using deep learning techniques. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-025-07236-
                                        4.

Peng:2025:PML

[612] Chenchen Peng, Zeyu Xia, Shengbo Tang, Yifei Guo, Canqun Yang, Tao Tang, and Yingbo Cui. PVGwfa: a multi-level parallel sequence-to-graph alignment algorithm. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07184-
                                        z.

Wang:2025:ADS

[613] Yingqing Wang, Yanhua Liang, Yue Huang, and Guihe Qin. An automated data stream analysis framework for Internet of Vehicles based on online ensemble learning and two-dimensional fractal dimension. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07227-
                                        5.

Shang:2025:VAS

[614] Jiandong Shang, Zhuxin Wen, Haobo Hua, Hengliang Guo, Gang Wu, Wenlong Fan, Yang Guo, and Guangsheng Qin. VBATS: an adaptive strategy for grouped GEMM on GPUs. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07195-
                                        w.

Ranilla-Cortina:2025:EHP

[615] Sandra Ranilla-Cortina, Pedro Alonso-Jordá, Jesús Vigo-Aguiar, and José Ranilla. The evolution of highperformance computing: how AI and quantum computing are reshaping supercomputing. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07241-
7.

Hoorfar:2025:PCT

[616] Hamid Hoorfar, Adam C. Puche, and Istvan Merchenthaler. Publisher correction: Thermal image edge detection for AI-powered medical research imaging. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07257-

z. See [565].

Azarkasb:2025:MTL

[617] Seyed Omid Azarkasb and Seyed Hossein Khasteh. The mapping trick: leveraging RoboSoccer obstacle avoidance and navigation for advanced task scheduling solutions in foggy IoE ecosystems. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07147-
4.

Zhang:2025:AYS

[618] Hongying Zhang, Wentao Liu, and Enyao Chen. ASOD-YOLOX: a study on small object detection in aerial images based on YOLOX. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07243-
                                        5.

Anonymous:2025:JNn

[619] Anonymous. Journal navigation. The Journal of Supercomputing, 81(5):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Yin:2025:IDH

[620] Jie Yin, Shuning Zhang, and Guanlong Deng. An improved discrete Harris Hawks optimization algorithm for the no-wait job shop problem to minimize total weighted tardiness. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07234-
                                        6.

Meng:2025:HBV

[621] Lei Meng, Yueqiang Xu, and Fuhong Lin. How to balance the verification burden: a multi-hierarchical aggregate signature for drone swarms. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07222-
                                        w.

Verma:2025:PML

[622] Gurudatta Verma and Tirath Prasad Sahu. PSI-MFS: lightweight multiobjective feature selection for enhanced multi-label classification. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07163-
4.

Ding:2025:ITA

[623] Zhiyi Ding, Lei Sun, and Ruiyang Ding. Improving the transferability of adversarial examples via the high-level interpretable features for object detection. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07225-
7.

Jain:2025:SAR

[624] Sarika Jain, Priyanka Sukul, and Sven Groppe. A scientometric analysis of reviews on the Internet of Things. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07230-w.

Cao:2025:DNB

[625] Buqing Cao, Yating Yi, and Bing Tang. A deep Q-network-based edge service offloading in cloud-edge-terminal environment. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07228-
                                        4.

Yang:2025:TEP

[626] Kang Yang, Lili Zhang, and Pei Yu. Traffic environmental protection edge computing: a monitoring algorithm and system of truck black smoke emission in complex scene. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07276-
                                        w.

Xue:2025:SRB

[627] Peng Xue, Qian Gao, and Jun Fan. Social recommendation based on contrastive learning of hypergraph convolution. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07143-8.

Dhrir:2025:MLD

[628] Hanen Dhrir, Maha Charfeddine, and Habib M. Kammoun. Machine learningand deep learning-based anomaly detection in firewalls: a survey. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07212-
                                        y.

Ayyadi:2025:ACR

[629] Armin Ayyadi and Arezoo Jahani. Addressing cost and resource variability for big data task scheduling in heterogeneous cloud environments. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07248-
0.

Park:2025:ABA

[630] Ju-Won Park and Taeyoung Hong.

Application I/O behavior analysis on leadership cluster system. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07235-
5.

Koiso:2025:PMA

[631] Norie Koiso, Yuki Takemoto, and Masami Takata. Proposed method of acquiring train data for early-modern Japanese printed character recognizers. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-024-
06866-4.

Deng:2025:HAY

[632] Yan Deng, Lidong Huang, and Shunxiang Shi. A heterogeneous attention YOLO model for traffic sign detection. The Journal of Supercomputing, 81(6): ??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07256-0.

Sui:2025:MAE

[633] Xiao-Fei Sui, Si-Wen Zhang, and Xue-Lian Bai. Mayfly algorithm with elementary functions and mathematical spirals for task scheduling in cloud computing system. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07231-
                                        9.

Wang:2025:PPS

[634] Guangchao Wang, Zhaojun Li, and Changyou Guo. PDPA: a part-of-speech-driven prompt and attention method for aspect sentiment triplet extraction. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07188-
                                        9.

Bai:2025:RMA

[635] Fenhua Bai, Hongye Xu, and Chi Zhang. RBC-MSS: asynchronous broadcasting protocol based on multisecret sharing. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07211-
                                        z.

Li:2025:MMO

[636] Wei Li, Zeming Ping, and Lei Wang. Multimodal multiobjective optimization algorithm with fine-grained special crowding distance and dual-space selection mechanism. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07221-
x.

Cuevas:2025:NMA

[637] Erik Cuevas, Oscar A. González-Sánchez, and Alma Rodríguez-Vazquez. A novel metaheuristic algorithm using structured population and virtual particles. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07194-
x.

Yang:2025:ISO

[638] Yahao Yang, Yu Liu, and Zhenlong Zhao. An improved snake optimization algorithm based on hybrid strategy. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07258-
y.

Buttner:2025:APP

[639] Markus Büttner, Christoph Alt, and Vadym Aizinger. Analyzing performance portability for a SYCL implementation of the 2D shallow water equations. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

                                        article/10.1007/s11227-025-07063-
                                        7.

Serajgah:2025:PFT

[640] Saman Amini Serajgah and Ali-Akbar Ahmadi. Prescribed finite-time adaptive fuzzy-based fault-tolerant control of robotic manipulators using dynamic scaling factor. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07270-
                                        2.

Li:2025:UTL

[641] Min Li, Shuming Zhou, and Gaolin Chen. A unified temporal link prediction framework based on nonnegative matrix factorization and graph regularization. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07217-
                                        7.

Nazir:2025:EEE

[642] Ahsan Nazir, Jingsha He, and Muhammad Salman Pathan. Empirical evaluation of ensemble learning and hybrid CNN–LSTM for IoT threat detection on heterogeneous datasets. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07255-
                                        1.

Alsofyani:2025:ECS

[643] May Alsofyani and Liqiang Wang.

Evaluating ChatGPT’s strengths and limitations for data race detection in parallel programming via prompt engineering. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07237-
3.

Kang:2025:EFX

[644] Sung Won Kang and Cheong Hee Park. Effective federated XGBoost learning for multi-class classification in Non–IID environments. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07245-
3.

Li:2025:FDF

[645] Boyang Li, Huihuang Zhao, and Leyi Li. FDPNet: Deep forgery detection by leveraging multi-scale selfforgery images generating. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07229-
3.

Gao:2025:MSF

[646] Jialin Gao, Liqiang Zhu, and Yao Wang. Multi-scale feature enhanced detection of foreign object intrusions on railways. The Journal of Supercomputing, 81(6):??, April 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07254-
2.

Zhang:2025:ESI

[647] Yong Cai Zhang, Yonghua Lin, and Yih Bing Chu. The ensemble of selfinformation-based feature selection for heterogeneous data via k-nearest neighborhood rough set model. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07244-
4.

Uzelaltinbulat:2025:DSU

[648] Selin Uzelaltinbulat, Yasemin Kucukciloglu, and Boran Sekeroglu. Diagnosis of sacroiliitis using MR images with a simplified custom deep learning model. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07280-
0.

Lashkari:2025:UEM

[649] Arash Habibi Lashkari, Mohammad-Moein Shafi, and Ashley Barkworth. Unveiling evasive malware behavior: toward generating a multi-sources benchmark dataset and evasive malware behavior profiling using network traffic and memory analysis. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/

article/10.1007/s11227-025-07267-
x.

Salehpour:2025:OID

[650] Arash Salehpour, Mohammad Ali Balafar, and Alireza Souri. An optimized intrusion detection system for resourceconstrained IoMT environments: enhancing security through efficient feature selection and classification. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07253-3.

Javeed:2025:UCD

[651] Muhammad Awais Javeed, Dawei Li, and Muhammad Awais Ashraf. Urban city data delivery optimization in VNDN using a route-based caching approach with content-type awareness in intelligent transportation system. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07250-
6.

Tsai:2025:MSD

[652] Yu-Shiuan Tsai, Chia-Tung Tsai, and Jian-Hong Huang. Multi-scale detection of underwater objects using attention mechanisms and normalized Wasserstein distance loss. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07251-
5.

Li:2025:ATS

[653] Wenjuan Li, Genyuan Yang, and Qiwen Ni. Adaptive two-stage task offloading based on meta reinforcement learning for mobile edge computing. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07274-
                                        y.

Meng:2025:MFF

[654] Chao Meng, Yuhong Du, and Guangyu Dong. A multi-feature fusion supervoxel clustering segmentation method based on energy function. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07220-
                                        y.

Battat:2025:TDO

[655] Nadia Battat and Abdallah Makhoul. A trust-driven optimization model for reliable authorization in Hadoop Environment. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07268-
                                        w.

Luo:2025:VLM

[656] Wenyu Luo, Changxing Hou, and Annan Xuan. Vision and LiDAR multi-modal fusion beam prediction method for millimeter-wave communication system. The Journal of Supercomputing, 81(6):??, April 2025.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07284-
w.

Emara:2025:MOT

[657] Farouk A. Emara, Ahmed A. A.

Gad-Elrab, and M. A. Abd El-Baky. Multi-objective task scheduling algorithm for load balancing in cloud computing based on improved Harris hawks optimization. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07091-
3.

Cai:2025:MTL

[658] Yuanyuan Cai, Fei Tong, and Haitao Xiong. Multi-task learning framework using tri-encoder with caption prompt for multimodal aspectbased sentiment analysis. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07252-
4.

Sun:2025:NID

[659] Xiaolong Sun, Zhengyao Gu, and Junwei Ye. Network intrusion detection based on feature fusion of attack dimension. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07271-
1.

Yuan:2025:RAS

[660] Yan Yuan, Jinwei Xu, and Yong Wang. Research and application of smart insole assisted gait recognition technology. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07226-
                                        6.

Shankani:2025:FMN

[661] Rana Shankani, Maedeh Khalifavi, and Amirhossein Nikoofard. F^{2} multisense: a novel approach to fuzzy fusion in multisensor data to improve saving energy in WBSN. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07233-
                                        7.

Almeida:2025:AIH

[662] Fernando Almeida and Edet Okon. Assessing the impact of high-performance computing on digital transformation: benefits, challenges, and size-dependent differences. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07281-
                                        z.

Yang:2025:ANC

[663] Tianyu Yang, Yanmei Hu, and Xiangtao Li. An adaptive network construction for single-cell clustering. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07263-
1.

Souza:2025:MAE

[664] Fillipe Souza, Eduardo Tavares, and Carlos Araújo. A modelling approach for estimating energy consumption of NoSQL-based storage systems. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07298-4.

Dhiman:2025:SAE

[665] Shalini Dhiman, Sumitra Nayak, and Swarnendu Kumar Chakraborty. A survey analyzing encryption schemes for IoT security measures. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07285-
9.

Lu:2025:SPL

[666] Hangqing Lu, Quan Yan, and Li Wang.

A single-point and length-representation-based model for nested named entity recognition. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07240-

Chen:2025:PSC

[667] Yulei Chen, Dongwei Guo, and Jianhua Chen. A provably secure certificateless online/offline signature scheme for internet of things. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07294-
                                        8.

Wang:2025:SFS

[668] Zhicheng Wang, Longge Wang, and Jinhu Wu. Semantic feature space construction for unsupervised fewshot image classification. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07315-
                                        6.

Xue:2025:RRR

[669] Qifan Xue, Feng Yang, and Weigong Zhang. Rethink Reynolds’ rules: flockinspired network for vehicle trajectory prediction. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07265-
                                        z.

Anonymous:2025:JNo

[670] Anonymous. Journal navigation. The Journal of Supercomputing, 81(6):??, April 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Thakkar:2025:MME

[671] Riddhi Thakkar and Madhuri Bhavsar.

MeghMesa: a multilevel elasticity for streaming applications in cloud. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07232-8.

Liu:2025:ISO

[672] Fangjun Liu, Huifu Zhang, and Shangfeng Mo. Instruction selection optimization for VLIW architecture based on classification node merging. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07323-6.

Zhou:2025:EAL

[673] Jinfeng Zhou, Xiaoqin Zeng, and Haoran Zhu. Enhancing aspect-level sentiment analysis through the integration of local context interdependencies and syntactic quality compensation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07301-
y.

Eldakhly:2025:OML

[674] Nabil M. Eldakhly. Optimized machine learning for real-time, non-invasive blood pressure monitoring. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07330-
                                        7.

Feng:2025:TTM

[675] Chunhui Feng, Yongxiang Zhong, and Xiaolong Liu. TMTC: trusted multimodal transformer classification framework for video frame deletion detection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07307-6.

Liu:2025:MMI

[676] Quankai Liu, Haifeng Sang, and Wangxing Chen. MGFormer: a multiinformation-based GRU-transformer network for pedestrian trajectory prediction. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07304-9.

Liu:2025:NFT

[677] Wenfei Liu, Jiafei Liu, and Qi Wang. A novel fault-tolerant technique for star graph-based interconnection networks. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07272-0.

Hong:2025:PUP

[678] Hanshu Hong, Yibo Sun, and Zhixin Sun. PB-UOKM: a policy-based updatable oblivious key management scheme for secure and practical data sharing in remote storage. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07332-
5.

Dixit:2025:QBC

[679] Vimal Dixit and Krishnan Rajkumar. QENN: breast cancer prediction using quantum-enhanced neural network. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07331-
6.

Xia:2025:CCL

[680] Limin Xia and Qiyue Xiao. Comprehensive context learning for twostage human-object interaction detection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07308-5.

Taher:2025:GFQ

[681] R. Ben Taher and M. Lassri. On the generalized Fibonacci quaternions and the linear difference equations with periodic coefficients. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07283-
x.

Erho:2025:TPJ

[682] Joseph Agaroghenefuoma Erho, Bunakiye Richard Japheth, and Juliana Iworikumo Consul. Theoretical perspective to the juggling sequence rotation performance pattern from system memory locality reference viewpoint. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07261-
                                        3.

Ao:2025:AVM

[683] Wenxiu Ao, Zhongmei Wang, and Liang Zheng. An audio-visual multimodal adaptive balanced learning method based on gradient modulation. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07311-w.

Misra:2025:QQE

[684] Shreshtha Misra and Poonam Rani.

QEKLR: quantum-enhanced kernel logistic regression for classification. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07266-y.

Xu:2025:ETO

[685] Tao Xu, Chaoyue Chen, and Dongdong Ma. Exponential-trigonometric optimization algorithm with multi-strategy fusion for UAV three-dimensional path planning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https:

//link.springer.com/article/10.
1007/s11227-025-07335-2.

Osaba:2025:EAQ

[686] Eneko Osaba, Esther Villar-Rodriguez, and Izaskun Oregi. Exploring the application of quantum technologies to industrial and real-world use cases. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07320-
9.

Liu:2025:SAD

[687] Tianwei Liu and Jingwen Luo. Semanticaware dynamic scene tracking and reconstruction for mobile robots leveraging visual SLAM. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07343-
2.

Wang:2025:DAT

[688] Yuan Wang, Turdi Tohti, and Askar Hamdulla. Domain-adaptive transfer network for visual-textual cross-domain sentiment classification. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07273-
z.

Li:2025:DUC

[689] Xinxin Li, Zhenping Lan, and Aixia Yuan. DPF-Unet: a CNN-swin transformer fusion network for 3D brain tumor segmentation in MRI images. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.springer.
                                        com/article/10.1007/s11227-025-
                                        07339-y.

Khan:2025:AZT

[690] Talha Ahmed Khan, Khizar Abbas, and Wang-Cheol Song. Accelerating zero-touch automation and optimization of beyond 5G services: deep learning and intent-based networking fusion. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07260-
                                        4.

Khomami:2025:CGGb

[691] Mohammad Mehdi Daliri Khomami, Ali Mohammad Saghiri, and Mohammad Reza Meybodi. Cellular goore game with multiple learning automata in each cell and its applications. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07249-
                                        z.

Chen:2025:ITA

[692] Jialong Chen, Jijun Tong, and Shudong Xia. An implicit tubular-aware network for coronary artery segmentation.

The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07312-9.

Dong:2025:OKE

[693] Yuyu Dong, Fang’ai Liu, and Xuejian Gao. Optimizing keyphrase extraction with dependency relation-aware attention graph convolutional networks. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07303-w.

Xu:2025:LTR

[694] Nengxiang Xu, Yuxiang Chen, and Nam Ling. N-Lock: a transactionreleased shard reconfiguration protocol with zero-knowledge proof. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07317-
4.

Wang:2025:MET

[695] Bing Wang, Chaoqi Cai, and Youming Zhang. MapsTSF: efficient traffic prediction via hybrid mamba 2transformer spatiotemporal modeling and cross adaptive periodic sparse forecasting. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07328-
1.

Sun:2025:MID

[696] Bo Sun, Zhuo Hao, and Jun He. Mutual introspective distillation for unbiased scene graph generation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07246-
                                        2.

Li:2025:YEO

[697] Yanshun Li, Quanfeng Zheng, and Mengwei Guo. YOLO-EDGE: an object detection algorithm for traffic scenarios. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07275-
                                        x.

Bain:2025:OAT

[698] Debanjali Bain and Biswanath Dutta. Ontological approach towards discovering and recommending COVID-19 therapeutics, risk factors, and drug interactions. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07302-
                                        x.

Bakhoda:2025:MOS

[699] Saeedeh Bakhoda, Mohammad Abdollahi Azgomi, and Mohammad Reza Ebrahimi Dishabi. Multi-objective scheduling of heterogeneous parallel systems using the VITS algorithm. The Journal of Supercomputing, 81(7):

??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07293-9.

Zhang:2025:ACD

[700] Zheqing Zhang, Hongzhi Li, and Kuan-Ching Li. AAMB: a cross-domain identity authentication scheme based on multilayered blockchains in IoMT. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07309-4.

Wang:2025:DDF

[701] Feng Wang, Chenglong Wang, and Yuxuan Zhang. Data-driven fuzzy logic control method for improved USV path planning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07318-
3.

Li:2025:ICO

[702] Lang Li, Li Hu, and Jin Li. Inverse correction-optimized vertical federated unlearning. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07310-
x.

Li:2025:SST

[703] Xiguang Li, Yuhan Zhu, and Liang Zhao. STGEN: spatio-temporal generalized aggregation networks for traffic accident prediction. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07288-
6.

Wang:2025:SIC

[704] Jun Wang, Chenhao Sun, and Xiaoli Li. Spatiotemporal information cooperative interaction network for video salient object detection. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07314-
7.

Czarnul:2025:ORA

[705] Paweł Czarnul, Marcel Antal, and Gordana Rakić. Optimization of resource-aware parallel and distributed computing: a review. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07295-
7.

He:2025:OMB

[706] Yunbin He, Chenglong Li, and Fuwei Guo. An oversampling method based on the weighting of minority class clusters. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07325-4.

Yaqoubi:2025:MCS

[707] Zahra Yaqoubi, Behrouz Shahgholi Ghahfarokhi, and Mojtaba Mahdavi. Mobile crowd sensing based spectrum monitoring with privacy protection and malicious behavior detection using hyperledger fabric and identity mixer. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07300-
z.

Gharehbagheri:2025:APS

[708] Parviz Gharehbagheri, Hamid Haj Seyyed Javadi, and Hamidreza Bakhshi. An alternative proposed single error detection algorithm for the single parity check algorithm. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07305-
8.

Ashi:2025:IDG

[709] H. A. Ashi and Noufe H. Aljahdaly.

Investigation of the damped geophysical KdV equation using the explicit exponential time differencing method. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07289-5.

Saad:2025:EMO

[710] Abla Saad, Osama Abdel-Raouf, and Ahmed Kafafy. Enhanced multiobjective scheduling for heterogeneous computing platforms using hybrid MOEAD with SA and TS-guided initialization strategies. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07278-
8.

Zhan:2025:CDR

[711] MengJun Zhan, XiaoYao Xie, and ZhongYuan Guo. Change detection of remote sensing building images based on dual-domain strip attention DML–UNet network. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07344-
1.

Alfawaz:2025:VFC

[712] Oruba Alfawaz, Ali A. El-Moursy, and Ahmed M. Khedr. VFCkM: a federated clustering framework based on k-means algorithm for vertically partitioned data with shared attributes. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07296-
6.

Verma:2025:AVL

[713] Priyanka Verma, Donna O’Shea, and John G. Breslin. ABIDS-VEM: leveraging an equilibrium optimizer and data ramification in association with ensemble learning for anomaly-based intrusion detection system. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07292-
w.

Masaud:2025:DCS

[714] Mohamed Ahmed Ali Masaud, Selçuk Alparslan Avci, and Javad Rahebi. Detecting cyberattacks in smart grids using VGG-16 and whale–fisher mantis optimization algorithm (WOA-FMO). The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07345-0.

Florez:2025:LCP

[715] Arantzazu Florez, Itsaso Rodríguez-Moreno, and Basilio Sierra. Lynxsight: change-point detection through different distance-based common spatial patterns. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07282-
y.

Liang:2025:CPC

[716] Bo Liang, Fujiang Yuan, and Jie Gao. Cs-pbft: a comprehensive scoringbased Practical Byzantine Fault Tolerance consensus algorithm. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL https://link.springer.com/
article/10.1007/s11227-025-07342-

He:2025:NCD

[717] Jiali He and Mideth B. Abisado. A novel conditional discrimination index approach for feature selection in partially labeled hybrid data. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07333-
                                        4.

Shang:2025:HWT

[718] Lei Shang, Huan Lei, and Wenyuan Yang. HIFNet: wavelet transformenhanced UAV object detection in complex conditions. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07338-
                                        z.

Nooraee:2025:TPO

[719] Mohsen Nooraee, Hamidreza Ghaffari, and Fateme Zarisfi Kermani. Tiny-ParsBERT: an optimized hybrid model for efficient sentiment analysis in Persian texts. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s11227-025-07297-
                                        5.

Derakhshan:2025:HNM

[720] Mohammad Hossein Derakhshan, Yadollah Ordokhani, and J. F. Gómez-Aguilar. A hybrid numerical method with high accuracy to solve a timespace diffusion model in terms of the Caputo and Riesz fractional derivatives. The Journal of Supercomputing, 81(7): ??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.
com/article/10.1007/s11227-025-
07287-7.

Anonymous:2025:JNp

[721] Anonymous. Journal navigation. The Journal of Supercomputing, 81(7):??, May 2025. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).